### v1

# Low Regularity of Self-Similar Solutions of Two-Dimensional Riemann problems with Shocks for the Isentropic Euler system

Abstract.
We are concerned with the low regularity of self-similar solutions of two-dimensional Riemann problems for the isentropic Euler system. We establish a general framework for the analysis of the local regularity of such solutions for a class of two-dimensional Riemann problems for the isentropic Euler system, which includes the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the four-shock Riemann problem. We prove that it is not possible that both the density and the velocity are in $H^{1}$ in the subsonic domain for the self-similar solutions of these problems in general. This indicates that the self-similar solutions of the Riemann problems with shocks for the isentropic Euler system are of much more complicated structure than those for the Euler system for potential flow; in particular, the density and the velocity are not necessarily continuous in the subsonic domain. The proof is based on a regularization of the isentropic Euler system to derive the transport equation for the vorticity, a renormalization argument extended to the case of domains with boundary, and DiPerna-Lions-type commutator estimates.
Key words and phrases:
Low regularity, transonic shock, free boundary, compressible flow, potential flow, self-similar, conservation laws, shock reflection-diffraction, regular shock reflection, Prandtl reflection, Lighthill diffraction, nonlinear, fine properties
2010 Mathematics Subject Classification:
Primary: 35M10, 35M12, 35B65, 35L65, 35J70, 76H05, 35L67, 35B45, 35B40, 35B36, 76N10; Secondary: 35R35, 35L03, 35L04, 35J67, 76L05, 76J20, 76N20, 76G25
1. Introduction
We are concerned with the regularity of self-similar solutions of the Riemann problems with shocks for the isentropic Euler system in a general setting, including several fundamental shock problems such as the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the four-shock Riemann problem. In 1860, Riemann first considered a special initial value problem with two constant states separated at the origin for the one-dimensional isentropic Euler system in [43] – now known as the Riemann problem; this Riemann problem has played a fundamental role in the mathematical theory of hyperbolic systems of conservation laws, since its solutions are building blocks and asymptotic attractors of general global entropy solutions. Since then, a systematic theory of one-dimensional hyperbolic systems of conservation laws has been established; see [5, 6, 11, 22, 30, 35, 41] and the reference cited therein. However, multi-dimensional Riemann problems are much more complicated and completely different from the one-dimensional case. Even for the two-dimensional Riemann problem with four constant states given in the four quadrants for the Euler system, nineteen genuinely different configurations have been identified; see [9, 10, 12, 36, 48, 47]. Since then, rigorous global results for the $2$-D four-quadrant Riemann problem for the Euler system were only done by Li-Zheng in [38, 39]. See also [20, 44] for the 2-D Riemann problem for Chaplygin gases.
On the other hand, the regular shock reflection problem is a different type of multi-dimensional Riemann problems – a lateral Riemann problem that involves the wedge boundaries. Shock reflection-diffraction phenomena were first presented by Ernst Mach [42] in 1878, and experimental, computational, and asymptotic analysis has shown that various patterns of shock reflection may occur, including regular and Mach reflection; see[15] and the references cited therein.
More recently, the existence of global regular reflection solutions was established for the Euler system for potential flow in [14, 15], and further properties of these solutions were proved in [1, 17]. In particular, the solution has a high regularity in the subsonic domain $\Omega$ (where the solution is not a constant state); see Figs. 4.24.2 in which density $\rho$ and velocity ${\mathbf{v}}$ are in $C^{\alpha}(\overline{\Omega})\cap C^{\infty}(\Omega)$ for some $\alpha\in(0,1)$. More precisely, solution $(\rho,{\mathbf{v}})$ is in a weighted Hölder space which implies that $(\rho,{\mathbf{v}})$ is in the Sobolev space $W^{1,p}(\Omega)$ for some $p>2$. In addition, the regularity of the curved reflected-diffracted shock was shown to be $C^{\infty}$ in the interior and $C^{2,\alpha}$ up to the endpoints, which is also expected to hold for the case of the isentropic Euler system. See also [2, 3, 26] for the Prandtl reflection-diffraction configuration, [13] for the Lighthill diffraction problem, and [12] for the four-shock Riemann problem.
In contrast, a remarkable phenomenon was first observed by Serre in [44] which showed by a formal calculation that, in the case of the isentropic Euler system, the regular shock reflection solutions develop vortical singularity, specifically that the vorticity cannot be in $L^{2}(\Omega)$. This implies that the velocity cannot be in $W^{1,2}(\Omega)$, i.e., $H^{1}(\Omega)$, which is lower than the regularity of the velocity for the case of potential flow discussed above. This in particular allows that the velocity can be discontinuous, which at least can not be excluded by the Sobolev embeddings. Since the calculation is formal, it is important to find out whether the low regularity indeed necessarily holds for the regular shock reflection solutions.
In this paper, under the natural assumptions on the regularity of the self-similar solutions in the subsonic domain $\Omega$ near the shocks, we rigorously prove that it is not possible indeed that both the density and the velocity are in $H^{1}(\Omega)$. The argument is based on the vortical singularity calculation in [44]. We first apply this calculation to the regularized solutions carefully for which the calculation can be rigorously justified; however, the additional error terms appear due to the regularization. Then we develop DiPerna-Lions-type commutator estimates to control the error terms when the regularization parameter tends to zero. With this, we then prove the lower regularity property of the self-similar solutions, by employing renormalization argument. These self-similar solutions include the above-mentioned shock reflection problems: the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the four-shock Riemann problem.
This indicates that the self-similar solutions of the Riemann problems with shocks for the isentropic Euler system have much more complicated structure than the corresponding solutions for the Euler system for potential flow; in particular, the density and the velocity are not necessarily continuous in $\Omega$, at least their continuity can not be obtained directly by the Sobolev embeddings. On the other hand, this argument allows the possibility that $(\rho,{\mathbf{v}})\in W^{1,p}(\Omega)$ for some or even all $p\in[1,2)$, in which case there are no shocks in $\Omega$.
This paper is organized as follows: In §2, we derive the isentropic Euler system in self-similar coordinates and then present the notion of entropy solutions in the new coordinates. In §3, we first formulate a general framework for analyzing the low regularity of entropy solutions of the Riemann problems in Definition 3.1 and then establish our main theorem of this paper, Theorem 3.2, for the entropy solutions. Then, in §4, we employ the main theorem, Theorem 3.2, for the general framework established in §3 to several fundamental transonic shock problems including the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the Riemann problem with four-shock interactions in §4.1–§4.4, respectively. The general framework and the main theorem, as well as ideas and approaches, developed in this paper should be useful for solving other similar low regularity problems for solutions of nonlinear partial differential equations.
2. The Isentropic Euler System and Entropy Solutions with Shocks
In this section, we first derive the isentropic Euler system in self-similar coordinates and then present the notion of entropy solutions, as well as the Rankine-Hugoniot conditions and the corresponding entropy condition across a shock (as a free boundary), in self-similar coordinates.
2.1. The Isentropic Euler System
As in [22, 44] (see also [15]), the isentropic Euler equations consist of the conservation laws of mass and momentum:
$\begin{cases}\partial_{t}\rho+\mbox{\rm div}(\rho{\mathbf{u}})=0,\\[2.84526pt] \partial_{t}(\rho{\mathbf{u}})+\mbox{\rm div}(\rho{\mathbf{u}}\otimes{\mathbf{u}})+\nabla p=0,\end{cases}$ (2.1)
where $\rho$ is the density, ${\mathbf{u}}=(u_{1},u_{2})$ is the velocity, and $p$ is the pressure. The constitutive relation between pressure $p$ and density $\rho$ is through the $\gamma$-law relation: $p=\frac{\rho^{\gamma}}{\gamma}$ after scaling, and the adiabatic exponent $\gamma>1$ is a given constant.
If an initial-boundary value problem is invariant under the self-similar scaling:
$({\mathbf{u}},p,\rho)=({\mathbf{u}},p,\rho)(\frac{\mathbf{x}}{t}),$ (2.2)
we introduce self-similar variables ${\boldsymbol{\xi}}=(\xi_{1},\xi_{2})=\frac{\mathbf{x}}{t}\in\mathbb{R}^{2}$ and the pseudo-velocity ${\mathbf{v}}={\mathbf{u}}-{\boldsymbol{\xi}}$. Then we obtain the isentropic Euler system for self-similar flow $(\rho,{\mathbf{v}})=(\rho,{\mathbf{v}})({\boldsymbol{\xi}})$ in the form:
$\mbox{\rm div}(\rho\,{\mathbf{v}})+2\rho=0,$ (2.3)
$\mbox{\rm div}(\rho\,{\mathbf{v}}\otimes{\mathbf{v}})+3\rho{\mathbf{v}}+\nabla p=0.$ (2.4)
If $(\rho,{\mathbf{v}})\in C^{1}$, then we can combine the equations above to rewrite (2.4) as
$({\mathbf{v}}\cdot\nabla){\mathbf{v}}+{\mathbf{v}}+\nabla h(\rho)=0,$ (2.5)
where $h(\rho)=\frac{\rho^{\gamma-1}}{\gamma-1}$ is the enthalpy. Note that the speed of sound is $c=\rho^{\frac{\gamma-1}{2}}=\sqrt{(\gamma-1)h(\rho)}$.
2.2. Entropy Solutions and the Rankine-Hugoniot Conditions
We consider solutions with shocks, which satisfy (2.3)–(2.4) in the following weak sense:
Definition 2.1.
Let $\Lambda\subset\mathbb{R}^{2}$ be a domain. Then $(\rho,{\mathbf{v}})\in L^{\infty}(\Lambda)$ is an entropy solution of system (2.3)–(2.4) if the following conditions hold:
  1. (i)
    $(\rho,{\mathbf{v}})$ is a weak solution: For any test functions $\phi\in C^{\infty}_{\rm c}(\Lambda)$ and $\zeta\in C^{\infty}_{\rm c}(\Lambda;\,\mathbb{R}^{2})$,
    $\int_{\Lambda}(\rho{\mathbf{v}}\cdot\nabla\phi-2\rho\phi)\,{\rm d}{\boldsymbol{\xi}}=0,$ (2.6)
    $\int_{\Lambda}\big{(}\rho{\mathbf{v}}\otimes{\mathbf{v}}:D\zeta-3\rho{\mathbf{v}}\cdot\zeta+p\,\mbox{\rm div}\zeta\big{)}\,{\rm d}{\boldsymbol{\xi}}=0,$ (2.7)
    where we have used the notation $A:B=\sum_{i,j=1}^{2}a_{ij}b_{ij}$ for $2\times 2$ matrices $A$ and $B$.
  2. (ii)
    $(\rho,{\mathbf{v}})$ satisfies the entropy condition: For any non-negative test function $\psi\in C^{\infty}_{\rm c}(\Lambda)$,
    $\int_{\Lambda}\Big{(}\big{(}\frac{1}{2}\rho|{\mathbf{v}}|^{2}+\rho e(\rho)+p(\rho)\big{)}{\mathbf{v}}\cdot\nabla\psi-2\big{(}\rho|{\mathbf{v}}|^{2}+\rho e(\rho)+p(\rho)\big{)}\psi\Big{)}\,{\rm d}{\boldsymbol{\xi}}\geq 0,$ (2.8)
    where the internal energy $e(\rho)$ is defined by $p(\rho)=\rho^{2}e^{\prime}(\rho)$, i.e., $e(\rho)=\frac{\rho^{\gamma-1}}{\gamma(\gamma-1)}$ for the polytropic case with $p(\rho)=\frac{\rho^{\gamma}}{\gamma}$.
Suppose that $S$ is a smooth curve in domain $\Lambda$. An entropy solution of (2.3)–(2.4), which is $C^{1}$ near and up to $S$ on both sides of $S$, satisfies the following Rankine-Hugoniot (R-H) conditions on $S$:
$[\rho{\mathbf{v}}\cdot\boldsymbol{\nu}]=0,\quad\,[(\rho{\mathbf{v}}\cdot\boldsymbol{\nu}){\mathbf{v}}+p\boldsymbol{\nu}]={\bf 0}\qquad\,\,\mbox{on $S$},$ (2.9)
[…]
If $\rho^{+}{\mathbf{v}}^{+}\cdot\boldsymbol{\nu}=0$ on $S$, then $\rho^{-}{\mathbf{v}}^{-}\cdot\boldsymbol{\nu}=0$ on $S$ by the first equality in (2.9), and $[p]=0$ from the second equality in (2.9) which implies that $[\rho]=0$. This discontinuity is called a vortex sheet. In this case, $[{\mathbf{v}}\cdot\boldsymbol{\tau}]\neq 0$ on $S$, unless the solution is continuous across $S$, where $\boldsymbol{\tau}$ is a unit tangent vector to $S$.
If $\rho^{+}{\mathbf{v}}^{+}\cdot\boldsymbol{\nu}\neq 0$ on $S$, then $\rho^{-}{\mathbf{v}}^{-}\cdot\boldsymbol{\nu}\neq 0$ on $S$ by the first equality in (2.9). This discontinuity is called a shock. In this case, $\rho{\mathbf{v}}\cdot\boldsymbol{\nu}$ is continuous across $S$ from the first equality in (2.9), and $[{\mathbf{v}}\cdot\boldsymbol{\tau}]=0$ and $[p]=-\rho{\mathbf{v}}\cdot\boldsymbol{\nu}[{\mathbf{v}}\cdot\boldsymbol{\nu}]$ from the second equality in (2.9). This implies that $[{\mathbf{v}}\cdot\boldsymbol{\nu}]\neq 0$, unless the solution is continuous across $S$, so that $[\rho]\neq 0$ from the first equality in (2.9). Also, it follows from $\rho^{\pm}>0$ on $S$ that $({\mathbf{v}}^{+}\cdot\boldsymbol{\nu})({\mathbf{v}}^{-}\cdot\boldsymbol{\nu})>0$ on $S$.
Thus, we have shown that the following properties hold, in addition to (2.9), in the case $\rho^{\pm}>0$ on $S$:
Shock: $({\mathbf{v}}^{+}\cdot\boldsymbol{\nu})({\mathbf{v}}^{-}\cdot\boldsymbol{\nu})>0,\;\;[{\mathbf{v}}\cdot\boldsymbol{\tau}]=0,\;\;[{\mathbf{v}}\cdot\boldsymbol{\nu}]\neq 0,\;\;[\rho]\neq 0;$ (2.10)
Vortex sheet: ${\mathbf{v}}^{+}\cdot\boldsymbol{\nu}={\mathbf{v}}^{-}\cdot\boldsymbol{\nu}=0,\;\;[{\mathbf{v}}\cdot\boldsymbol{\tau}]\neq 0,\;\;[\rho]=0.$ (2.11)
Furthermore, the entropy condition (2.8) is required across shock $S$ separating the smooth states $(\rho^{\pm},{\mathbf{v}}^{\pm})$ defined in domains $\Lambda^{\pm}$ with $\rho^{\pm}>0$. Then it follows from the direct calculation through (2.8)–(2.10) and the choice of orientation of the unit normal ${\boldsymbol{\nu}}$ on $S$ to point from $\Lambda^{-}$ to $\Lambda^{+}$ that, on $S$,
$\mbox{If }{\mathbf{v}}^{-}\cdot{\boldsymbol{\nu}}>0,\;\;\mbox{ then }\;{\mathbf{v}}^{-}\cdot{\boldsymbol{\nu}}>{\mathbf{v}}^{+}\cdot{\boldsymbol{\nu}}>0,\,\,\rho^{-}<\rho^{+},\;\;{\mathbf{v}}^{-}\cdot{\boldsymbol{\nu}}>c^{-},\;\;{\mathbf{v}}^{+}\cdot{\boldsymbol{\nu}}<c^{+},$ (2.12)
where the last two inequalities are shown e.g., in [44, Theorem 2.2] (the argument is given there for steady solutions; the proof applies to the self-similar case because the self-similar Rankine-Hugoniot conditions are the same), since $p(\rho)$ is increasing and convex, implied by $p(\rho)=\frac{\rho^{\gamma}}{\gamma}$ with $\gamma>1$.
We also give a definition of entropy solutions of system (2.3)–(2.4) in $\Lambda$ with the slip boundary conditions:
[…]
where ${\boldsymbol{\nu}}$ is the outer normal to $\partial\Lambda$. The motivation is the following: Suppose that $(\rho,{\mathbf{v}})\in C^{1}$ and $\partial\Lambda\in{\rm Lip}$ satisfy (2.3)–(2.4) in $\Lambda$ and (2.13). Then it follows that (2.6) and
$\int_{\Lambda}\big{(}\rho{\mathbf{v}}\otimes{\mathbf{v}}:D\zeta-3\rho{\mathbf{v}}\cdot\zeta+p\,\mbox{\rm div}\zeta\big{)}\,{\rm d}{\boldsymbol{\xi}}-\int_{\partial\Lambda}p\,\zeta\cdot{\boldsymbol{\nu}}\,{\rm d}l=0$ (2.14)
are satisfied for any test function $\zeta\in C^{\infty}_{\rm c}(\mathbb{R}^{2};\,\mathbb{R}^{2})$. Note that here the test function is not required to vanish on $\partial\Lambda$. Based on that, we define the notion of entropy solutions of the boundary value problem (2.3)–(2.4) and (2.13) as follows:
Definition 2.2.
Let $\Lambda\subset\mathbb{R}^{2}$ be a domain with a Lipschitz boundary $\partial\Lambda$. Let $(\rho,{\mathbf{v}})\in L^{\infty}(\Lambda)$, and let $\rho\in BV_{\rm loc}(\Lambda\cap{\mathcal{N}}_{r}(\partial\Lambda))$ for some $r>0$. Then $(\rho,{\mathbf{v}})$ is an entropy solution of system (2.3)–(2.4) with slip boundary condition (2.13) if (2.6), (2.14), and (2.8) are satisfied for any test functions $\phi\in C^{\infty}_{\rm c}(\mathbb{R}^{2})$, $\zeta\in C^{\infty}_{\rm c}(\mathbb{R}^{2};\,\mathbb{R}^{2})$, and $\psi\in C^{\infty}_{\rm c}(\Lambda)$ with $\psi\geq 0$, respectively.
3. Low Regularity of Self-Similar Solutions of the Riemann Problems with Shocks for the Isentropic Euler System
In this section, we first formulate a general framework for analyzing the low regularity of entropy solutions (i.e., self-similar solutions of admissible structure) of the Riemann problems in Definition 3.1, motivated by the solutions of the physically fundamental Riemann problems described in §4 below. Then we establish our main theorem, Theorem 3.2, for the entropy solutions $(\rho,{\mathbf{v}})$ in the general framework, which will be applied to understanding the low regularity of the solutions of the Riemann problems including the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the Riemann problem with four shock interactions. These are achieved by carefully analyzing the vorticity function $\omega:=\partial_{1}v_{2}-\partial_{2}v_{1}$ for the pseudo-velocity ${\mathbf{v}}=(v_{1},v_{2})$.
Figure 3.1. The Riemann Problem in a General Setting
More precisely, as shown in Fig. 3.1, we consider a Riemann problem in a self-similar wedge domain $\Lambda$ in $\mathbb{R}^{2}$ with the wedge-vertex being the origin, which also includes the case of the whole space. That is, in polar coordinates,
either Λ=2 or $\mbox{either $\,\,\Lambda=\mathbb{R}^{2}\,\,\,$ or $\,\,\Lambda=\big{\{}(r,\theta)\;:\;\theta^{-}<\theta<\theta^{+}\big{\}}$}.$ (3.1)
For a Riemann problem in a domain $\Lambda$ whose boundary contains a wedge-boundary, the initial data are the given constant velocity and density in each sub-sector of $\Lambda$ such that the Rankine-Hugoniot conditions hold on the lines separating them. Motivated by the expected configurations of self-similar solutions of the Riemann problems in §4, we consider solutions in self-similar coordinates of the following form, in terms of the density and pseudo-velocity $(\rho,{\mathbf{v}})$ with $\mathcal{N}_{r}(\Gamma):=\{{\boldsymbol{\xi}}\,:\,\mbox{\rm dist}\{{\boldsymbol{\xi}},\Gamma\}<r\}$:
Definition 3.1.
We say that the vector function $(\rho,{\mathbf{v}})$ on $\Lambda$ is an entropy solution of a Riemann problem of admissible structure if $(\rho,{\mathbf{v}})\in L^{\infty}(\Lambda)$, with $\rho\in BV_{\rm loc}(\Lambda\cap{\mathcal{N}}_{r}(\partial\Lambda))$ for some $r>0$, is an entropy solution of system (2.3)–(2.4) with the slip boundary condition (2.13) in the sense of Definition 2.2, which satisfies the following properties:
  1. (i)
    There exists an open, bounded, connected set $\Omega\subset\Lambda$, an integer $M\geq 1$, and open connected sets $\Lambda_{i}$, $i=1,\cdots,M$, pairwise disjoint, such that
    $\overline{\Lambda}\setminus\Omega=\cup_{i=1}^{M}\overline{\Lambda_{i}},$
    and $(\rho,{\mathbf{v}})$ is a constant state in each $\Lambda_{i}$, i.e., $(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{i},{\mathbf{u}}_{i}-{\boldsymbol{\xi}})$ in $\Lambda_{i}$, where $\rho_{i}>0$ is a constant and ${\mathbf{u}}_{i}$ is a constant vector.
  2. (ii)
    If, for $i\neq j$, sets $\Lambda_{i}$ and $\Lambda_{j}$ have common boundary within $\Lambda$; that is, if $\partial\Lambda_{i}\cap\partial\Lambda_{j}\cap\Lambda\neq\emptyset$, then the corresponding constant states are not equal to each other: $(\rho_{i},{\mathbf{v}}_{i})\neq(\rho_{j},{\mathbf{v}}_{j})$.
  3. (iii)
    $\partial\Omega$ is Lipschitz. Denote
    $\Gamma^{\mbox{\rm\footnotesize ext}}=\partial\Omega\cap\partial\Lambda,\qquad\Gamma^{\mbox{\rm\footnotesize int}}=\partial\Omega\cap\Lambda,$
    so that $\partial\Omega=\Gamma^{\mbox{\rm\footnotesize ext}}\cup\Gamma^{\mbox{\rm\footnotesize int}}$. It is possible that $\Gamma^{\mbox{\rm\footnotesize ext}}=\emptyset$ (in particular, this is true when $\partial\Lambda=\emptyset$).
  4. (iv)
    If $\partial\Lambda\neq\emptyset$, then ${\Gamma^{\mbox{\rm\footnotesize ext}}}=\cup_{i=1}^{N_{1}}\overline{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}}$, where ${N_{1}}\geq 1$ and each $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ is a relatively open segment of straight line. Segments $\{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\}_{i=1}^{N_{1}}$ are disjoint and, if $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ and $\Gamma^{\mbox{\rm\footnotesize ext}}_{j}$ have a common endpoint, then the interior angle (for $\Omega$) is within $(0,2\pi)$ at that point. Since $\Lambda$ is a wedge domain with the wedge-vertex at the origin, only one pair among segments $\{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\}$ may have a common endpoint that is $\mathcal{P}_{0}=(0,0)$.
  5. (v)
    $\overline{\Gamma^{\mbox{\rm\footnotesize int}}}=\cup_{i=1}^{N_{2}}\overline{\Gamma^{\mbox{\rm\footnotesize int}}_{i}}$, where ${N_{2}}\geq 1$ and each $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ is a relatively open segment of curve, and segments $\{\Gamma^{\mbox{\rm\footnotesize int}}_{i}\}_{i=1}^{N_{2}}$ are disjoint. Each $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ is $C^{2}$ in its relative interior and $C^{1}$ up to the endpoints. Moreover, $\mathcal{P}_{0}\notin\overline{\Gamma^{\mbox{\rm\footnotesize int}}}$.
  6. (vi)
    ${\mathbf{v}}\in C\big{(}{\mathcal{N}}_{\sigma}(\Gamma^{\mbox{\rm\footnotesize int}})\cap\overline{\Omega}\big{)}$ for some $\sigma>0$.
  7. (vii)
    $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is a shock and ${\mathbf{v}}\cdot{\boldsymbol{\nu}}\leq-C^{-1}$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, and ${\mathbf{v}}\cdot{\boldsymbol{\nu}}\leq 0$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ for $i=2,\cdots,{N_{2}}$, where ${\mathbf{v}}$ is taken from the $\Omega$–side and ${\boldsymbol{\nu}}$ denotes the outer unit normal with respect to $\Omega$.
  8. (viii)
    There exists a point $\hat{P}$ in the relative interior of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ such that the curvature of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is non-zero at $\hat{P}$, and $({\mathbf{v}}\cdot{\boldsymbol{\tau}})(\hat{P})\neq 0$, where ${\boldsymbol{\tau}}(\hat{P})$ is a unit tangent vector to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at $\hat{P}$.
Then we have the following main theorem.
Theorem 3.2 (Main Theorem).
Let $(\rho,{\mathbf{v}})$ be a solution of a Riemann problem of admissible structure in the sense of Definition 3.1. Assume that $(\rho,{\mathbf{v}})$ satisfy
  1. (i)
    $(\rho,{\mathbf{v}})\in C^{1}(({\mathcal{N}}_{\sigma}(\Gamma^{\mbox{\rm\footnotesize int}})\cap\overline{\Omega})\setminus\partial_{\rm p}\Gamma^{\mbox{\rm\footnotesize int}})\cap C^{0,1}({\mathcal{N}}_{\sigma}(\Gamma^{\mbox{\rm\footnotesize int}})\cap\overline{\Omega})$ for some $\sigma>0$, where $\partial_{\rm p}\Gamma^{\mbox{\rm\footnotesize int}}$ denotes the set of endpoints of the curve segments $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ for $i=1,\cdots,{N_{2}}$.
  2. (ii)
    $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
  3. (iii)
    The flow is subsonic on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ from the $\Omega$–side: $|{\mathbf{v}}|<c$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, where $c=\rho^{\frac{\gamma-1}{2}}$ is the speed of sound.
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
Proof.
We divide the proof into six steps.
1. We prove the theorem by establishing that, under the assumption:
$(\rho,{\mathbf{v}})\in H^{1}(\Omega),$ (3.2)
the calculations of vortical singularity can rigorously be justified, which leads to the contradiction.
Under assumption (3.2), vorticity $\omega=\nabla\times{\mathbf{v}}$ satisfies $\omega\in L^{2}(\Omega)$ and
$X:=\frac{\omega}{\rho}\in L^{2}(\Omega),$ (3.3)
where we have used (3.2) and assumption (ii). In addition, using (3.2) and assumptions (i)–(ii), we see that
The left-hand sides of equations (2.3) and (2.5) are in $L^{2}(\Omega)$ and the equations hold a.e. in $\Omega$. (3.4)
2. We first formally derive the equations and identities that vorticity $\omega$ satisfies, and then prove them rigorously.
Taking the curl of (2.5), we formally obtain the equation:
${\mathbf{v}}\cdot\nabla\omega+(1+\mbox{\rm div}\,{\mathbf{v}})\,\omega=0.$ (3.5)
Combining with the first equation of (2.3), we formally have
${\mathbf{v}}\cdot\nabla\big{(}\frac{\omega}{\rho}\big{)}=\frac{\omega}{\rho}.$ (3.6)
If $f\in C^{1}(\mathbb{R})$, multiplying the last equation by $f^{\prime}(\frac{\omega}{\rho})$ and then combining with the first equation of (2.3), we formally derive
$\mbox{\rm div}\big{(}\rho f(\frac{\omega}{\rho}){\mathbf{v}}\big{)}=\rho g(\frac{\omega}{\rho}),$ (3.7)
where
$g(s):=sf^{\prime}(s)-2f(s).$ (3.8)
Now we are going to show that equation (3.7) holds under the present assumptions in the weak sense defined as follows: Using notation (3.3), equation (3.7) can be written as
$\mbox{\rm div}\left(\rho f(X){\mathbf{v}}\right)=\rho g(X)$ (3.9)
for any $f\in C^{1}(\mathbb{R})$ with $g$ defined by (3.8). From the weak form (2.6) of the conservation law of mass, which holds for all $\phi$ specified in Definition 2.2, and the regularity assumption in (3.2), we obtain that $\rho{\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ holds ${\mathcal{H}}^{1}$a.e. on $\Gamma^{\mbox{\rm\footnotesize ext}}$ in the sense of traces. Then, using assumption (ii), we recover condition (2.13) on $\partial\Lambda\cap\partial\Omega$:
${\mathbf{v}}\cdot{\boldsymbol{\nu}}=0\qquad\,\,{\mathcal{H}}^{1}\mbox{--a.e. on }\;\Gamma^{\mbox{\rm\footnotesize ext}}.$ (3.10)
We next show that, under the present assumptions, (3.9) holds weakly in $\Omega$ in the sense that
$\int_{\Omega}\big{(}\rho f(X){\mathbf{v}}\cdot\nabla\zeta+\rho g(X)\zeta\big{)}{\rm d}{\boldsymbol{\xi}}-\int_{\Gamma^{\mbox{\rm\footnotesize int}}}\rho f(X){\mathbf{v}}\cdot{\boldsymbol{\nu}}\,\zeta\,{\rm d}l=0\qquad\mbox{ for all }\;\zeta\in C^{\infty}(\mathbb{R}^{2})$ (3.11)
for any $f\in C^{1}(\mathbb{R}^{1})$ with $\|f^{\prime}\|_{C^{0,1}(\mathbb{R}^{1})}<\infty$.
Notice that the boundary integral in (3.11) is taken along a part of $\partial\Omega$. Formally, the boundary integral along the remaining part $\Gamma^{\mbox{\rm\footnotesize ext}}$ of $\partial\Omega$ is expected to vanish by (3.10). Equation (3.11) shows that
$\rho f(X){\mathbf{v}}\cdot{\boldsymbol{\nu}}=0\qquad\mbox{on $\Gamma^{\mbox{\rm\footnotesize ext}}$ in the weak sense.}$.
Furthermore, the integrand in the boundary integral in (3.11) is well-defined by assumption (i) and Definition 3.1(v).
3. To prove (3.11), we introduce smooth approximations of $(\rho,{\mathbf{v}})$. Since we work in the bounded domain $\Omega$ and need the boundary condition ${\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ on $\Gamma^{\mbox{\rm\footnotesize ext}}$ and other properties to hold for the approximating functions for the argument below, we first construct a specific sequence of smooth approximations of $(\rho,{\mathbf{v}})$.
We first extend ${\mathbf{v}}$ by the reflection across the straight boundary segments $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$, $i=1,\cdots,{N_{1}}$, so that the normal component is extended by the odd reflection and condition (3.10) is used to conclude that the extended function remains in $H^{1}$, and the tangential component is extended by the even reflection.
Denote by ${\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}$ the unit inner normal vector to $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ with respect to $\Lambda$, $i=1,\cdots,{N_{1}}$, respectively, and by ${\boldsymbol{\tau}}_{i}^{\mbox{\rm\footnotesize ext}}$ the unit tangential vectors to $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$.
For each $r>0$ and $i=1,\cdots,{N_{1}}$, we first define an extension of ${\mathbf{v}}$ from ${\mathcal{N}}_{r}(\Gamma^{\mbox{\rm\footnotesize ext}}_{i})\cap\Omega$ across $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ by reflection. However, there is the following issue: Denote
$W_{i}^{r}:=\big{\{}{\boldsymbol{\xi}}-s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}\;:\;{\boldsymbol{\xi}}\in\Gamma^{\mbox{\rm\footnotesize ext}}_{i},\;s\in(-r,0),\;{\boldsymbol{\xi}}+s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}\in\Omega\big{\}},$
the image of ${\mathcal{N}}_{r}(\Gamma^{\mbox{\rm\footnotesize ext}}_{i})\cap\Omega$ under the reflection across $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$. We note that, if $\mathcal{P}_{0}=(0,0)$ is a common endpoint of $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ and $\Gamma^{\mbox{\rm\footnotesize ext}}_{j}$, and the interior angle $\theta_{\mathcal{P}_{0}}$ of $\Lambda$ at $\mathcal{P}_{0}$ is larger than $\pi$, i.e., $\theta_{\mathcal{P}_{0}}\in(\pi,2\pi)$, then $W_{i}^{r}\cap W_{j}^{r}\neq\emptyset$ for any $r>0$. Moreover, if $\theta_{\mathcal{P}_{0}}\in(\frac{3}{2}\pi,2\pi)$, then $W_{i}^{r}\cap\Omega\neq\emptyset$ for any $r>0$. Of course, in general, the extension of ${\mathbf{v}}$ by reflection across $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ into $W_{i}^{r}$ does not match with ${\mathbf{v}}$ in $W_{i}^{r}\cap\Omega$ or with extension of ${\mathbf{v}}$ by reflection across $\Gamma^{\mbox{\rm\footnotesize ext}}_{j}$ in $W_{i}^{r}\cap W_{j}^{r}$, so the region needs to be restricted.
In the case above, for each $r>0$, we have
$(W_{i}^{r}\cap\Omega)\setminus B_{Lr}(\mathcal{P}_{0})=\emptyset,\quad(W_{i}^{r}\cap W_{j}^{r})\setminus B_{Lr}(\mathcal{P}_{0})=\emptyset\qquad\mbox{for some }L=L(\theta_{\mathcal{P}_{0}})\geq 0.$ (3.12)
In fact, this is true with $L=\mbox{cosec}\,{\theta_{\mathcal{P}_{0}}}$ if $\theta_{\mathcal{P}_{0}}\in(\pi,2\pi)$ and $L=0$ if $\theta_{\mathcal{P}_{0}}\in(0,\pi]$. Also, there exists $r_{1}>0$ such that, for $i=1,\cdots,{N_{1}}$,
$\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\cap B_{(L+1)r_{1}}(\mathcal{P}_{0})=\emptyset\qquad\mbox{if $\mathcal{P}_{0}$ is not an endpoint of }\Gamma^{\mbox{\rm\footnotesize ext}}_{i}.$ (3.13)
Now, for each $r\in(0,r_{1}]$, define
$V_{i}^{r}:=W_{i}^{r}\setminus\overline{B_{Lr}(\mathcal{P}_{0})},$
which is an open connected set. We extend ${\mathbf{v}}$ to $V_{i}^{r}$ by
${\mathbf{v}}({\boldsymbol{\xi}}-s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}):=-({\mathbf{v}}\cdot{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}})({\boldsymbol{\xi}}+s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}){\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}+({\mathbf{v}}\cdot{\boldsymbol{\tau}}_{i}^{\mbox{\rm\footnotesize ext}})({\boldsymbol{\xi}}+s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}){\boldsymbol{\tau}}_{i}^{\mbox{\rm\footnotesize ext}}$ (3.14)
for all ${\boldsymbol{\xi}}\in\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ and $s\in(0,r)$ such that ${\boldsymbol{\xi}}+s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}\in\Omega\setminus\overline{B_{Lr}(\mathcal{P}_{0})}$. From the definitions of $W_{i}^{r}$ and $V_{i}^{r}$, since segment $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ lies on the line passing through $\mathcal{P}_{0}$, the expression above defines ${\mathbf{v}}$ on the whole region $V_{i}^{r}$. Also, it follows that, for all ${\boldsymbol{\xi}}\in\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ and $s\in(0,r)$ such that ${\boldsymbol{\xi}}+s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}\in\Omega\setminus\overline{B_{Lr}(\mathcal{P}_{0})}$,
$\begin{split}&({\mathbf{v}}\cdot{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}})({\boldsymbol{\xi}}-s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}):=-({\mathbf{v}}\cdot{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}})({\boldsymbol{\xi}}+s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}),\\[2.84526pt] &({\mathbf{v}}\cdot{\boldsymbol{\tau}}_{i}^{\mbox{\rm\footnotesize ext}})({\boldsymbol{\xi}}-s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}):=({\mathbf{v}}\cdot{\boldsymbol{\tau}}_{i}^{\mbox{\rm\footnotesize ext}})({\boldsymbol{\xi}}+s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}),\end{split}$ (3.15)
that is, ${\mathbf{v}}\cdot{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}$ is extended by the odd reflection, and ${\mathbf{v}}\cdot{\boldsymbol{\tau}}_{i}^{\mbox{\rm\footnotesize ext}}$ is extended by the even reflection across $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$.
Similarly, for each $i=1,\cdots,{N_{1}}$ and $r\in(0,r_{1}]$, we extend $\rho$ to $V_{i}^{r}$ by the even reflection across $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$: For all ${\boldsymbol{\xi}}\in\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ and $s\in(0,r)$ such that ${\boldsymbol{\xi}}+s{\boldsymbol{\nu}}_{i}^{\mbox{\rm\footnotesize ext}}\in\Omega\setminus\overline{B_{Lr}(\mathcal{P}_{0})}$,
$\rho({\boldsymbol{\xi}}-s{\boldsymbol{\nu}}):=\rho({\boldsymbol{\xi}}+s{\boldsymbol{\nu}}),$ (3.16)
where we recall that $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ is a relatively open line segment.
Using (3.2) and assumption (ii), and noting that ${\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ to make the odd extension of ${\mathbf{v}}\cdot{\boldsymbol{\nu}}$ across $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$, we find that the extended $(\rho,{\mathbf{v}})$ satisfies $(\rho,{\mathbf{v}})\in H^{1}((\Omega\cup\overline{V_{i}^{r_{1}}})^{0})\cap L^{\infty}(\Omega\cup\overline{V_{i}^{r_{1}}})$ and $\rho\geq C_{0}^{-1}$ on $\Omega\cup\overline{V_{i}^{r_{1}}}$.
Combining the extensions for $i=1,\cdots,{N_{1}}$, and using (3.12)–(3.13), we obtain that, for each $r\in(0,r_{1}]$, $(\rho,{\mathbf{v}})$ is extended to the domain:
$\Omega^{\rm ext}_{r}=\big{(}\Omega\cup(\cup_{i=1}^{N_{1}}\overline{V_{i}^{r}})\big{)}^{0},$ (3.17)
and the extension satisfies $(\rho,{\mathbf{v}})\in H^{1}(\Omega^{\rm ext}_{r})\cap L^{\infty}(\Omega^{\rm ext}_{r})$ and $\rho\geq C_{0}^{-1}$ on $\Omega^{\rm ext}_{r}$.
Moreover, by (3.4) and the explicit structure (3.14)–(3.16) of the extension, it follows that property (3.4) holds for the extended $(\rho,{\mathbf{v}})$ in $\Omega^{\rm ext}_{r}$ for any $r\in(0,r_{1}]$:
The left-hand sides of Eqs. (2.3) and (2.5) are (3.18)
in $L^{2}(\Omega^{\rm ext}_{r})$, and the equations hold a.e. in $\Omega^{\rm ext}_{r}$.
Indeed, if ${\boldsymbol{\xi}}\in\Gamma^{\mbox{\rm\footnotesize ext}}$ and $s\in(0,r)$, and if $(\rho,{\mathbf{v}})$ is differentiable at ${\boldsymbol{\xi}}+s{\boldsymbol{\nu}}$, then clearly $(\rho,{\mathbf{v}})$ is differentiable at ${\boldsymbol{\xi}}-s{\boldsymbol{\nu}}$, and (2.3) and (2.5) hold at ${\boldsymbol{\xi}}-s{\boldsymbol{\nu}}$, which can be seen by the explicit calculation or using the standard symmetries of the isentropic Euler system (2.3)–(2.4).
Thus, we have shown the following results in this step:
Lemma 3.3.
Under assumptions (i)–(ii) of Theorem 3.2 and (3.2), there exists $r_{1}>0$ depending only on $\Omega$ such that, for any $r\in(0,r_{1}]$, there exists an extension of $(\rho,{\mathbf{v}})$ into $\Omega^{\rm ext}_{r}$, still denoted as $(\rho,{\mathbf{v}})$, such that
  1. (a)
    $(\rho,{\mathbf{v}})\in H^{1}(\Omega^{\rm ext}_{r})\cap L^{\infty}(\Omega^{\rm ext}_{r})$ with $\rho\geq C_{0}^{-1}$ in $\Omega^{\rm ext}_{r};$
  2. (b)
    For any ${\boldsymbol{\xi}}\in\Gamma^{\mbox{\rm\footnotesize ext}}$ and $s\in(0,r)$ such that ${\boldsymbol{\xi}}+s{\boldsymbol{\nu}}\in\Omega\setminus\overline{B_{Lr}(\mathcal{P}_{0})}$ for $L$ from (3.12), the following odd/even reflection properties hold:
    $\begin{cases}({\mathbf{v}}\cdot{\boldsymbol{\nu}})({\boldsymbol{\xi}}-s{\boldsymbol{\nu}})=-({\mathbf{v}}\cdot{\boldsymbol{\nu}})({\boldsymbol{\xi}}+s{\boldsymbol{\nu}}),\\[2.84526pt] ({\mathbf{v}}\cdot{\boldsymbol{\tau}})({\boldsymbol{\xi}}-s{\boldsymbol{\nu}})=({\mathbf{v}}\cdot{\boldsymbol{\tau}})({\boldsymbol{\xi}}+s{\boldsymbol{\nu}}),\\[2.84526pt] \rho({\boldsymbol{\xi}}-s{\boldsymbol{\nu}})=\rho({\boldsymbol{\xi}}+s{\boldsymbol{\nu}}).\end{cases}$ (3.19)
  3. (c)
    Property (3.18) holds.
4. We now show that equation (3.9) is satisfied in the weak sense:
Lemma 3.4.
Under the assumptions of Theorem 3.2 and (3.2), equation (3.9) holds weakly in the sense (3.11) for each $f\in C^{1}(\mathbb{R})$ with $\|f^{\prime}\|_{C^{0,1}(\mathbb{R})}<\infty$ and $g$ defined by (3.8).
This can be proved as follows: We reduce $r_{1}$ to obtain, in addition to (3.13), that, for $j=1,\cdots,{N_{2}}$,
$\Gamma^{\mbox{\rm\footnotesize int}}_{j}\cap B_{(L+1)r_{1}}(\mathcal{P}_{0})=\emptyset\qquad\mbox{and}\qquad r_{1}<\sigma,$ (3.20)
where we have used that $\mathcal{P}_{0}\notin\overline{\Gamma^{\mbox{\rm\footnotesize int}}}$ by Definition 3.1(v), $\sigma$ is from assumption (i) of Theorem 3.2, and $L$ is from (3.12). Fix $f\in C^{1}(\mathbb{R})$ with $\|f^{\prime}\|_{C^{0,1}(\mathbb{R})}\leq C$, and let $g$ be defined by (3.8).
4.1. We first prove (3.11) in the case when the smooth function $\zeta$ satisfies that there exists $r\in(0,r_{1})$ such that
$\zeta\equiv 0\qquad\mbox{in ${\mathcal{N}}_{r}(\Gamma^{\mbox{\rm\footnotesize int}})\cup B_{2Lr}(\mathcal{P}_{0})$}.$ (3.21)
Thus, we need to consider $(\rho,{\mathbf{v}})$ only in the region:
$\Omega^{r}:=\Omega\setminus\big{(}\overline{{\mathcal{N}}_{r}(\Gamma^{\mbox{\rm\footnotesize int}})}\cup\overline{B_{2Lr}(\mathcal{P}_{0})}\big{)}.$ (3.22)
Note that $\Omega^{r}\Subset\Omega^{\rm ext}_{r}$. Let
$\delta:=\mbox{\rm dist}(\partial\Omega^{\rm ext}_{r},\;\Omega^{r}).$
Then $\delta\in(0,r]$. We now mollify $({\mathbf{v}},\omega)$ and $\rho$ in $\Omega^{r}$, by using the extension of $(\rho,{\mathbf{v}})$ into $\Omega^{\rm ext}_{r}$ constructed in Lemma 3.3 and the corresponding extension of $\omega=\partial_{1}v_{2}-\partial_{2}v_{1}$. In order to achieve that the mollified ${\mathbf{v}}$ satisfies the boundary condition (3.10), we use $\eta\in C^{\infty}_{\rm c}(\mathbb{R}^{2})$ with $\int_{\mathbb{R}^{2}}\eta({\boldsymbol{\xi}})\,{\rm d}{\boldsymbol{\xi}}=1$ of form $\eta({\boldsymbol{\xi}})=g(|{\boldsymbol{\xi}}|)$ for some $g\in C^{\infty}_{\rm c}(\mathbb{R})$, for example, the standard mollifier. Then we define $\eta_{\varepsilon}({\boldsymbol{\xi}})=\frac{1}{\varepsilon^{2}}\eta(\frac{{\boldsymbol{\xi}}}{\varepsilon})$ for $\varepsilon>0$, and denote $F_{\varepsilon}:=F*\eta_{\varepsilon}$ for various functions $F$, specifically
${\mathbf{v}}_{\varepsilon}={\mathbf{v}}*\eta_{\varepsilon},\quad\omega_{\varepsilon}=\omega*\eta_{\varepsilon},\quad\rho_{\varepsilon}=\rho*\eta_{\varepsilon}\quad(i(\rho))_{\varepsilon}=i(\rho)*\eta_{\varepsilon}\qquad\,\,\mbox{for $\varepsilon\in(0,\frac{\delta}{2})$}.$
Now, since $(\rho,{\mathbf{v}})\in H^{1}(\Omega^{\rm ext}_{r})\cap L^{\infty}(\Omega^{\rm ext}_{r})$ with $\rho\geq C^{-1}$ by Lemma 3.3(a)–(b), and $\Omega^{r}\Subset\Omega^{\rm ext}_{r}$, we obtain that, for each $\varepsilon\in(0,\frac{\delta}{2})$,
$\rho_{\varepsilon},\,{\mathbf{v}}_{\varepsilon},\,(i(\rho))_{\varepsilon}\in H^{1}(\Omega^{r})\cap L^{\infty}(\Omega^{r}),\quad\|(\rho_{\varepsilon}\,{\mathbf{v}}_{\varepsilon})\|_{L^{\infty}(\Omega^{r})}\leq\|(\rho\,{\mathbf{v}})\|_{L^{\infty}(\Omega^{\rm ext}_{r})},$ (3.23)
$(\rho_{\varepsilon},\,{\mathbf{v}}_{\varepsilon})\to(\rho,\,{\mathbf{v}})\,\,\,\,\mbox{in $H^{1}(\Omega^{r})$},\quad\,\omega_{\varepsilon}\to\omega\,\,\,\,\mbox{in $L^{2}(\Omega^{r})\qquad\,\,$ as $\varepsilon\to 0$}.$
Moreover, from (3.14)–(3.15) and $\eta({\boldsymbol{\xi}})=g(|{\boldsymbol{\xi}}|)$, we have
${\mathbf{v}}_{\varepsilon}\cdot{\boldsymbol{\nu}}=0\qquad\mbox{on $\cup_{i=1}^{N_{1}}\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\cap\partial\Omega^{r}\quad$ for all $\varepsilon\in(0,\delta)$}.$ (3.24)
Now, from (3.18), we see that, in $\Omega^{r}$,
$\mbox{\rm div}(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon})+2\rho_{\varepsilon}+r^{(1)}_{\varepsilon}=0,$ (3.25)
$({\mathbf{v}}_{\varepsilon}\cdot\nabla){\mathbf{v}}_{\varepsilon}+{\mathbf{v}}_{\varepsilon}+\nabla(i(\rho))_{\varepsilon}+{\bf r}^{(2)}_{\varepsilon}=0,$ (3.26)
where
$r^{(1)}_{\varepsilon}=(\mbox{\rm div}\,(\rho{\mathbf{v}}))_{\varepsilon}-\mbox{\rm div}\,(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon}),\qquad{\bf r}^{(2)}_{\varepsilon}=\big{(}({\mathbf{v}}\cdot\nabla){\mathbf{v}}\big{)}_{\varepsilon}-({\mathbf{v}}_{\varepsilon}\cdot\nabla){\mathbf{v}}_{\varepsilon}.$
The functions on the left-hand side of (3.25)–(3.26) are smooth. Taking the curl of (3.26), we obtain
${\mathbf{v}}_{\varepsilon}\cdot\nabla\omega_{\varepsilon}+(1+\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon})\omega_{\varepsilon}+\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}=0\qquad\,\,\mbox{in $\Omega^{r}$}.$ (3.27)
Denote
$X^{(\varepsilon)}:=\frac{\omega_{\varepsilon}}{\rho_{\varepsilon}}.$
Using (3.23) and the lower bound of $\rho$ in $\Omega^{\rm ext}_{r}$ by Lemma 3.3(a), we obtain that, for $\varepsilon\in(0,\frac{\delta}{2})$,
$X^{(\varepsilon)}\in L^{2}(\Omega^{r}),\qquad\;X^{(\varepsilon)}\to X\,\,\,\,\mbox{in $L^{2}(\Omega^{r})\,\,\,$ as $\varepsilon\to 0$}.$ (3.28)
We apply the definition of $X^{(\varepsilon)}$ in the first equality, along with equations (3.25) and (3.27) in the second equality, to compute:
$\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon}\cdot\nabla X^{(\varepsilon)}$ $={\mathbf{v}}_{\varepsilon}\cdot\nabla\omega_{\varepsilon}-\nabla\rho_{\varepsilon}\cdot{\mathbf{v}}_{\varepsilon}X^{(\varepsilon)}$
$=\rho_{\varepsilon}X^{(\varepsilon)}+r^{(1)}_{\varepsilon}-\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}.$
From this and (3.25), we have
$\mbox{\rm div}\big{(}\rho_{\varepsilon}f(X^{(\varepsilon)}){\mathbf{v}}_{\varepsilon}\big{)}$ $=f^{\prime}(X^{(\varepsilon)})\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon}\cdot\nabla X^{(\varepsilon)}+f(X^{(\varepsilon)})\mbox{\rm div}(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon})$
$=\rho_{\varepsilon}\big{(}X^{(\varepsilon)}f^{\prime}(X^{(\varepsilon)})-2f(X^{(\varepsilon)})\big{)}+{\mathcal{R}}_{\varepsilon},$ (3.29)
where
${\mathcal{R}}_{\varepsilon}=f^{\prime}(X^{(\varepsilon)})\big{(}r^{(1)}_{\varepsilon}-\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}\big{)}-f(X^{(\varepsilon)})r^{(1)}_{\varepsilon}.$ (3.30)
Recalling definition (3.8) of $g(\cdot)$, we rewrite (3.29) as
$\mbox{\rm div}\big{(}\rho_{\varepsilon}f(X^{(\varepsilon)}){\mathbf{v}}_{\varepsilon}\big{)}$ $=\rho_{\varepsilon}g(X^{(\varepsilon)})+{\mathcal{R}}_{\varepsilon}\qquad\mbox{in $\Omega^{r}$}.$
Let $\zeta\in C^{\infty}(\mathbb{R}^{2})$ satisfy (3.21). Multiply the last equation by $\zeta$, integrate over $\Omega$, and integrate by parts via using (3.21) and (3.24) to obtain
$\int_{\Omega^{r}}\big{(}\rho_{\varepsilon}f(X^{(\varepsilon)}){\mathbf{v}}_{\varepsilon}\cdot\nabla\zeta+\big{(}\rho_{\varepsilon}g(X^{(\varepsilon)})+{\mathcal{R}}_{\varepsilon}\big{)}\zeta\big{)}\,{\rm d}\xi=0,$ (3.31)
where we have used (3.21)–(3.22) to restrict the domain to $\Omega^{r}$.
4.2. To send $\varepsilon\to 0$ in (3.31), we note the following facts: Since $f\in C^{1}(\mathbb{R})$ with $\|f^{\prime}\|_{C^{0,1}(\mathbb{R})}<\infty$, then, by (3.8), $g\in C(\mathbb{R})$ with $\mbox{Lip}(g)<\infty$ on $\mathbb{R}$. It follows from (3.3) and (3.28) that, for all $\varepsilon\in(0,\frac{\delta}{2})$,
$\|f(X^{(\varepsilon)})\|_{L^{2}(\Omega^{r})}+\|f^{\prime}(X^{(\varepsilon)})\|_{L^{\infty}(\Omega^{r})}+\|g(X^{(\varepsilon)})\|_{L^{2}(\Omega^{r})}\leq C,$ (3.32)
and
$(f(X^{(\varepsilon)}),\,g(X^{(\varepsilon)}))\to\,(f(X),g(X))\qquad\mbox{in $L^{2}(\Omega^{r})\,\,$ as $\varepsilon\to 0$}.$ (3.33)
Next, we show that
${\mathcal{R}}_{\varepsilon}\to 0\qquad\mbox{in $L^{1}(\Omega^{r})\;\;$ as $\varepsilon\to 0$}.$ (3.34)
From (3.30) and (3.32) and the fact that $\Omega^{r}$ is a bounded domain, in order to prove (3.34), it suffices to show that
$r^{(1)}_{\varepsilon}\to 0\qquad\qquad\mbox{in $L^{2}(\Omega^{r})\;\;$ as $\varepsilon\to 0$},$ (3.35)
$\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}\to 0\qquad\,\mbox{in$L^{1}(\Omega^{r})\;\;$ as $\varepsilon\to 0$}.$ (3.36)
To show (3.35), we first note that $\Omega^{r}\Subset\Omega^{\rm ext}_{r}$ and $\rho{\mathbf{v}}\in H^{1}(\Omega^{\rm ext}_{r})\cap L^{\infty}(\Omega^{\rm ext}_{r})$ by Lemma 3.3(a), so that
$(\mbox{\rm div}\,(\rho{\mathbf{v}}))_{\varepsilon}-\mbox{\rm div}\,(\rho{\mathbf{v}})\to 0\qquad\mbox{in $L^{2}(\Omega^{r})\;$ as $\varepsilon\to 0$}.$
Thus, it remains to show that
$\mbox{\rm div}\,(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon})-\mbox{\rm div}\,(\rho{\mathbf{v}})\to 0\qquad\mbox{in $L^{2}(\Omega^{r})\;$ as $\varepsilon\to 0$}.$ (3.37)
We first show that
$\mbox{\rm div}\,(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon})-\mbox{\rm div}\,(\rho{\mathbf{v}}_{\varepsilon})\to 0\qquad\mbox{in $L^{2}(\Omega^{r})\;$ as$\varepsilon\to 0$}.$
Indeed,
$\mbox{\rm div}\,(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon})-\mbox{\rm div}\,(\rho{\mathbf{v}}_{\varepsilon})=\nabla(\rho_{\varepsilon}-\rho)\cdot{\mathbf{v}}_{\varepsilon}+(\rho_{\varepsilon}-\rho)(\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon}-\mbox{\rm div}\,{\mathbf{v}})+(\rho_{\varepsilon}-\rho)\mbox{\rm div}\,{\mathbf{v}}.$
In the argument below, we use (3.23). We see that $\nabla(\rho_{\varepsilon}-\rho)\cdot{\mathbf{v}}_{\varepsilon}\to 0$ in $L^{2}(\Omega^{r})$ as $\varepsilon\to 0$, because $\nabla\rho_{\varepsilon}\to\nabla\rho$ in $L^{2}(\Omega^{r})$ and ${\mathbf{v}}_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$. Also, $(\rho_{\varepsilon}-\rho)(\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon}-\mbox{\rm div}\,{\mathbf{v}})\to 0$ in $L^{2}(\Omega^{r})$, because $\rho-\rho_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$ and $\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon}\to\mbox{\rm div}\,{\mathbf{v}}$ in $L^{2}(\Omega^{r})$. Finally, $(\rho_{\varepsilon}-\rho)\mbox{\rm div}\,{\mathbf{v}}\to 0$ in $L^{2}(\Omega^{r})$, because $\rho-\rho_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$ and converge to zero a.e. in $\Omega_{r}$, and $\mbox{\rm div}\,{\mathbf{v}}\in L^{2}(\Omega^{r})$, so that
$\int_{\Omega_{r}}(\rho_{\varepsilon}-\rho)^{2}(\mbox{\rm div}\,{\mathbf{v}})^{2}\,{\rm d}\xi\to 0\qquad\mbox{as $\varepsilon\to 0$},$
by the dominated convergence theorem. The convergence:
$\mbox{\rm div}\,(\rho{\mathbf{v}}_{\varepsilon})-\mbox{\rm div}\,(\rho{\mathbf{v}})\to 0\qquad\mbox{in $L^{2}(\Omega^{r})\;\;$ as $\varepsilon\to 0$}$
can be shown similarly. This completes the proof of (3.37), which leads to (3.35).
Now we show (3.36). Note that
$\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}=\partial_{\xi_{1}}\big{(}({\mathbf{v}}\cdot\nabla v_{2})_{\varepsilon}-({\mathbf{v}})_{\varepsilon}\cdot\nabla(v_{2})_{\varepsilon}\big{)}-\partial_{\xi_{2}}\big{(}({\mathbf{v}}\cdot\nabla v_{1})_{\varepsilon}-({\mathbf{v}})_{\varepsilon}\cdot\nabla(v_{1})_{\varepsilon}\big{)}.$
Then (3.36) follows from Lemma A.1 in Appendix A with $p=q=2$, $b=v_{j}$, and $u=\partial_{j}v_{k}$ for the corresponding $j,k=1,2$, and $i=3-k$.
Combining the results above, (3.34) is now proved. Then, sending $\varepsilon\to 0$ in (3.31) and using (3.33)–(3.34), we obtain
$\int_{\Omega}\big{(}\rho f(X){\mathbf{v}}\cdot\nabla\zeta+\rho g(X)\zeta\big{)}\,{\rm d}{\boldsymbol{\xi}}=0,$
which is equivalent to (3.11), by using (3.21).
4.3. Next, we prove (3.11) in the case when the smooth function $\zeta$ satisfies:
$\zeta\equiv 0\qquad\mbox{in $\Omega\setminus{\mathcal{N}}_{r}(\Gamma^{\mbox{\rm\footnotesize int}})\,\,$ for some $r\in(0,r_{1})$.}$ (3.38)
Then, by (3.20) and assumption (i) of Theorem 3.2, equation (3.9) and the boundary conditions (3.10) hold classically on $\mbox{supp}(\zeta)\cap\Omega$ (for the equation, the argument is given from (3.5) to (3.7)). Then, multiplying (3.9) by $\zeta$, integrating over $\Omega$, and integrating by parts in the first term with the use of (3.10), we obtain (3.11) for $\zeta$ satisfying (3.38).
4.4. Combining the two previous cases, we obtain (3.11) for all smooth $\zeta$ satisfying
$\zeta\equiv 0\qquad\mbox{in $\Omega\cap B_{r}(\mathcal{P}_{0})\,\,$ for some $r>0$.}$ (3.39)
This in particular implies the following: if $\mathcal{P}_{0}\notin\partial\Omega$, then (3.11) holds for all smooth $\zeta$, since we can modify $\zeta$ outside $\Omega$ in this case so that the modified function $\zeta$ satisfies (3.39) for some $r>0$, and this modification clearly does not affect (3.11) for $\zeta$.
Thus, the remaining proof is for the case that $\mathcal{P}_{0}\in\partial\Omega$. Note that this means that $\mathcal{P}_{0}$ is a common endpoint of some of $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ and $\Gamma^{\mbox{\rm\footnotesize ext}}_{j}$.
4.5. Now we consider the case that $\mathcal{P}_{0}$ is a common endpoint of some of $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ and $\Gamma^{\mbox{\rm\footnotesize ext}}_{j}$, and fix $\zeta\in C^{\infty}(\mathbb{R}^{2})$. Let $\psi\in C^{\infty}(\mathbb{R}^{2})$ be such that
$0\leq\psi\leq 1\,\,$ on $\mathbb{R}^{2}$,        $\psi\equiv 0\,\,$ on $B_{1}$,        $\psi\equiv 1$ on $\mathbb{R}^{2}\setminus B_{2}$.
Let $\psi^{r}({\boldsymbol{\xi}})=\psi(\frac{{\boldsymbol{\xi}}}{r})$ for $r>0$. In particular,
$\psi^{r}\equiv 0\;\;\mbox{in $B_{r}$},\qquad\psi^{r}\equiv 1\;\;\mbox{in $\mathbb{R}^{2}\setminus B_{2r}$},\qquad|D\psi^{r}|\leq\frac{C}{r},\qquad\mbox{supp}(D\psi^{r})\subset B_{2r}.$ (3.40)
Here and below, the universal constant $C$ is independent of $r$, which may be different at different occurrence. Then, for any small $r>0$, function $\zeta\psi^{r}$ satisfies (3.39), so (3.11) holds with the test function $\zeta\psi^{r}$ instead of $\zeta$. Thus, we have
$\int_{\Omega}\big{(}\rho f(X)({\mathbf{v}}\cdot\nabla\zeta)+\rho g(X)\zeta\big{)}\psi^{r}\,{\rm d}{\boldsymbol{\xi}}+\int_{\Omega}\rho f(X)({\mathbf{v}}\cdot\nabla\psi^{r})\zeta\,{\rm d}{\boldsymbol{\xi}}$
$\quad-\sum_{i=1}^{N_{2}}\int_{\Gamma^{\mbox{\rm\footnotesize int}}_{i}}\rho f(X)({\mathbf{v}}\cdot{\boldsymbol{\nu}})\zeta\psi^{r}\,{\rm d}l=0.$ (3.41)
We estimate the second integral in (3), by using (3.3), (3.40), the boundedness of ${\mathbf{v}}$ by assumption (ii), and that $f^{\prime}(X)$ is bounded on $\mathbb{R}$:
$\bigg{|}\int_{\Omega}\rho f(X)({\mathbf{v}}\cdot\nabla\psi^{r})\zeta\,{\rm d}{\boldsymbol{\xi}}\bigg{|}\leq\frac{C}{r}\int_{\Omega\cap B_{2r}}(1+|X|)\,{\rm d}{\boldsymbol{\xi}}\leq C\bigg{(}\int_{\Omega\cap B_{2r}}(1+|X|^{2})\,{\rm d}{\boldsymbol{\xi}}\bigg{)}^{\frac{1}{2}}\to 0\quad\,\,\mbox{as }r\to 0.$
Notice that $\psi^{r}\equiv 1$ in $\mathbb{R}^{2}\setminus B_{2r}$ and $|\psi^{r}|\leq 1$ in $B_{2r}$, $(\rho,{\mathbf{v}})$ are bounded and $X\in L^{2}(\Omega)$, and $|(f(X),g(X))|\leq C(1+|X|)$ from the assumptions of $f(X)$ and (3.8). Then, denoting by $L_{r}$ the difference between the first term in (3.11) and the first term in (3), we have
$|L_{r}|=\bigg{|}\int_{\Omega\cap B_{2r}}\big{(}\rho f(X)({\mathbf{v}}\cdot\nabla\zeta)+\rho g(X)\zeta\big{)}(1-\psi_{r})\,{\rm d}{\boldsymbol{\xi}}\bigg{|}\leq C\int_{\Omega\cap B_{2r}}(1+|X|)\,{\rm d}{\boldsymbol{\xi}}\to 0\quad\,\,\mbox{as }r\to 0.$
For the boundary integral, we obtain that, for each $i\in\{1,\cdots,{N_{2}}\}$,
$\int_{\Gamma^{\mbox{\rm\footnotesize int}}_{i}}\rho f(X)({\mathbf{v}}\cdot{\boldsymbol{\nu}})\zeta\psi^{r}\,{\rm d}l\rightarrow\int_{\Gamma^{\mbox{\rm\footnotesize int}}_{i}}\rho f(X)({\mathbf{v}}\cdot{\boldsymbol{\nu}})\zeta\,{\rm d}l\qquad\,\mbox{as }r\to 0.$
Indeed, $\mathcal{P}_{0}\notin\overline{\Gamma^{\mbox{\rm\footnotesize int}}_{i}}$ by Definition 3.1(v), so that $\psi^{r}\equiv 1$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ if $r<\frac{1}{2}\mbox{\rm dist}(\mathcal{P}_{0},\Gamma^{\mbox{\rm\footnotesize int}}_{i})$.
Combining the convergence facts shown above and sending $r\to 0$ in (3), we conclude (3.11).
5. We now show that vorticity $\omega$ (and thus $X$) is smooth and not identically zero on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$. Recall that $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ denotes a relatively open curve segment and $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is a shock by Definition 3.1(vii).
Lemma 3.5.
Under the assumptions of Theorem 3.2, vorticity $\omega$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ from the $\Omega$–side is continuous and not identically zero on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, and is bounded on $\overline{\Gamma^{\mbox{\rm\footnotesize int}}_{1}}$.
This can be proved as follows: First, it follows directly from the regularity of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ in Definition 3.1(v) and assumption (i) of Theorem 3.2 that $\omega$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ from the $\Omega$–side is continuous on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ and bounded on $\overline{\Gamma^{\mbox{\rm\footnotesize int}}_{1}}$. Then, in the rest of the proof, its suffices to show that $\omega$ is not identically zero on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$.
Since equations (2.3)–(2.4) and the Rankine-Hugoniot conditions (2.9) are invariant under the coordinate rotation and translation, at any fixed point $P\in\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, we can choose the coordinates ${\boldsymbol{\xi}}$ such that the $\xi_{1}$–direction is tangent to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at $P$. Then $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is the graph of a function $f_{s}$ locally, i.e., $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\{{\boldsymbol{\xi}}\,:\,\xi_{2}=f_{s}(\xi_{1})\}$ locally near $P$. Thus, $f_{s}$ is in $C^{2}$ in a neighborhood of $P$ by Definition 3.1(v), and we can obtain that $f^{\prime}_{s}=0$ at point $P$ by choosing the appropriate coordinate system. Note that $(f^{\prime}_{s},-1)$ is a normal of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, so that the Rankine-Hugoniot conditions (2.9) become
$\begin{cases}(\rho v_{1}-\rho_{1}v_{1}^{-})f_{s}^{\prime}-(\rho v_{2}-\rho_{1}v_{2}^{-})=0,\\[2.84526pt] (\rho v_{1}^{2}+p-\rho_{1}(v_{1}^{-})^{2}-p_{1})f^{\prime}_{s}-(\rho v_{1}v_{2}-\rho_{1}v_{1}^{-}v_{2}^{-})=0,\\[2.84526pt] (\rho v_{1}v_{2}-\rho_{1}v_{1}^{-}v_{2}^{-})f_{s}^{\prime}-(\rho v_{2}^{2}+p-\rho_{1}(v_{2}^{-})^{2}-p_{1})=0,\end{cases}$ (3.42)
where ${\mathbf{v}}^{-}=(v_{1}^{-},v_{2}^{-})$ and ${\mathbf{v}}=(v_{1},v_{2})$. Taking the tangential derivative $\partial_{\boldsymbol{\tau}}:=\partial_{\xi_{1}}+f_{s}^{\prime}\partial_{\xi_{2}}$ of the Rankine-Hugoniot conditions (3.42) along $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ and using the condition that $f_{s}^{\prime}=0$ at $P$, we obtain that, at point $P$,
$\begin{cases}(\rho v_{1}-\rho_{1}v_{1}^{-})f_{s}^{\prime\prime}-(\rho v_{2})_{\xi_{1}}=0,\\[2.84526pt] (\rho v_{1}^{2}+p-\rho_{1}(v_{1}^{-})^{2}-p_{1})f^{\prime\prime}_{s}-(\rho v_{1}v_{2})_{\xi_{1}}-\rho_{1}v_{2}^{-}=0,\\[2.84526pt] (\rho v_{2}^{2})_{\xi_{1}}+c^{2}\rho_{\xi_{1}}=0.\end{cases}$
By equations (2.3)–(2.4) and the definition: $\omega=(v_{1})_{\xi_{2}}-(v_{2})_{\xi_{1}}$, it follows from a straightforward but long calculation that
$(v_{1})_{\xi_{1}}=$ $-1-\frac{c^{2}+v_{2}^{2}}{\rho|{\mathbf{v}}|^{2}}v_{1}\rho_{\xi_{1}}+\frac{c^{2}-v_{2}^{2}}{\rho|{\mathbf{v}}|^{2}}v_{2}\rho_{\xi_{2}}-\frac{v_{1}v_{2}}{|{\mathbf{v}}|^{2}}\omega,$
$(v_{1})_{\xi_{2}}=$ $-\frac{c^{2}-v_{1}^{2}}{\rho|{\mathbf{v}}|^{2}}v_{2}\rho_{\xi_{1}}-\frac{c^{2}-v_{2}^{2}}{\rho|{\mathbf{v}}|^{2}}v_{1}\rho_{\xi_{2}}+\frac{v_{1}^{2}}{|{\mathbf{v}}|^{2}}\omega,$
$(v_{2})_{\xi_{1}}=$ $-\frac{c^{2}-v_{1}^{2}}{\rho|{\mathbf{v}}|^{2}}v_{2}\rho_{\xi_{1}}-\frac{c^{2}-v_{2}^{2}}{\rho|{\mathbf{v}}|^{2}}v_{1}\rho_{\xi_{2}}-\frac{v_{2}^{2}}{|{\mathbf{v}}|^{2}}\omega,$
$(v_{2})_{\xi_{2}}=$ $-1+\frac{c^{2}-v_{1}^{2}}{\rho|{\mathbf{v}}|^{2}}v_{1}\rho_{\xi_{1}}-\frac{c^{2}+v_{1}^{2}}{\rho|{\mathbf{v}}|^{2}}v_{2}\rho_{\xi_{2}}+\frac{v_{1}v_{2}}{|{\mathbf{v}}|^{2}}\omega.$
Therefore, by a long computation, we obtain that $(\rho_{\xi_{1}},\rho_{\xi_{2}},\omega)$ satisfy the linear system at $P$:
$\begin{pmatrix}a_{1}&b_{1}&c_{1}\\ a_{2}&b_{2}&c_{2}\\ a_{3}&b_{3}&c_{3}\end{pmatrix}\begin{pmatrix}\rho_{\xi_{1}}\\ \rho_{\xi_{2}}\\ \omega\end{pmatrix}=\begin{pmatrix}d_{1}f^{\prime\prime}_{s}\\ d_{2}f^{\prime\prime}_{s}\\ 0\end{pmatrix},$ (3.43)
where
$(a_{1},b_{1},c_{1})=((c^{2}-2v_{1}^{2}-v_{2}^{2})v_{2},\,(c^{2}-v_{2}^{2})v_{1},\,\rho v_{2}^{2}),$
[…]
$(a_{3},b_{3},c_{3})=(3v_{1}^{2}v_{2}^{2}+v_{1}^{2}c^{2}+v_{2}^{4}-v_{2}^{2}c^{2},\,-2(c^{2}-v_{2}^{2})v_{1}v_{2},\,-2\rho v_{2}^{3}),$
and
$d_{1}=-|{\mathbf{v}}|^{2}\big{(}\rho v_{1}-\rho_{1}v_{1}^{-}\big{)},\qquad d_{2}=-|{\mathbf{v}}|^{2}\big{(}\rho v_{1}^{2}+p-\rho_{1}(v_{1}^{-})^{2}-p_{1}\big{)}.$
Notice that
$\begin{vmatrix}a_{1}&b_{1}&c_{1}\\ a_{2}&b_{2}&c_{2}\\ a_{3}&b_{3}&c_{3}\end{vmatrix}=\rho v_{2}(c^{2}-v_{2}^{2})^{2}|{\mathbf{v}}|^{4}\neq 0,$
where we have used the fact that $v_{2}={\mathbf{v}}\cdot\boldsymbol{\nu}\neq 0$ at $P$ by the entropy condition and the ellipticity assumption that $|{\mathbf{v}}|<c$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ in Theorem 3.2. Therefore, from (3.43),
$\omega=\frac{d_{1}\begin{vmatrix}a_{2}&b_{2}\\ a_{3}&b_{3}\end{vmatrix}-d_{2}\begin{vmatrix}a_{1}&b_{1}\\ a_{3}&b_{3}\end{vmatrix}}{\rho v_{2}(c^{2}-v_{2}^{2})^{2}|{\mathbf{v}}|^{4}}f_{s}^{\prime\prime}.$
Since $f^{\prime}_{s}=0$ at $P$, it follows from (3.42) that
$\rho v_{2}-\rho_{1}v_{2}^{-}=0,\,\,\,\,v_{1}=v_{1}^{-},\,\,\,\,\rho v_{2}^{2}+p-\rho_{1}(v_{2}^{-})^{2}-p_{1}=0\qquad\,\,\mbox{at }P.$ (3.44)
Then
$\omega=\frac{v_{1}\big{(}(\rho-\rho_{1})v_{2}^{2}+(p-p_{1})\big{)}}{\rho v_{2}}f_{s}^{\prime\prime}.$ (3.45)
Notice that ${\mathbf{v}}\cdot{\boldsymbol{\nu}}\leq-C^{-1}$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ by Definition 3.1(vii), where ${\mathbf{v}}$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is taken from the $\Omega$–side, and ${\boldsymbol{\nu}}$ is the outer normal with respect to $\Omega$. By (2.10), this implies that ${\mathbf{v}}\cdot{\boldsymbol{\nu}}<0$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ for ${\mathbf{v}}$ taken from the $\Lambda\setminus\Omega$–side. Using the entropy condition (2.12) on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, we obtain that $\rho>\rho_{1}$ and $p>p_{1}$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$. Moreover, $v_{1}(\hat{P})=({\mathbf{v}}\cdot{\boldsymbol{\tau}})(\hat{P})\neq 0$ for point $\hat{P}\in\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ specified in Definition 3.1(viii). Therefore, $\omega\neq 0$ at $\hat{P}$. Then the vorticity is not identically zero on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$.
6. Using Lemma 3.4, we can formally choose $f(s)=s^{2}$ with $g(s)=0$ by (3.8), and use $\zeta\equiv 1$ to obtain that, by (3.11).
$00=-\int_{\Gamma_{\rm shock}}\big{|}\frac{\omega}{\rho}\big{|}^{2}\rho{\mathbf{v}}\cdot{\boldsymbol{\nu}}\,{\rm d}l>$ 0 .
The strict inequality above follows from Definition 3.1(vii) and Lemma 3.5. This (formally) shows that assumption (3.2) leads to a contradiction. A minor technical point is that function $f(s)=s^{2}$ does not satisfy the assumption of Lemma 3.4: $\|f^{\prime}\|_{C^{0,1}(\mathbb{R})}<\infty$.
To make this rigorously, we approximate $f(s)=s^{2}$ by the functions that satisfy the assumptions of Lemma 3.4 and verify the limit process of this approximation. More specifically, for any $M>1$, define $f_{M}\in C^{1}(\mathbb{R})$ by $f_{M}(0)=0$ and $f^{\prime}_{M}(t)=2\min\{|t|,M\}\mathrm{sign}(t)$. Then
$f_{M}(t)=\begin{cases}t^{2}\qquad&\mbox{if }|t|\leq M,\\ M^{2}+2M(|t|-M)\qquad&\mbox{if }|t|>M.\end{cases}$
It follows that $\|f^{\prime}_{M}\|_{C^{0,1}(\mathbb{R})}<\infty$ and the function defined by (3.8) is
$g_{M}(t)=\begin{cases}0&\mbox{if }|t|\leq M,\\ 2(M^{2}-M|t|)\qquad&\mbox{if }|t|>M.\end{cases}$ (3.46)
Now (3.11) holds with $f_{M}$, $g_{M}$, and any $\zeta\in C^{\infty}(\mathbb{R}^{2})$. Choosing $\zeta\equiv 1$, then we have
$\int_{\Omega}\rho g_{M}(X)\,{\mathrm{d}}{\boldsymbol{\xi}}=\int_{\Gamma^{\mbox{\rm\footnotesize int}}}\rho f_{M}(X)\,({\mathbf{v}}\cdot{\boldsymbol{\nu}})\,{\rm d}l.$ (3.47)
We send $M\to\infty$. Clearly, $g_{M}(X)\to 0$ pointwise in $\Omega$. Also, from (3.46), we obtain
$|g_{M}(t)|\leq 2t^{2}\qquad\mbox{ for all $M>1$ and $t\in\mathbb{R}$}.$
Thus, using that $X\in L^{2}(\Omega)$ and $\rho\in L^{\infty}(\Omega)$, we have
$\int_{\Omega}\rho g_{M}(X)\,{\mathrm{d}}{\boldsymbol{\xi}}\to 0\qquad\mbox{as $M\to\infty$},$
by the dominated convergence theorem. Also, since $X\in L^{\infty}(\Gamma^{\mbox{\rm\footnotesize int}})$ by assumptions (i)–(ii), then it follows from the explicit form of $f_{M}$ that
$\int_{\Gamma^{\mbox{\rm\footnotesize int}}}\rho f_{M}(X)\,({\mathbf{v}}\cdot{\boldsymbol{\nu}})\,{\rm d}l=\int_{\Gamma^{\mbox{\rm\footnotesize int}}}\rho|X|^{2}\,({\mathbf{v}}\cdot{\boldsymbol{\nu}})\,{\rm d}l\qquad\mbox{for all $M\geq\|X\|_{L^{\infty}(\Gamma^{\mbox{\rm\footnotesize int}})}$}.$
Thus, sending $M\to\infty$ in (3.47), we obtain
$\int_{\Gamma^{\mbox{\rm\footnotesize int}}}\rho|X|^{2}\,({\mathbf{v}}\cdot{\boldsymbol{\nu}})\,{\rm d}l=0.$
Since $\rho\in[C_{0}^{-1},\,C_{0}]$ by assumption (ii) and ${\mathbf{v}}\cdot{\boldsymbol{\nu}}\leq 0$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$, for $i=2,\cdots,{N_{2}}$, from Definition 3.1(vii), we obtain
$\int_{\Gamma^{\mbox{\rm\footnotesize int}}_{1}}\rho|X|^{2}\,({\mathbf{v}}\cdot{\boldsymbol{\nu}})\,{\rm d}l\geq 0.$
This is a contradiction since $\rho\in[C_{0}^{-1},\,C_{0}]$, ${\mathbf{v}}\cdot{\boldsymbol{\nu}}\leq-C^{-1}$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ by Definition 3.1(vii), and $\omega$ (and thus $X$) is continuous and not identically zero on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ by Lemma 3.5.
This completes the proof of Theorem 3.2. ∎
4. Applications to Transonic Shock problems
In this section, we employ the general framework of Definition 3.1 and the main theorem, Theorem 3.2, established in §3 to analyze the low regularity of entropy solutions of several transonic shock problems including the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the Riemann problem with four-shock interactions.
First of all, in verifying the conditions of Definition 3.1, condition (viii) is often difficult to be verified directly. The following lemma is useful for that; in fact, it is used in all the applications we describe below.
Lemma 4.1.
Let $(\rho,{\mathbf{v}})$ be a solution of a Riemann problem which satisfies conditions (i)–(vii) of Definition 3.1 and assumption (i) of Theorem 3.2. Assume that
  1. (a)
    $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\partial\Omega\cap\partial\Lambda_{j}$ for some $j\in\{1,\cdots,M\}$, where $\Lambda_{j}$ are defined in Definition 3.1(i).
  2. (b)
    $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is not a segment of straight line. In particular, denoting by $P_{1}$ and $P_{2}$ the endpoints of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, there exists a point $P^{*}\in\overline{\Gamma_{1}^{int}}\setminus\{P_{1}\}$ such that $\boldsymbol{\tau}(P_{1})\neq\pm\boldsymbol{\tau}(P^{*})$, where $\boldsymbol{\tau}(\cdot)$ is a unit tangent vector to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at a point.
  3. (c)
    $({\mathbf{v}}\cdot\boldsymbol{\tau})(P_{1})\neq 0$.
Then condition (viii) of Definition 3.1 is satisfied.
Proof.
By re-indexing sets $\Lambda_{j}$, we can assume that $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\partial\Omega\cap\partial\Lambda_{1}$. Then $(\rho_{1},{\mathbf{v}}_{1})$ is the uniform state in $\Lambda_{1}$, where $\rho_{1}$ is constant and ${\mathbf{v}}_{1}=(u_{1}^{(1)},u_{2}^{(1)})-{\boldsymbol{\xi}}$ with a constant state $(u_{1}^{(1)},u_{2}^{(1)})$. This state is called state $(1)$.
We show the existence of a point $\hat{P}$ in the relative interior of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ such that the curvature of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is non-zero at $\hat{P}$ and $({\mathbf{v}}\cdot\boldsymbol{\tau})(\hat{P})\neq 0$. Denote by $S$ the line tangential to $\overline{\Gamma^{\mbox{\rm\footnotesize int}}_{1}}$ at $P_{1}$.
Denote by $Q$ the intersection point of line $S$ and line $L$ through center $O_{1}=(u_{1}^{(1)},u_{2}^{(1)})$ of state $(1)$ perpendicular to $S$. Note that, for any point $P$, ${\mathbf{v}}_{1}=O_{1}-P$ so that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(Q)=0$ and $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(P)\neq 0$ for all $P\in S\setminus{Q}$, where we recall that ${\mathbf{v}}_{1}$ is the pseudo-velocity of the uniform state in $\Lambda_{1}$ which also defines ${\mathbf{v}}_{1}$ on the whole $\mathbb{R}^{2}$.
Denote by $\hat{Q}$ the point on $S$ such that $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ coincides with $S$ between points $P_{1}$ and $\hat{Q}$, but not on any larger interval extended through $\hat{Q}$. Note that it is possible that $\hat{Q}=P_{1}$, but $\hat{Q}\neq P^{*}$ since the tangential line to $\overline{\Gamma^{\mbox{\rm\footnotesize int}}_{1}}$ at $P^{*}$ is not parallel to $L$ by our condition (b).
If $\hat{Q}\neq Q$, then $({\mathbf{v}}\cdot\boldsymbol{\tau})(\hat{Q})=({\mathbf{v}}_{1}\cdot\boldsymbol{\tau})(\hat{Q})\neq 0$, where ${\mathbf{v}}$ is the velocity on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ from the $\Omega$–side and we have used (2.10). Also, from the definition of $\hat{Q}$, in any neighborhood of $\hat{Q}$, there exists a point $P\in\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ with nonzero curvature. Thus, it follows from the $C^{2}$–regularity of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ in its relative interior (by condition (v) of Definition 3.1) and the continuity of ${\mathbf{v}}$ in $\Omega$ near and up to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ (by assumption (i) of Theorem 3.2) that there exists a point $\hat{P}\in(\Gamma^{\mbox{\rm\footnotesize int}}_{1})^{0}$ with non-zero curvature and $({\mathbf{v}}\cdot\boldsymbol{\tau})(\hat{P})\neq 0$.
Therefore, the remaining possibility is that $\hat{Q}=Q$. Note that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(Q)=0$. Moreover, using condition (c), the Rankine-Hugoniot conditions, and the regularity of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ and of $(\rho,{\mathbf{v}})$ given in condition (v) of Definition 3.1 and assumption (i) of Theorem 3.2, we see that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(P_{1})=({\mathbf{v}}\cdot\boldsymbol{\tau}_{S})(P_{1})\neq 0$. This implies that $Q\neq P_{1}$. Also, since $\hat{Q}\neq P^{*}$ as we discussed above, then $Q\neq P^{*}$ for the present case. Thus, for the present case $Q=\hat{Q}$, it follows that $Q$ is an interior point of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, and the part of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ between $P_{1}$ and $Q$ lies on the straight line $S$. In particular, line $S$ is tangential to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at $Q$. We now shift and rotate the coordinates to have the origin at $Q$ and the coordinates $\xi_{1}$ and $\xi_{2}$ to be along $S$ and $L$, respectively. Then $O_{1}=(0,\tilde{v}_{1})$ for some $\tilde{v}_{1}\in\mathbb{R}$ (in fact, $\tilde{v}_{1}\neq 0$ by condition (v) of Definition 3.1, but this will not be used below). To fix notation, let the $\xi_{1}$–axis along $S$ be oriented so that $P_{1}=(\xi_{P_{1}},0)$ with $\xi_{P_{1}}<0$. Since $S=\{\xi_{2}=0\}$ is tangential to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at $Q=(0,0)$, then curve $\Gamma^{\mbox{\rm\footnotesize int}}_{1}\cap B_{r}(Q)$ is a graph for some $r>0$: There exists $f\in C^{2}(\mathbb{R})$ such that
$\Gamma^{\mbox{\rm\footnotesize int}}_{1}\cap B_{r}(Q)=\big{\{}(\xi_{1},f(\xi_{1}))\;:\;\xi_{1}\in(a,b)\big{\}}\qquad\mbox{for some $a<0$, $b>0$, and $f\equiv 0$ on $(a,0)$},$
where the last assertion holds because $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ lies on $S$ between $P_{1}$ and $Q$. Thus, $f^{\prime}(0)=f^{\prime\prime}(0)=0$ so that $|f^{\prime}(\xi_{1})|\leq O(\varepsilon)\xi_{1}$ for all $\xi_{1}\in(-\varepsilon,\varepsilon)$, where $O(\varepsilon)\to 0$ as $\varepsilon\to 0^{+}$. For any $P=(\xi_{1},f(\xi_{1}))\in\Gamma^{\mbox{\rm\footnotesize int}}_{1}\cap B_{r}(Q)$, we have
${\mathbf{v}}_{1}(P)=O_{1}-P=(-\xi_{1},\,\tilde{v}_{2}-f(\xi_{1})),\qquad\boldsymbol{\tau}(P)=\frac{(1,\,f^{\prime}(\xi_{1}))}{\sqrt{1+(f^{\prime}(\xi_{1}))^{2}}}.$
Then
$({\mathbf{v}}_{1}\cdot\boldsymbol{\tau})(P)=\frac{-\xi_{1}+(\tilde{v}_{2}-f(\xi_{1}))f^{\prime}(\xi_{1})}{\sqrt{1+(f^{\prime}(\xi_{1}))^{2}}}=\frac{-\xi_{1}+O(\varepsilon)\xi_{1}}{\sqrt{1+(f^{\prime}(\xi_{1}))^{2}}}\neq 0\qquad\mbox{for $\xi_{1}\in(0,\varepsilon)$ if $\varepsilon$ is small.}$.
Thus, ${\mathbf{v}}\cdot\boldsymbol{\tau}={\mathbf{v}}_{1}\cdot\boldsymbol{\tau}\neq 0$ at any $P=(\xi_{1},f(\xi_{1}))$ with $\xi_{1}\in(0,\varepsilon)$. Since $\hat{Q}=Q$, i.e., for every $\varepsilon>0$, there exists a point $\xi_{1}\in(0,\varepsilon)$ such that $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ has non-zero curvature at $P=(\xi_{1},f(\xi_{1}))$, it follows that there exists a point $\hat{P}$, at which the tangential velocity and the curvature of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ are nonzero, in the present case. This completes the proof. ∎
4.1. Lower Regularity of the Regular Shock Reflection Solutions for the Isentropic Euler System
The first Riemann problem we address is the regular shock reflection problem for the isentropic Euler system (2.1). When a plane incident shock $S_{0}:=\Gamma_{\rm shock}^{0}$ hits a two-dimensional wedge, a shock reflection-diffraction configuration takes shape. The incident shock $S_{0}$ separates two constant states: state $(0)$ with velocity ${\mathbf{u}}^{(0)}=(0,0)$ and density $\rho_{0}$ ahead of the shock, and state (1) with velocity ${\mathbf{u}}^{(1)}=(u^{(1)}_{1},0)$ and density $\rho_{1}$ behind the shock, where $\rho_{1}>\rho_{0}$, and ${u}^{(1)}_{1}>0$ is determined by $(\rho_{0},\rho_{1},\gamma)$ through the Rankine-Hugoniot conditions on $S_{0}$. The incident shock $S_{0}$ moves in the direction of the $x_{1}$–axis and hits the wedge vertex at the initial time. The slip boundary condition ${\mathbf{u}}\cdot\boldsymbol{\nu}=0$ is prescribed on the wedge boundary, where ${\mathbf{u}}$ is the velocity of gas. Since state (1) does not satisfy the boundary condition, the shock reflection-diffraction configuration occurs at later time, which is self-similar. Depending on the flow parameters and the wedge angle, there may be various patterns of shock reflection-diffraction configurations, including Regular Reflection and Mach Reflection.
The regular reflection problem is a lateral Riemann problem in the region
[…]
where +$\mathbb{R}^{2}_{+}=\mathbb{R}^{2}\cap\{x_{1}>0\}$. We seek functions $(\rho,{\mathbf{u}})(\mathbf{x},t)$ satisfying system (2.1) in $\Lambda$ with the boundary condition ${\mathbf{u}}\cdot{\boldsymbol{\nu}}=0$ on $\partial\Lambda$ and the initial data:
$(\rho,{\mathbf{u}})(\mathbf{x},0)=\begin{cases}(\rho_{0},{\mathbf{u}}^{(0)})\qquad&\mbox{if }\;\mathbf{x}\in\Lambda\cap\{x_{1}<0\},\\ (\rho_{1},{\mathbf{u}}^{(1)})\qquad\,&\mbox{if }\;\mathbf{x}\in\Lambda\cap\{x_{1}>0\}.\end{cases}$
This initial-boundary value problem is invariant under scaling (2.2), so we seek self-similar solutions $(\rho,{\mathbf{v}})=(\rho,{\mathbf{v}})({\boldsymbol{\xi}})$, where the self-similar variables ${\boldsymbol{\xi}}$ and the pseudo-velocity ${\mathbf{v}}={\mathbf{u}}-{\boldsymbol{\xi}}$ are introduced in §2.1.
Figure 4.1. Supersonic regular reflection
Figure 4.2. Subsonic regular reflection
First, consider the problem with an assumption on the symmetry with respect to the $x_{1}$–axis. Then we can consider only the upper half-plane $\{x_{2}>0\}$ and prescribe the slip boundary condition ${\mathbf{u}}\cdot{\boldsymbol{\nu}}=0$ on the symmetry line $\{x_{2}=0\}$, so that there is only one reflection point $P_{0}$ to be considered (see Figs 4.24.2).
The regular shock reflection-diffraction configuration is characterized by the fact that the reflection occurs at point $P_{0}$ of the intersection of the incident shock with the wedge boundary. Figs. 4.24.2 show the structure of regular shock reflection-diffraction configurations in self-similar coordinates.
A necessary condition for the existence of a regular reflection-diffraction configuration is the existence of the constant state $(2)$ and the reflected shock line such that state $(2)$ satisfies both the slip boundary condition on the wedge and the Rankine-Hugoniot conditions (2.9)–(2.10) on the shock with state $(1)$ across the reflected shock line written at $P_{0}$. These conditions lead to a system of algebraic equations for the constant velocity and density of state $(2)$. Moreover, the entropy condition (2.8) becomes an inequality in terms of the parameters of states $(1)$ and $(2)$.
It is well-known (see e.g. [15, Chapter 18] for the full Euler system case; the argument for the isentropic Euler system case is similar) that, given the parameters of states $(0)$ and $(1)$, there exists a detachment angle $\theta_{\rm w}^{\rm d}\in(0,\frac{\pi}{2})$ such that the system of algebraic equations for parameters of state (2) has two solutions for each wedge angle $\theta_{\rm w}\in(\theta_{\rm w}^{\rm d},\frac{\pi}{2})$ such that the entropy condition (2.8) is satisfied for the resulting two-shock configuration. These two solutions become equal when $\theta_{\rm w}=\theta_{\rm w}^{\rm d}$. Thus, two types of two-shock configurations occur at $P_{0}$ for each $\theta_{\rm w}\in(\theta_{\rm w}^{\rm d},\frac{\pi}{2})$. For such $\theta_{\rm w}$, state $(2)$ with the smaller density is called a weak state $(2)$. It is expected that the weak state $(2)$ is physical, while the strong state $(2)$ is not stable as the wedge angle tends to $\frac{\pi}{2}$ (as shown in [14] for the potential flow case). In the case of the potential flow model, the global existence of regular shock reflection solutions for all $\theta_{\rm w}\in(\theta_{\rm w}^{\rm d},\frac{\pi}{2})$ with $(\rho,{\mathbf{u}})$ at $P_{0}$ determined by the weak states $(2)$ has been established in [14, 15]. For the full or isentropic Euler system, the existence of regular reflection solutions is an outstanding open problem. From now on, state $(2)$ always refers to the weak state $(2)$, which is unique for each $\theta_{\rm w}\in(\theta_{\rm w}^{\rm d},\frac{\pi}{2})$.
Furthermore, it follows from the Rankine-Hugoniot conditions (2.9)–(2.10) on the straight shock $S_{1}:=\Gamma_{\rm shock}^{1}$ between states $(1)$ and $(2)$, the slip boundary condition on the wedge for state $(2)$, ${\mathbf{v}}_{1}({\boldsymbol{\xi}})=(u_{1}^{(1)},0)-{\boldsymbol{\xi}}$, and ${\mathbf{v}}_{2}({\boldsymbol{\xi}})={\mathbf{u}}_{2}-{\boldsymbol{\xi}}$ that
The shock line $S_{1}$ between states $(1)$ and $(2)$ is not vertical for all $\theta_{\rm w}\in(\theta_{\rm w}^{\rm d},\frac{\pi}{2})$. (4.1)
Moreover, from the entropy condition (2.12) on $S_{1}$, we have
$\rho_{2}>\rho_{1}.$ (4.2)
Depending on the wedge angle, state $(2)$ can be either supersonic or subsonic at $P_{0}$, i.e., either $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$ or the opposite inequality holds, where $c_{2}=\rho_{2}^{\frac{\gamma-1}{2}}$ is the (constant) speed of sound of state (2). Moreover, for $\theta_{\rm w}$ near $\frac{\pi}{2}$ (resp. for $\theta_{\rm w}$ near $\theta_{\rm w}^{\rm d}$), state $(2)$ is supersonic (resp. subsonic) at $P_{0}$. The type of state $(2)$ at $P_{0}$ for a given wedge angle $\theta_{\rm w}$ determines the type of reflection, supersonic or subsonic, as shown in Fig. 4.2 and Fig. 4.2, respectively.
Definition 4.2.
$(\rho,{\mathbf{v}})\in L^{\infty}(\Lambda)$, with $\rho\in BV_{\rm loc}(\Lambda\cap{\mathcal{N}}_{r}(\partial\Lambda))$ for some $r>0$, is called an entropy solution of the regular shock reflection problem if $(\rho,{\mathbf{v}})$ is an entropy solution of system (2.3)–(2.4) with slip boundary condition (2.13) in the sense of Definition 2.2, which satisfies the asymptotic conditions:
$\lim_{R\to\infty}\|(\rho,{\mathbf{v}})-(\bar{\rho},\bar{{\mathbf{v}}})\|_{0,\Lambda\setminus B_{R}(0)}=0,$
where
$(\bar{\rho},\bar{{\mathbf{v}}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for}\,\,\,\xi_{1}>\xi_{1}^{0},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for}\,\,\,\xi_{1}<\xi_{1}^{0},\end{cases}$
and $\xi_{1}^{0}>0$ is the location of the incident shock $S_{0}$ on the self-similar plane.
Next, we define the points and lines in Figs. 4.24.2. The incident shock $S_{0}$ is line $\{\xi_{1}=\xi_{1}^{0}\}$ with $\xi_{1}^{0}=\frac{\rho_{1}{u}^{(1)}_{1}}{\rho_{1}-\rho_{0}}>0$. The center, $O_{2}={\mathbf{u}}^{(2)}=(u_{1}^{(2)},u_{2}^{(2)})$, of the sonic circle $B_{c_{2}}(O_{2})$ of state $(2)$ lies on the wedge boundary between the reflection point $P_{0}$ and the wedge vertex $P_{3}$ for both the supersonic and subsonic cases.
Then, for the supersonic case, i.e., when $|D\varphi_{2}(P_{0})|=|P_{0}O_{2}|>c_{2}$ so that $P_{0}\notin\overline{B_{c_{2}}(O_{2})}$, we denote by $P_{4}$ the upper point of intersection of $\partial B_{c_{2}}(O_{2})$ with the wedge boundary such that $O_{2}\in P_{3}P_{4}$. Also, the sonic circle $\partial B_{c_{2}}(O_{2})$ of state $(2)$ intersects line $S_{1}$, and one of the points of intersection, $P_{1}\in\Lambda$, is such that segment $P_{0}P_{1}$ is outside $B_{c_{2}}(O_{2})$. Denote the arc of $\partial B_{c_{2}}(O_{2})$ by $\Gamma_{\text{\rm sonic}}=P_{1}P_{4}$. The curved part of the reflected-diffracted shock is $\Gamma_{\text{\rm shock}}=P_{1}P_{2}$, where $P_{2}\in\{\xi_{2}=0\}$. Then we denote the line segments $\Gamma_{\rm sym}:=P_{2}P_{3}$ and $\Gamma_{\text{\rm wedge}}:=P_{3}P_{4}$. The lines and curves $\Gamma_{\text{\rm shock}}$, $\Gamma_{\text{\rm sonic}}$, $\Gamma_{\rm sym}$, and $\Gamma_{\text{\rm wedge}}$ do not have common points, except their endpoints $P_{1},\cdots,P_{4}$. Thus, $\Gamma_{\text{\rm shock}}\cup\Gamma_{\text{\rm sonic}}\cup\Gamma_{\rm sym}\cup\Gamma_{\text{\rm wedge}}$ is a closed curve without self-intersection. Denote by $\Omega$ the bounded domain restricted by this curve.
For the subsonic/sonic case, i.e., when $|D\varphi_{2}(P_{0})|=|P_{0}O_{2}|\leq c_{2}$ so that $P_{0}\in\overline{B_{c_{2}}(O_{2})}$, the curved reflected-diffracted shock is $\Gamma_{\text{\rm shock}}=P_{0}P_{2}$, which does not have common interior points with the line segments $\Gamma_{\rm sym}=P_{2}P_{3}$ and $\Gamma_{\text{\rm wedge}}=P_{0}P_{3}$. Then $\Gamma_{\text{\rm shock}}\cup\Gamma_{\rm sym}\cup\Gamma_{\text{\rm wedge}}$ is a closed curve without self-intersection, and $\Omega$ is the bounded domain restricted by this curve.
Furthermore, in some parts of the argument below, it is convenient to extend problem (2.3)–(2.4) and (2.13), given in $\Lambda$ by even reflection about the $\xi_{1}$–axis, i.e., defining
$(\rho^{\rm ext},{\mathbf{v}}^{\rm ext})(-\xi_{1},\xi_{2}):=(\rho^{\rm ext},{\mathbf{v}}^{\rm ext})(\xi_{1},\xi_{2})\qquad\,\,\,\mbox{for any ${\boldsymbol{\xi}}=(\xi_{1},\xi_{2})\in\Lambda$}.$
Then $(\rho^{\rm ext},{\mathbf{v}}^{\rm ext})$ is defined in region $\Lambda^{\rm ext}$ obtained from $\Lambda$ by adding the reflected region $\Lambda^{-}$, i.e., $\Lambda^{\rm ext}=\Lambda\cup\{(\xi_{1},0)\;:\;\xi_{1}<0\}\cup\Lambda^{-}$. In a similar way, region $\Omega$ and curves $\Gamma_{\text{\rm shock}}\subset\partial\Omega$ and $P_{0}P_{2}$ can be extended into the corresponding region $\Omega^{\rm ext}$ and curves $\Gamma_{\text{\rm shock}}^{\rm ext}\subset\partial\Omega^{\rm ext}$ and $P_{0}P_{2}P_{0}^{\rm ext}$.
Now we give the definition of a global regular shock reflection solution. The intuition for the definition is the following: The regular shock reflection solution is an entropy solution of the regular shock reflection problem in the sense of Definition 4.2 which has the structure shown in Figs. 4.24.2, where $(\rho,{\mathbf{v}})$ coincides with those of states $(0)$, $(1)$, and $(2)$ in their respective regions. As we discussed above, the necessary condition is the existence of state $(2)$, which means that the wedge angle satisfies $\theta_{\rm w}\in(\theta_{\rm w}^{\rm d},\frac{\pi}{2})$. Moreover, it is expected that the solution is relatively regular in $\Omega$. However, we show below that it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$. On the other hand, from the physical/computational experiments and the theoretical results in the case of potential flow, it is expected that $\Gamma_{\rm shock}$ is a smooth curve and $(\rho,{\mathbf{v}})$ is smooth near and up to $\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}$ in $\Omega$. Moreover, the regularity discussed is expected for the shock reflection-diffraction configuration extended to $\{\xi_{1}<0\}$ by the even reflection about the $\xi_{1}$–axis (since this is the original shock reflection-diffraction configuration). In particular, the extended shock curve $\Gamma_{\text{\rm shock}}^{\rm ext}$ is smooth, which shows that $\Gamma_{\rm shock}$ must be orthogonal to $\Gamma_{\rm sym}$ at $P_{2}$. Then, noting that the pseudo-velocity of state $(1)$ is ${\mathbf{v}}_{1}({\boldsymbol{\xi}})=(u_{1}^{(1)},0)-{\boldsymbol{\xi}}$ with $u_{1}^{(1)}>0$ and $\{P_{2}\}=\overline{\Gamma}_{\rm shock}\cap\overline{\Gamma}_{\rm sym}\subset\{\xi_{1}<0\}$, it follows that vector ${\mathbf{v}}_{1}$ on $\Gamma_{\rm shock}$ near $P_{2}$ points into $\Omega$. It is then expected that this holds on the whole shock $\Gamma_{\rm shock}$, unless the jumps of the velocity and the density across the reflected-diffracted shock are degenerate at some points (which are not expected). Then it follows from (2.10) for $\Gamma_{\rm shock}$ that ${\mathbf{v}}$ on $\Gamma_{\rm shock}$ from the $\Omega$–side also points into $\Omega$. On the sonic arc (for the supersonic reflection), the jump of the velocity is not expected, and velocity ${\mathbf{v}}_{2}$ of state $(2)$ on $\Gamma_{\rm sonic}$ points into $\Omega$ (as ${\mathbf{v}}_{2}$ points along the radial direction of the sonic circle of state $(2)$ towards its center $O_{2}\in P_{3}P_{4}$). Therefore, it is expected that
${\mathbf{v}}\cdot{\boldsymbol{\nu}}\leq-C^{-1}\qquad\mbox{ on $\overline{\Gamma}_{\rm shock}\cup\overline{\Gamma}_{\rm sonic}$}$ (4.3)
for some $C>0$, where ${\mathbf{v}}$ on the curves is taken from the $\Omega$–side and ${\boldsymbol{\nu}}$ is the outer normal with respect to $\Omega$.
Based on the remarks above, we define the notion of regular shock reflection solutions:
Definition 4.3.
Fix the wedge angle $\theta_{\rm w}\in(\theta_{\rm w}^{\rm d},\frac{\pi}{2})$, and let domain $\Lambda=\Lambda(\theta_{\rm w})$ as defined above. An entropy solution $(\rho,{\mathbf{v}})$ of the regular shock reflection problem in the sense of Definition 4.2 is called a regular shock reflection solution if $(\rho,{\mathbf{v}})$ satisfies the following additional properties:
If state $(2)$ for $\theta_{\rm w}$ is supersonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$, the solution has the supersonic reflection structure as on Fig. 4.2. If state $(2)$ for $\theta_{\rm w}$ is subsonic or sonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|\leq c_{2}$, the solution has the subsonic reflection structure as on Fig. 4.2. More specifically,
  1. (i)
    The extended reflected-diffracted shock curve $P_{0}P_{2}P_{0}^{\rm ext}$ is $C^{1}$ up to its endpoints.
  2. (ii)
    $(\rho,{\mathbf{v}})$ is continuous in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})$ for the supersonic or sonic reflections, and in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock})$ for the subsonic reflection, for some $r>0$.
  3. (iii)
    The solution coincides with states $(0)$, $(1)$, and $(2)$ in their respective regions: for the supersonic reflection case,
    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for $\xi_{1}>\xi_{1}^{0}$ and $\xi_{2}>\xi_{1}\tan\theta_{\rm w}$},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for $\xi_{1}<\xi_{1}^{0}$ and above curve $P_{0}P_{1}P_{2}$},\\[2.84526pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in $P_{0}P_{1}P_{4}$},\end{cases}$
    where $\xi_{1}^{0}>0$ is the location of the incident shock $S_{0}$ on the self-similar plane; and for the subsonic or sonic reflection case,
    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for $\xi_{1}>\xi_{1}^{0}$ and $\xi_{2}>\xi_{1}\tan\theta_{\rm w}$},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for $\xi_{1}<\xi_{1}^{0}$ and above curve $P_{0}P_{2}$},\end{cases}$
    and $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{0}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{2},{\mathbf{v}}_{2})(P_{0})$.
  4. (iv)
    (4.3) holds for some $C>0$, where ${\mathbf{v}}$ on the curves is taken from the $\Omega$–side and ${\boldsymbol{\nu}}$ is the outer normal with respect to $\Omega$.
  5. (v)
    The flow is pseudo-subsonic in $\Omega$ on and near $\Gamma_{\rm shock}$, except the except the sonic point $P_{1}$ for the supersonic or sonic reflections, i.e., $|{\mathbf{v}}|<c$ on $\overline{\Gamma}_{\rm shock}\setminus\{P_{1}\}$ for supersonic and sonic reflections, and on $\overline{\Gamma}_{\rm shock}$ for strictly subsonic reflections, where ${\mathbf{v}}$ on $\Gamma_{\rm shock}$ is computed from the $\Omega$–side.
Remark 4.4.
The curve, $\Gamma_{\rm shock}$, cannot be a straight segment. Indeed, if the shock is a straight segment, then it lies on a vertical line passing through $P_{2}$, since the tangent line to $\Gamma_{\rm shock}$ at $P_{2}$ is vertical by condition (i) of Definition 4.3. On the other hand, the tangent to $\Gamma_{\rm shock}$ at $P_{1}$ for the supersonic reflection and at $P_{0}$ for the subsonic reflection is tangent to the straight shock $S_{1}$ between state $(1)$ and $(2)$, where, for the subsonic case, this follows from the property: $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{0}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{2},{\mathbf{v}}_{2})(P_{0})$ in Definition 4.3(iii) by using the Rankine-Hugoniot conditions (2.9)–(2.10) on the shock. As we have shown above, the straight shock between states $(1)$ and $(2)$ is not vertical. This shows that $\Gamma_{\rm shock}$ cannot lie in a straight line.
Remark 4.5.
For the supersonic reflections, it follows from conditions (ii)–(iii) of Definition 4.3 and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ are continuous across the sonic arc $P_{1}P_{4}=\overline{\Gamma_{\rm sonic}}$:
$\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{2},{\mathbf{v}}_{2}(P))\qquad\mbox{for any $P\in\overline{\Gamma_{\rm sonic}}$}.$ (4.4)
Theorem 4.6.
Let $(\rho,{\mathbf{v}})$ be a regular shock reflection solution in the sense of Definition 4.3. Assume that $(\rho,{\mathbf{v}})$ satisfy the following:
  1. (i)
    The reflected-diffracted shock $\Gamma_{\rm shock}=P_{1}P_{2}$ is $C^{2}$ in its relative interior, and $C^{1}$ up to endpoints $P_{1}$, where $P_{1}$ is replaced by $P_{0}$ if the subsonic shock reflection occurs at $P_{0}$;
  2. (ii)
    $(\rho,{\mathbf{v}})\in C^{1}\left(({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega})\setminus\{P_{1}\}\right)\cap C^{0,1}({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega})$ for some $\sigma>0$, where $P_{1}$ is replaced by $P_{0}$ if the subsonic shock reflection occurs at $P_{0}$;
  3. (iii)
    $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
The proof of Theorem 4.6 will be given for the more general non-symmetric case below (Theorem 4.10), based on the observation in Remark 4.8.
Next, for the non-symmetric regular shock reflection problem, based on the argument near the reflection point $P_{0}$ for the symmetric case above, there are four configurations depending on whether the solution at $P_{0}$ and $P_{1}$ is subsonic or supersonic. As two examples of them, Fig. 4.4 and Fig. 4.4 exhibit the structures of two supersonic regular reflections and two subsonic regular reflections, respectively. In light of Definition 4.3, we define the regular reflection solution for the non-symmetric case. Let the wedge angle $\theta_{\mathrm{w}}=\theta_{\mathrm{w}}^{1}+\theta_{\mathrm{w}}^{2}$ with $\theta_{\mathrm{w}}^{i}\in(\theta_{\mathrm{w}}^{\mathrm{d}},\frac{\pi}{2})$, where $\theta_{\mathrm{w}}^{i}$ is the angle between $\Gamma_{\mathrm{w}}^{i}$ and the $\xi_{1}$–axis for $i=1,2$. Let
$\Lambda=\mathbb{R}^{2}_{+}\setminus\big{\{}\boldsymbol{\xi}\;:\;\xi_{1}>0,\;-\xi_{1}\tan\theta_{\rm w}^{2}<\xi_{2}<\xi_{1}\tan\theta_{\rm w}^{1}\big{\}}.$
Let $\xi_{1}^{0}>0$ be the location of the incident shock $S_{0}$ on the self-similar plane.
Figure 4.3. Non-symmetric supersonic regular reflection
Figure 4.4. Non-symmetric subsonic regular reflection
We call $(\rho,{\mathbf{v}})\in L^{\infty}(\Lambda)$, with $\rho\in BV_{\rm loc}(\Lambda\cap{\mathcal{N}}_{r}(\partial\Lambda))$ for some $r>0$, is an entropy solution of the non-symmetric regular shock reflection problem if $(\rho,{\mathbf{v}})$ is an entropy solution of system (2.3)–(2.4) with the slip boundary condition (2.13) in the sense of Definition 2.2, which satisfies the asymptotic conditions:
$\lim_{R\to\infty}\|(\rho,{\mathbf{v}})-(\bar{\rho},\bar{{\mathbf{v}}})\|_{0,\Lambda\setminus B_{R}(0)}=0,$
where
$(\bar{\rho},\bar{{\mathbf{v}}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for $\xi_{1}>\xi_{1}^{0}$},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for $\xi_{1}<\xi_{1}^{0}$}.\end{cases}$
Then we can define the non-symmetric regular shock reflection solution.
Definition 4.7.
For given angles $\theta_{\mathrm{w}}^{i}$ with $\theta_{\mathrm{w}}^{i}\in(\theta_{\mathrm{w}}^{\mathrm{d}},\frac{\pi}{2})$ for $i=1,2$, an entropy solution $(\rho,\mathbf{v})$ is called a non-symmetric regular shock reflection solution of the non-symmetric regular shock reflection problem if $(\rho,\mathbf{v})$ satisfies the following additional properties:
  1. (i)
    If state $(2)$ at $P_{0}$ and state $(3)$ at $P_{1}$ are both supersonic, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$ and $|{\mathbf{v}}_{2}(P_{1})|>c_{3}$, the solution has the supersonic reflection structure as in Fig. 4.4 at both points $P_{0}$ and $P_{1}$. If state $(2)$ at $P_{0}$ and state $(3)$ at $P_{1}$ are both subsonic or sonic, i.e., $|{\mathbf{v}}_{2}(P_{0})|\leq c_{2}$ and $|{\mathbf{v}}_{2}(P_{1})|\leq c_{3}$, the solution has the subsonic reflection structure as in Fig. 4.4.
  2. (ii)
    The reflected-diffracted shock curve $P_{0}P_{1}$ is $C^{1}$ up to its endpoints.
  3. (iii)
    $(\rho,{\mathbf{v}})$ is continuous in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}^{2}\cup\Gamma_{\rm sonic}^{3})$ for the supersonic reflections, and in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock})$ for the subsonic reflection, for some $r>0$.
  4. (iv)
    The solution coincides with states $(0)$, $(1)$, $(2)$, and $(3)$ in their respective regions:
    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for}\,\,\,\xi_{1}>\xi_{1}^{0},\,\,\,\xi_{2}>\xi_{1}\tan\theta_{\rm w}^{1},\\[2.84526pt] (\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for}\,\,\,\xi_{1}>\xi_{1}^{0},\,\,\,\xi_{2}<-\xi_{1}\tan\theta_{\rm w}^{2},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for}\,\,\,\xi_{1}<\xi_{1}^{0},\,\,\,\mbox{and the left to curve }\;P_{0}P_{1},\\[2.84526pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in}\,\,\,\Lambda_{2}\,\,\,\mbox{if state $(2)$ at $P_{0}$ is supersonic},\\[2.84526pt] (\rho_{3},{\mathbf{v}}_{3})\qquad\mbox{in}\,\,\,\Lambda_{3}\,\,\,\mbox{if state $(3)$ at $P_{1}$ is supersonic},\end{cases}$
    where $\xi_{1}^{0}>0$ is the location of the incident shock $S_{0}$ on the self-similar plane, and
    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{0}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{2},{\mathbf{v}}_{2})(P_{0})\qquad\mbox{if state $(2)$ at $P_{0}$ is subsonic or sonic},$
    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{1}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{3},{\mathbf{v}}_{3})(P_{1})\qquad\mbox{if state $(3)$ at $P_{1}$ is subsonic or sonic}.$
  5. (v)
    (4.3) holds for some $C>0$, where $\Gamma_{\rm sonic}$ is replaced by $\Gamma_{\rm sonic}^{2}\cup\Gamma_{\rm sonic}^{3}$, ${\mathbf{v}}$ on the curves is taken from the $\Omega$–side, and ${\boldsymbol{\nu}}$ is the outer normal with respect to $\Omega$.
  6. (vi)
    The flow is pseudo-subsonic in $\Omega$ on and near $\Gamma_{\rm shock}$, except for the sonic points $P_{2}$ and $P_{3}$ if they exist, point $P_{0}$ if state $(2)$ is sonic at $P_{0}$, and point $P_{1}$ if state $(3)$ is sonic at $P_{1}$.
Remark 4.8.
Similarly as stated in Remark 4.4, curve $\Gamma_{\rm shock}$ cannot be a straight segment, which will be addressed in more detail in Step 3 of the proof of Theorem 4.10 below. Moreover, if $\theta^{1}_{\mathrm{w}}=\theta^{2}_{\mathrm{w}}$, the symmetric regular shock reflection problem is a special case of the non-symmetric regular shock reflection problem, by defining $(\rho^{\rm ext},{\mathbf{v}}^{\rm ext})(\xi_{1},-\xi_{2})):=(\rho^{\rm ext},{\mathbf{v}}^{\rm ext})(\xi_{1},\xi_{2})$ for any ${\boldsymbol{\xi}}=(\xi_{1},\xi_{2})\in\Lambda$. Therefore, we only prove the low regularity of the non-symmetric regular reflection solution below, since the low regularity for the symmetric case follows directly.
Remark 4.9.
For the supersonic reflections, it follows from conditions (ii)–(iv) of Definition 4.7 and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ are continuous across the sonic arcs $\Gamma_{\rm sonic}^{2}\cup\Gamma_{\rm sonic}^{3}$:
$\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{2},{\mathbf{v}}_{i})(P)\qquad\mbox{for any }\;P\in\overline{\Gamma_{\rm sonic}^{i}}\quad\mbox{if }\;|{\mathbf{v}}_{i}(P_{0})|>c_{i}\;\;\mbox{for }\;i=2,3.$
Theorem 4.10.
Let $(\rho,{\mathbf{v}})$ be a non-symmetric regular shock reflection solution in the sense of Definition 4.7. Assume that $(\rho,{\mathbf{v}})$ satisfy the following:
  1. (i)
    The reflected-diffracted shock $\Gamma_{\rm shock}=P_{2}P_{3}$ is $C^{2}$ in its relative interior, and $C^{1}$ up to endpoints $P_{2}$ and $P_{3}$, where $P_{2}$ (or $P_{3}$) is replaced by $P_{0}$ (or $P_{1}$) if the subsonic shock reflection occurs at $P_{0}$ (or $P_{1}$);
  2. (ii)
    $(\rho,{\mathbf{v}})\in C^{1}\left(({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega})\setminus\{P_{2},\,P_{3}\}\right)\cap C^{0,1}({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega})$ for some $\sigma>0$, where $P_{2}$ (or $P_{3}$) is replaced by $P_{0}$ (or $P_{1}$) if the subsonic shock reflection occurs at $P_{0}$ (or $P_{1}$);
  3. (iii)
    $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
Proof.
It suffices to check that the assumptions of Theorem 3.2 are satisfied. We divide the proof into four steps.
1. Since the cases with one supersonic reflection and one subsonic reflection can be treated similarly, we focus only on the non-symmetric supersonic reflection solution (see Fig. 4.4) and the non-symmetric subsonic reflection solution (see Fig. 4.4) in the proof below.
2. We first show that the non-symmetric regular reflection solutions satisfy conditions (i)–(vii) of Definition 3.1 and assumptions (i)–(iii) of Theorem 3.2.
The properties described in Definition 3.1(i) with $M=5$ for the non-symmetric supersonic reflection solutions and with $M=3$ for the non-symmetric subsonic reflection solutions by Definition 4.7(iv), and with $\Lambda_{j}$ being the region of possible state $(0)$, $(1)$, $(2)$, or $(3)$. Moreover, condition (ii) of Definition 3.1 also holds for the regular shock reflection solutions because states $(0)$, $(1)$, $(2)$, and $(3)$ are four different constant states, specifically $\rho_{0}<\rho_{1}<\rho_{2}$ and $\rho_{0}<\rho_{1}<\rho_{3}$.
The properties in Definition 3.1(iii) hold, i.e., $\partial\Omega$ is a Lipschitz curve, by Definition 4.7(ii) and the fact that the rest of $\partial\Omega$ consists of the straight segments and the arcs of the circles for the supersonic reflection, and the angles of vertices $P_{j},j=1,2,\cdots,6$, for the non-symmetric supersonic reflection and of vertices $P_{0}$, $P_{1}$, and $P_{4}$ for the non-symmetric subsonic reflection are all within $(0,\pi)$.
For the non-symmetric supersonic reflection, ${N_{1}}=2$ and ${N_{2}}=3$, with $\Gamma^{\mbox{\rm\footnotesize ext}}_{1}=\Gamma_{\rm wedge}^{1}=P_{4}P_{5}$, $\Gamma^{\mbox{\rm\footnotesize ext}}_{2}=\Gamma_{\rm wedge}^{2}=P_{4}P_{6}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}=P_{2}P_{3}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm sonic}^{2}=P_{2}P_{5}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{3}=P_{3}P_{6}$.
For the non-symmetric subsonic reflection, ${N_{1}}=2$ and ${N_{2}}=1$, with $\Gamma^{\mbox{\rm\footnotesize ext}}_{1}=\Gamma_{\rm wedge}^{1}=P_{0}P_{4}$, $\Gamma^{\mbox{\rm\footnotesize ext}}_{2}=\Gamma_{\rm wedge}^{2}=P_{1}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}=P_{0}P_{1}$.
In both the non-symmetric supersonic and subsonic reflection cases, all the requirements of Definition 3.1(iv)–(v) hold, by the regularity of $\Gamma_{\rm shock}$ given in Definition 4.7(ii) and the facts that the angles in the corner points of $\Omega$ are within $(0,\pi)$ for the non-symmetric regular reflection except for point $P_{4}$, where condition (3.1) holds with $\theta^{-}=\theta_{\rm w}^{1}$ and $\theta^{+}=2\pi-\theta_{\rm w}^{2}$. Also, point $P_{4}=(0,0)$ for the non-symmetric regular reflection solution is point $\mathcal{P}_{0}=(0,0)$ as described in Definition 3.1(iv). Moreover, $P_{4}$ is the common point of $\Gamma^{\mbox{\rm\footnotesize ext}}_{1}=\Gamma_{\rm wedge}^{1}$ and $\Gamma^{\mbox{\rm\footnotesize ext}}_{2}=\Gamma_{\rm wedge}^{2}$ for both the regular supersonic and subsonic reflections, as described in Definition 3.1(iv).
Now we show that the requirements of Definition 3.1(vii) hold. Clearly, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}$ is a shock, and ${\mathbf{v}}\cdot{\boldsymbol{\nu}}\leq-C^{-1}$ where ${\mathbf{v}}$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is taken from the $\Omega$–side and ${\boldsymbol{\nu}}$ is the outer normal with respect to $\Omega$ by property (v) of Definition 4.7 of the non-symmetric regular reflection solution. Also, for the non-symmetric supersonic reflection, $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm sonic}^{2}$ and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{3}$ are an arc of the sonic circles of state $(2)$ and state $(3)$, respectively, which imply that ${\mathbf{v}}_{2}\cdot{\boldsymbol{\nu}}=-|{\mathbf{v}}_{2}|=-c_{2}<0$ on $\Gamma_{\rm sonic}^{2}$ and ${\mathbf{v}}_{3}\cdot{\boldsymbol{\nu}}=-|{\mathbf{v}}_{3}|=-c_{3}<0$ on $\Gamma_{\rm sonic}^{3}$. Using the boundary conditions that ${\mathbf{v}}={\mathbf{v}}_{2}$ on $\Gamma_{\rm sonic}^{2}$ and ${\mathbf{v}}={\mathbf{v}}_{3}$ on $\Gamma_{\rm sonic}^{3}$ by Definition 4.7(iv), we obtain that
${\mathbf{v}}\cdot{\boldsymbol{\nu}}=-c_{2}<0\qquad\mbox{on $\Gamma_{\rm sonic}^{2}$ and ${\mathbf{v}}\cdot{\boldsymbol{\nu}}=-c_{3}<0$ on $\Gamma_{\rm sonic}^{3}$}.$
Thus, the requirements of Definition 3.1(vii) hold.
Next, since $\overline{\Gamma^{\mbox{\rm\footnotesize int}}}=\overline{\Gamma_{\rm shock}}\cup\overline{\Gamma_{\rm sonic}^{1}}\cup\overline{\Gamma_{\rm sonic}^{2}}$, assumptions (ii)–(iii) of Theorem 4.10 imply that assumptions (i)–(ii) of Theorem 3.2 hold for the regular shock reflection solution.
Therefore, we have shown that conditions (i)–(vii) of Definition 3.1 and assumptions (i)–(ii) of Theorem 3.2 hold.
3. It remains to show that the requirements in Definition 3.1(viii) hold. This is achieved by the use of Lemma 4.1. Thus, it remains to show that conditions (a)–(c) of Lemma 4.1 are satisfied below.
Condition (a) is satisfied since $\Lambda_{1}$ is the region of state $(1)$ as shown in Figs. 4.44.4 above, and $\Gamma_{\rm shock}=\partial\Lambda_{1}\cap\partial\Omega$ in both the supersonic and subsonic cases.
We now check condition (b) of Lemma 4.1. Assume that $\Gamma_{\rm shock}$ is a straight segment. Recall that $S_{1}=\Gamma_{\rm shock}^{1}$ is the straight shock between states $(1)$ and $(2)$, i.e., the line passing through points $P_{0}$ and $P_{2}$ for the supersonic reflection, and through $P_{0}$ for the subsonic reflection. Using Definition 4.7(iv), we obtain that $(\rho,{\mathbf{v}})(P_{2})=(\rho_{2},{\mathbf{v}}_{2})(P_{2})$ for the non-symmetric supersonic reflection, and $(\rho,{\mathbf{v}})(P_{0})=(\rho_{2},{\mathbf{v}}_{2})(P_{0})$ for the non-symmetric subsonic reflection, where $(\rho,{\mathbf{v}})$ in both cases is computed from the $\Omega$–side. Thus, the tangent line to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}$ at the upper endpoint, i.e., at $P_{2}$ for the non-symmetric supersonic reflection and at $P_{0}$ for the non-symmetric subsonic reflection, is line $S_{1}$. Similarly, we have the tangent line to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}$ at the lower endpoint, i.e., at $P_{3}$ for the non-symmetric supersonic reflection and at $P_{1}$ for the non-symmetric subsonic reflection, is line $S_{2}=\Gamma_{\rm shock}^{2}$. Hence, if $\Gamma_{\rm shock}$ is a straight segment, then $\Gamma_{\rm shock}$ lies on both lines $S_{1}$ and $S_{2}$; in particular, these lines coincide. It follows that $S_{1}=S_{2}$ is the line passing through $P_{0}$ and $P_{1}$, and $\Gamma_{\rm shock}$ lies within interval $P_{0}P_{1}$. However, the wedge is convex since $\theta_{\mathrm{w}}^{i}\in(0,\frac{\pi}{2})$ for $i=1,2$, so $P_{0}P_{1}$ lies within the wedge, thus outside $\Lambda$. It follows that $\Gamma_{\rm shock}\subset P_{0}P_{1}$ lies outside $\Lambda$, which contradicts the structure of regular reflection-diffraction configuration. That is, the assumption that $\Gamma_{\rm shock}$ is a straight segment leads to a contradiction, which verifies (b) of Lemma 4.1.
Next, we show condition (c) of Lemma 4.1 is satisfied.
Consider first the case of reflections that are supersonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$; see Fig. 4.4. Then we need to show that $({\mathbf{v}}\cdot{\boldsymbol{\tau}}_{\rm shock})(P_{2})\neq 0$. Assume this is not true, then $({\mathbf{v}}\cdot{\boldsymbol{\tau}}_{\rm shock})(P_{2})=0$. Note that ${\boldsymbol{\tau}}_{\rm shock}(P_{2})={\boldsymbol{\tau}}_{S_{1}}$ and ${\mathbf{v}}_{|\Omega}(P_{2})={\mathbf{v}}_{2}(P_{2})$ by Definition 4.7(ii, iv) so that ${\mathbf{v}}_{2}(P_{2})\cdot{\boldsymbol{\tau}}_{S_{1}}=0$ and
$|{\mathbf{v}}(P_{2})\cdot{\boldsymbol{\nu}}_{\rm shock}|=|{\mathbf{v}}_{2}(P_{2})\cdot{\boldsymbol{\nu}}_{S_{1}}|=|{\mathbf{v}}_{2}(P_{2})|=c_{2},$
which contradicts the last inequality in (2.12). Thus condition (c) of Lemma 4.1 is proved in the case when $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$.
In the case of reflections that are sonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|=c_{2}$ (see Fig. 4.4), the argument is the same as above with only notational change: we use point $P_{0}$ here instead of point $P_{2}$.
It remains to consider the case of reflections that are subsonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|<c_{2}$; see Fig. 4.4. Assume that $({\mathbf{v}}\cdot\boldsymbol{\tau}_{\rm shock})(P_{0})=0$. Recall that $(\rho,{\mathbf{v}})(P_{0})=(\rho_{2},{\mathbf{v}}_{2})(P_{0})$ for the subsonic reflection and $S_{1}$ is tangent to $\Gamma_{\rm shock}$ at $P_{0}$, as we have shown above. Thus, from $({\mathbf{v}}\cdot\boldsymbol{\tau}_{\rm shock})(P_{0})=0$, we obtain that $({\mathbf{v}}_{2}\cdot\boldsymbol{\tau}_{S_{1})}(P_{0})=0$. Since $S_{1}$ is the line shock between states $(1)$ and $(2)$, the last equality implies by (2.10) that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S_{1})}(P_{0})=0$, so that line $L$ through centers $O_{1}=(u_{1}^{(1)},0)$ and $O_{2}=(u_{1}^{(2)},u_{2}^{(2)})$ of states $(1)$ and $(2)$ (which is orthogonal to $S_{1}$) intersects $S_{1}$ at $P_{0}$. Since $P_{0}\in\Gamma_{\rm wedge}$, $O_{2}\in L\cap\Gamma_{\rm wedge}$ (where $O_{2}\in\Gamma_{\rm wedge}$ because ${\mathbf{v}}_{2}\cdot{\boldsymbol{\nu}}=0$ on $\Gamma_{\rm wedge}$), and $L\cap\Gamma_{\rm wedge}=\{P_{0}\}$, it follows that $O_{2}=P_{0}$ so that ${\mathbf{v}}_{2}(P_{0})=\bf{0}$. Since ${\mathbf{v}}_{2}(P_{0})=\bf{0}$, it follows from the Rankine-Hugoniot conditions (2.9)–(2.10) on $S_{1}$ between states $(1)$ and $(2)$ that ${\mathbf{v}}_{1}(P_{0})=\bf{0}$, that is, $O_{1}=P_{0}$. However, this is not true since $O_{1}=(u_{1}^{(1)},0)$ for $u_{1}^{(1)}>0$, while $P_{0}=(l\cos\theta_{\rm w}^{1},l\sin\theta_{\rm w}^{1})$ for some $l>0$. This contradiction shows that $({\mathbf{v}}\cdot\boldsymbol{\tau}_{\rm shock})(P_{0})\neq 0$, i.e., condition (c) of Lemma 4.1 holds for the subsonic reflection at $P_{0}$.
Now all the conditions of Lemma 4.1 are verified for the regular reflection solutions. Applying Lemma 4.1, we obtain that Definition 3.1(viii) holds.
4. Now, all the conditions of Theorem 3.2 are verified for the regular reflection solutions. Then the conclusion of Theorem 4.10 follows from Theorem 3.2. ∎
4.2. Lower Regularity of the Prantl-Meyer Reflection Solutions for Supersonic Flows past a Solid Ramp
The second example is the Prandtl reflection problem for the isentropic Euler system (2.1). This is of a self-similar structure that occurs when a $2$-D supersonic flow with density $\rho_{\infty}>0$ and velocity ${\mathbf{v}}_{\infty}=(u_{\infty},0)$, $u_{\infty}>0$, along the wedge-axis hits the wedge in the direction at $t=0$. See Figs. 4.64.6; also see Bae-Chen-Feldman [2, 3] and Elling-Liu [26].
Figure 4.5. Prandtl supersonic shock reflection
Figure 4.6. Prandtl subsonic shock reflection
Consider the problem in the self-similar coordinates ${\boldsymbol{\xi}}=(\xi_{1},\xi_{2})$ in the region:
$\Lambda=\mathbb{R}^{2}_{+}\setminus\big{\{}{\boldsymbol{\xi}}:\,\xi_{2}>\max(0,\xi_{1}\tan\theta_{\rm w})\big{\}}.$
We seek global entropy solutions of the boundary value problem in the sense of Definition 2.2.
First, a similar argument as made for the regular shock reflection at the reflection point $P_{0}$ (see Figs. 4.24.2) yields that, at wedge-vertex $P_{1}$, for a given uniform incoming flow $(\rho_{\infty},{\mathbf{v}}_{\infty})$, there is a detachment angle $\theta_{\rm w}^{\rm d}\in(0,\frac{\pi}{2})$ such that the system of algebraic equations (2.9) and (2.13) for state $(O)$ has two solutions for each wedge-angle $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$ such that the entropy condition is satisfied for the two-shock configuration. The weak state $(O)$ with the smaller density is expected to be physical. So we always refer to state $(O)$ as the weak state $(O)$. Similarly, it follows from the Rankine-Hugoniot conditions (2.9) and the slip boundary condition (2.13) that
The shock line $\Gamma_{\rm shock}^{O}$ between states $(O)$ and ()($\infty$)( ∞ ) is not parallel to $\Gamma_{\rm wedge}$ for all $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$. (4.5)
Depending on the wedge-angle, state $(O)$ can be either supersonic or subsonic at $P_{1}$. It determines the type of the reflection, supersonic or subsonic, as shown in Fig. 4.6 and Fig. 4.6, respectively.
Second, by a straightforward computation, we know that there exists a unique constant state $(N)$, which determines the normal reflection of state ()($\infty$)( ∞ ) from the wedge boundary $\Gamma_{\rm wedge}$ so that state $(N)$ satisfies the slip boundary condition (2.13) along $\Gamma_{\rm wedge}$ and the Rankine-Hugoniot conditions (2.9) along a straight line $\Gamma_{\rm shock}^{N}$, which lies in $\Lambda$ and is parallel to $\Gamma_{\rm wedge}$.
Hinted by the solution structures given in [2, 3, 26] for the potential flow, for any given wedge-angle $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$, an entropy solution $(\rho,{\mathbf{v}})$ of the boundary value problem in the sense of Definition 2.2 is called a regular Prandtl-Meyer reflection solution for the isentropic Euler system (2.3)–(2.4) if $(\rho,{\mathbf{v}})$ satisfies the following further properties:
  1. (i)
    If state $(O)$ at $P_{1}$ is supersonic, i.e., $|{\mathbf{v}}_{O}(P_{1})|>c_{O}$, the solution has the supersonic reflection structure as in Fig. 4.6 at point $P_{1}$. If state $(O)$ at $P_{1}$ is subsonic or sonic, i.e., $|{\mathbf{v}}_{O}(P_{1})|\leq c_{O}$, the solution has the subsonic reflection structure as in Fig. 4.6.
  2. (ii)
    The reflected shock curve $\Gamma_{\rm shock}$ (i.e., $P_{4}P_{5}$ for the supersonic reflection and $P_{1}P_{4}$ for the subsonic reflection) is $C^{1}$ up to its endpoints and is $C^{2}$ in its relative interior.
  3. (iii)
    $(\rho,{\mathbf{v}})$ is continuous in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}^{N}\cup\Gamma_{\rm sonic}^{O})$ when $|{\mathbf{v}}_{O}(P_{1})|>c_{O}$, and in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}^{N})$ when $|{\mathbf{v}}_{O}(P_{1})|\leq c_{O}$ for some $r>0$.
  4. (iv)
    The solution coincides with states ()($\infty$)( ∞ ), $(O)$, and $(N)$ in their respective regions. Specifically,
    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{\infty},{\mathbf{v}}_{\infty})\qquad\,\,&\mbox{in}\,\,\,\Lambda_{\infty},\\[1.42262pt] (\rho_{N},{\mathbf{v}}_{N})\qquad\,\,&\mbox{in}\,\,\,\Lambda_{N},\\[1.42262pt] (\rho_{O},{\mathbf{v}}_{O})\qquad\,\,&\mbox{in}\,\,\,\Lambda_{O}\,\,\,\,\mbox{if state (O) at $P_{1}$ is supersonic},\end{cases}$
    and
    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{1}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{O},{\mathbf{v}}_{O})(P_{1})\qquad\mbox{if state (O) at $P_{1}$ is subsonic or sonic}.$
  5. (v)
    ${\mathbf{v}}\cdot\nu\leq-C^{-1}$ on $\overline{\Gamma_{\rm shock}}\cup\overline{\Gamma_{\rm sonic}^{N}}\cap\overline{\Gamma_{\rm sonic}^{O}}$ for some $C>0$, where ${\mathbf{v}}$ on the curves is taken from the $\Omega$–side and ${\boldsymbol{\nu}}$ is the outer normal with respect to $\Omega$.
  6. (vi)
    The flow is pseudo-subsonic in $\Omega$ on and near $\Gamma_{\rm shock}$, except for the sonic points $P_{4}$ and $P_{5}$ if they exist, or point $P_{1}$ if state $(O)$ is sonic at $P_{1}$.
We remark that, for the supersonic reflections, it follows from (ii)–(iv) of the definition above and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ are continuous across the sonic arcs $\Gamma_{\rm sonic}^{N}$ and $\Gamma_{\rm sonic}^{O}$.
Then, following the arguments as for the proof of Theorem 4.10, we have the following theorem on the lower regularity of the regular Prandtl-Meyer shock reflection solutions:
Theorem 4.11.
Let $(\rho,{\mathbf{v}})$ be a regular Prandtl-Meyer shock reflection solution for the isentropic Euler system such that
  1. (i)
    $(\rho,{\mathbf{v}})\in C^{1}\big{(}({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}^{N}\cup\Gamma_{\rm sonic}^{O})\cap\overline{\Omega})\setminus\{P_{4},\,P_{5}\}\big{)}\cap C^{0,1}\big{(}{\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}^{N}\cup\Gamma_{\rm sonic}^{O})\cap\overline{\Omega}\big{)}$ for some $\sigma>0$, where $P_{5}$ is replaced by $P_{1}$ if the subsonic shock reflection occurs at $P_{1}$;
  2. (ii)
    $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
Proof.
For the supersonic reflection, $M=3$, ${N_{1}}=1$, and ${N_{2}}=3$ with $\Lambda_{1}=\Lambda_{N}$, $\Lambda_{2}=\Lambda_{\infty}$, $\Lambda_{3}=\Lambda_{O}$, $\Gamma^{\mbox{\rm\footnotesize ext}}_{1}=\Gamma_{\rm wedge}=P_{2}P_{3}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}=P_{4}P_{5}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm sonic}^{N}=P_{3}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{O}=P_{2}P_{5}$.
For the subsonic reflection, $M=2$, ${N_{1}}=1$, and ${N_{2}}=2$ with $\Lambda_{1}=\Lambda_{N}$, $\Lambda_{2}=\Lambda_{\infty}$, $\Gamma^{\mbox{\rm\footnotesize ext}}_{1}=\Gamma_{\rm wedge}=P_{1}P_{3}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}=P_{1}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm sonic}^{N}=P_{3}P_{4}$.
Then we can follow the proof of Theorem 4.10 to show that the regular solutions satisfy the assumptions of Theorem 3.2 similarly. The only (slight) difference is in showing that condition (b) of Lemma 4.1 holds. We need to show that $\Gamma_{\rm shock}$ is not a straight segment. Let us first consider the supersonic Prandtl reflection case; see Fig. 4.6. Arguing as in the proof of Theorem 4.10, we show that, at point $P_{4}$ (resp. $P_{5}$), curve $\Gamma_{\rm shock}$ is tangential to line $\Gamma_{\rm shock}^{N}$ (resp. $\Gamma_{\rm shock}^{O}$). If $\Gamma_{\rm shock}$ is a straight segment, we obtain that $\Gamma_{\rm shock}^{N}$ and $\Gamma_{\rm shock}^{O}$ lie in the same line. However, $\Gamma_{\rm shock}^{N}$ is parallel to $\Gamma_{\rm wedge}$ with a positive distance from it, while $\Gamma_{\rm shock}^{O}$ passes through point $P_{1}\in\Gamma_{\rm wedge}$. This contradiction shows that $\Gamma_{\rm shock}$ is not a straight segment. In the case of subsonic or sonic Prandtl reflection as in Fig. 4.6, we argue similarly, except we consider point $P_{1}$ instead of point $P_{5}$ and use the fact that, in the subsonic or sonic case, $\Gamma_{\rm shock}$ is parallel to $\Gamma_{\rm shock}^{O}$ at $P_{1}$ by the last equality in condition (iv) of the definition of admissible solutions. Then we conclude in the same way as in the case of supersonic Prandtl reflection. Thus, we do not repeat the similar arguments for the proof. ∎
4.3. Lower Regularity of the Shock Diffraction Solutions of the Lighthill Problem
The third problem is the Lighthill problem, i.e., the shock diffraction problem, for the isentropic Euler system. As discussed in [16] for the potential flow and shown in Fig. 4.7, the Lighthill shock diffraction problem arises as a straight incident shock passes through a wedge stepping down.
Figure 4.7. Lighthill shock diffraction problem
Initially, we consider two piecewise constant Riemann data with the left state $(1)$: $(\rho_{1},u_{1}^{(1)},0)$ for $u_{1}^{(1)}>0$ and the right state $(0)$: $(\rho_{0},0,0)$, separated by a vertical shock, which hits the wedge at the wedge-corner $P_{1}$. Then the Lighthill shock diffraction problem evolves in a self-similar structure as time goes on. In the self-similar coordinates ${\boldsymbol{\xi}}$, the incident shock $\Gamma_{\rm shock}^{1}$ is given by $\xi_{1}=\xi_{1}^{0}$. By a straightforward calculation, it follows from the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12) that the location of the incident shock satisfies $0<\xi_{1}^{0}<c_{1}$, where $c_{1}=\rho_{1}^{\frac{\gamma-1}{2}}$ is the sonic speed of state $(1)$. Thus, as shown in Fig. 4.7, the incident shock $\Gamma_{\rm shock}^{1}$ interacts with the sonic circle $\Gamma_{\rm sonic}$ of state $(1)$ and becomes a transonic shock $\Gamma_{\rm shock}$, and then $\Gamma_{\rm shock}$ meets wedge $\Gamma_{\rm wedge}^{0}$ perpendicularly.
Let $\Lambda_{0}$, $\Lambda_{1}$, and $\Omega$ be defined, as shown in Fig. 4.7. Let $\Lambda=\overline{\Lambda_{0}\cup\Lambda_{1}\cup\Omega}$. An entropy solution $(\rho,{\mathbf{v}})$ of the boundary value problem in $\Lambda$ in the sense of Definition 2.2 is called a regular shock diffraction solution of the Lighthill problem for the isentropic Euler system if $(\rho,{\mathbf{v}})$ satisfies the following further properties:
  1. (i)
    The diffracted shock curve $\Gamma_{\rm shock}=P_{2}P_{3}$ is $C^{1}$ up to its endpoints and is $C^{2}$ in its relative interior. Its tangent is perpendicular to $\Gamma_{\rm wedge}^{0}$ at $P_{2}$ and is vertical at $P_{3}$.
  2. (ii)
    $(\rho,{\mathbf{v}})$ is continuous in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})$ for some $r>0$.
  3. (iii)
    The solution coincides with states $(0)$ and $(1)$ in their respective regions. Specifically,
    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{in $\Lambda_{0}$},\\[1.42262pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{in $\Lambda_{1}$},\end{cases}$
    where ${\mathbf{v}}_{0}=-{\boldsymbol{\xi}}$ and ${\mathbf{v}}_{1}=(u_{1}^{(1)},0)-{\boldsymbol{\xi}}$.
  4. (iv)
    ${\mathbf{v}}\cdot\nu\leq-C^{-1}$ on $\overline{\Gamma_{\rm shock}}\cup\overline{\Gamma_{\rm sonic}}$ for some $C>0$, where ${\mathbf{v}}$ on the curves is taken from the $\Omega$–side and ${\boldsymbol{\nu}}$ is the outer normal with respect to $\Omega$.
  5. (v)
    The flow is pseudo-subsonic in $\Omega$ on and near $\Gamma_{\rm shock}$ except for the sonic points $P_{3}$.
Let $M=2$, ${N_{1}}=2$, and ${N_{2}}=2$ with $\Lambda_{1}=\Lambda_{0}$, $\Lambda_{2}=\Lambda_{1}$, $\Gamma^{\mbox{\rm\footnotesize ext}}_{1}=\Gamma_{\rm wedge}^{1}=P_{1}P_{4}$, $\Gamma^{\mbox{\rm\footnotesize ext}}_{2}=\Gamma_{\rm wedge}^{0}=P_{1}P_{2}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}=P_{2}P_{3}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm sonic}=P_{3}P_{4}$. Then, following the proof of Theorem 4.10, we have the following theorem on the lower regularity of the regular shock diffraction solutions.
Theorem 4.12.
Let $(\rho,{\mathbf{v}})$ be a regular shock diffraction solution of the Lighthill problem for the isentropic Euler system and satisfy the following:
  1. (i)
    $\rho,{\mathbf{v}}\in C^{1}\left(({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega})\setminus\{P_{3}\}\right)\cap C^{0,1}({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega})$ for some $\sigma>0$;
  2. (ii)
    $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
Since the proof argument is similar to the one for Theorem 4.11, we omit the details.
4.4. Lower Regularity of the Riemann Solutions with Four-Shock Interactions
The final problem is the Riemann problem with four-shock interaction structure for the isentropic Euler system. As discussed in [12] for the Riemann problem with four-shock interactions for the potential flow, initially in the ${\mathbf{x}}$–coordinates, the scale-invariant domains $\Lambda_{i}\subset\mathbb{R}^{2}$, for $i=1,2,3,4$, are defined by
$\Lambda_{1}=\big{\{}{\mathbf{x}}\in\mathbb{R}^{2}\,:\,-\theta_{14}<\theta<\theta_{12}\big{\}},\qquad\qquad\quad\Lambda_{2}=\big{\{}{\mathbf{x}}\in\mathbb{R}^{2}\,:\,\theta_{12}<\theta<\pi-\theta_{32}\big{\}},$
$\Lambda_{3}=\big{\{}{\mathbf{x}}\in\mathbb{R}^{2}\,:\,\pi-\theta_{32}<\theta<\pi+\theta_{34}\big{\}},\qquad\,\Lambda_{4}=\big{\{}{\mathbf{x}}\in\mathbb{R}^{2}\,:\,\pi+\theta_{34}<\theta<2\pi-\theta_{14}\big{\}},$
where $\theta$ is the polar angle of point ${\mathbf{x}}\in\mathbb{R}^{2}$ and the four parameters $\theta_{12},\theta_{32},\theta_{34},\theta_{14}\in(0,\frac{\pi}{2})$. On each $\Lambda_{i}$, suitable constant states $(i)$ with values $(\rho_{i},{\mathbf{v}}_{i})$ are given for the Riemann initial data such that any two neighboring states are connected by exactly one planar shock discontinuity, which satisfies the Rankine-Hugoniot conditions and the entropy condition for the isentropic Euler system. As introduced in [12], under the symmetry assumption that $\theta_{12}=\theta_{14}=\theta_{1}$, $\theta_{32}=\theta_{34}=\theta_{2}$, and states in $\Lambda_{2}$ and $\Lambda_{4}$ are the same constants, and under a structure assumption that one forward shock is generated between states $(1)$ and $(j)$ and one backward shock is generated between states $(3)$ and $(j)$, governed by the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12), we expect the Riemann problem develops a Riemann solution with the structure as shown in Fig. 4.8.
Figure 4.8. Riemann problem with four-shock interactions
On the symmetry line that is the dashed line shown in Fig. 4.8, the velocity satisfies (2.13) with $\boldsymbol{\nu}$ being the vertical direction. At the interaction points $P_{1}$ and $P_{4}$, based on an argument similar to the one for the regular shock reflection problem at the reflection point $P_{0}$, with the symmetry line and state $(2)$ in Fig. 4.8 corresponding to $\Gamma_{\rm wedge}$ and state $(1)$ in Fig. 4.2, there exists a detachment angle $\theta^{\rm d}\in(0,\frac{\pi}{2})$ depending on the data such that, for $\theta_{1},\theta_{2}\in(0,\theta_{\rm w}^{\rm d})$, there exist two constant states $(5)$ and $(6)$ in $\Lambda_{5}$ and $\Lambda_{6}$, respectively, satisfying the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12). One corresponds to the strong shock and the other to the weak shock. We always select the weak one because it expects to be stable in general. Then there exists a sonic angle $\theta^{\rm s}\in(0,\theta^{\rm d}]$ such that the state at $P_{1}$ (or $P_{4}$) is pseudo-supersonic if $\theta_{1}\in(0,\theta^{\rm s})$ (or $\theta_{2}\in(0,\theta^{\rm s})$) and is pseudo-subsonic if $\theta_{1}\in(\theta^{\rm s},\theta^{\rm d}]$ (or $\theta_{2}\in(\theta^{\rm s},\theta^{\rm d}]$).
When $\theta_{1}\in(0,\theta^{\rm s})$ at $P_{1}$, there exist points $P_{2}$ and $P_{6}$ on $\Gamma_{\rm sonic}^{1}$ (the sonic curve of state $(6)$) such that shocks $P_{1}P_{2}$ and $P_{1}P_{6}$ are straight shocks and the state in $\Lambda_{6}$ with boundaries $P_{1}P_{2}$, $\Gamma_{\rm sonic}^{1}$, and $P_{1}P_{6}$ is the constant state $(6)$. Similarly, when $\theta_{2}\in(0,\theta^{\rm s})$ at $P_{4}$, there exist points $P_{3}$ and $P_{5}$ on $\Gamma_{\rm sonic}^{2}$ (the sonic curve of state $(5)$) such that shocks $P_{3}P_{4}$ and $P_{4}P_{5}$ are straight shocks and the state in $\Lambda_{5}$ with boundaries $P_{3}P_{4}$, $\Gamma_{\rm sonic}^{2}$, and $P_{4}P_{5}$ is the constant state $(5)$. On the other hand, when $\theta_{1}\in[\theta^{\rm s},\theta^{\rm d}]$ at $P_{1}$ (or when $\theta_{2}\in[\theta^{\rm s},\theta^{\rm d}]$ at $P_{4}$), the sonic arc, the straight shocks, and $\Lambda_{6}$ (or $\Lambda_{5}$) disappear near point $P_{1}$ (or $P_{4}$), and state $(6)$ (or state $(5)$) is the limit of solution $(\rho,{\mathbf{v}})$ in $\Omega$ as the point tends to $P_{1}$ (or $P_{4}$). It is direct to see that
The tangents to Γshock1 at P2 and P3 are not parallel to each $\mbox{\it The tangents to $\Gamma_{\rm shock}^{1}$ at $P_{2}$ and $P_{3}$ are not parallel to each other},$ (4.6)
where $P_{2}$ is replaced by $P_{1}$ if state $(6)$ is subsonic and $P_{3}$ is replaced by $P_{4}$ if state $(5)$ is subsonic.
Let $\Lambda_{i}$ for $i=1,\cdots,6$, if they exist, and let $\Omega$ be defined as shown in Fig. 4.8. Let $\Lambda=\mathbb{R}^{2}$. An entropy solution $(\rho,{\mathbf{v}})$ of the boundary value problem in $\Lambda$ in the sense of Definition 2.2 is called a regular Riemann solution with four-shock interactions for the isentropic Euler system if $(\rho,{\mathbf{v}})$ satisfies the following further properties:
  1. (i)
    The shock curves $\Gamma_{\rm shock}^{1}=P_{2}P_{3}$ and $\Gamma_{\rm shock}^{2}=P_{5}P_{6}$ are $C^{1}$ up to their endpoints and are $C^{2}$ in their relative interiors. Their tangents satisfy (4.6).
  2. (ii)
    $(\rho,{\mathbf{v}})$ is continuous in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{2})$ for some $r>0$.
  3. (iii)
    The solution coincides with the corresponding constant states in their respective regions $\Lambda_{i}$, for $i=1,\cdots,6$, if they exist: Specifically,
    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{in}\,\,\,\Lambda_{1},\\[1.42262pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in}\,\,\,\Lambda_{2},\\[1.42262pt] (\rho_{3},{\mathbf{v}}_{3})\qquad\mbox{in}\,\,\,\Lambda_{3},\\[1.42262pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in}\,\,\,\Lambda_{4},\\[1.42262pt] (\rho_{5},{\mathbf{v}}_{5})\qquad\mbox{in}\,\,\,\Lambda_{5},\quad\mbox{if state (5) at $P_{4}$ is supersonic}\\[1.42262pt] (\rho_{6},{\mathbf{v}}_{6})\qquad\mbox{in}\,\,\,\Lambda_{6},\quad\mbox{if state (6) at $P_{1}$ is supersonic},\end{cases}$
    where
    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{4}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{5},{\mathbf{v}}_{5})(P_{4})\qquad\mbox{if state (5) at $P_{4}$ is subsonic or sonic},$
    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{1}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{6},{\mathbf{v}}_{6})(P_{1})\qquad\mbox{if state (6) at $P_{1}$ is subsonic or sonic}.$
  4. (iv)
    ${\mathbf{v}}\cdot\nu\leq-C^{-1}$ on $\overline{\Gamma_{\rm shock}^{1}}\cup\overline{\Gamma_{\rm sonic}^{1}}\cup\overline{\Gamma_{\rm shock}^{2}}\cup\overline{\Gamma_{\rm sonic}^{2}}$ for some $C>0$, where ${\mathbf{v}}$ on the curves is taken from the $\Omega$–side and ${\boldsymbol{\nu}}$ is the outer normal with respect to $\Omega$.
  5. (v)
    The flow is pseudo-subsonic in $\Omega$ on and near $\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}$, except for possible sonic points $P_{2}$, $P_{3}$, $P_{5}$, and $P_{6}$.
If state $(5)$ at $P_{4}$ and state $(6)$ at $P_{1}$ are both supersonic, then $M=6$, ${N_{1}}=0$, and ${N_{2}}=4$ with $\Lambda_{i}$ (for $i=1,\cdots,6$) being given as in Fig. 4.8, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{2}P_{3}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{5}P_{6}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{1}=P_{2}P_{6}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{4}=\Gamma_{\rm sonic}^{2}=P_{3}P_{5}$.
If state $(5)$ at $P_{4}$ and state $(6)$ at $P_{1}$ are both subsonic, then $M=4$, ${N_{1}}=0$, and ${N_{2}}=2$ with $\Lambda_{i}$ (for $i=1,2,3,4$) being given as in Fig. 4.8, $\Lambda_{5}$ and $\Lambda_{6}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{1}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{4}P_{1}$.
If state $(5)$ at $P_{4}$ is supersonic and state $(6)$ at $P_{1}$ is subsonic, then $M=5$, ${N_{1}}=0$, and ${N_{2}}=2$ with $\Lambda_{i}$ (for $i=1,\cdots,5$) being given as in Fig. 4.8, $\Lambda_{6}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{1}P_{3}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{5}P_{1}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{2}=P_{3}P_{5}$.
It state $(5)$ at $P_{4}$ is subsonic and state $(6)$ at $P_{1}$ is supersonic, then $M=5$, ${N_{1}}=0$ and ${N_{2}}=3$ with $\Lambda_{i}$ (for $i=1,\cdots,6$) being given as in Fig. 4.8, $\Lambda_{5}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{2}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{4}P_{6}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{1}=P_{2}P_{6}$.
Then, following the proof of Theorem 4.10, we have the following theorem on the lower regularity of the regular Riemann solution, whose proof is omitted since it is similar to the one for Theorem 4.10:
Theorem 4.13.
Let $(\rho,{\mathbf{v}})$ be a regular Riemann solution with four-shock interactions for the isentropic Euler system and further satisfy:
  1. (i)
    $(\rho,{\mathbf{v}})\in C^{1}\big{(}({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega})\setminus\{P_{2},P_{3},P_{5},P_{6}\}\big{)}\cap C^{0,1}\big{(}{\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega}\big{)}$ for some $\sigma>0$, where $\Gamma_{\rm sonic}^{i}=\emptyset$ when it does not exist for $i=1,2$;
  2. (ii)
    $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
Appendix A DiPerna-Lions-Type Commutator Estimates
In this appendix, we show the DiPerna-Lions-type commutator estimates, which have been used in the proof of Theorem 3.2 in §3.
Lemma A.1.
Let $\Omega\subset\mathbb{R}^{n}$ be open and bounded, and $n\geq 2$. For $\varepsilon>0$, define $\eta_{\varepsilon}(\mathbf{x})=\frac{1}{\varepsilon^{n}}\eta(\frac{\mathbf{x}}{\varepsilon})$ for a mollifier kernel $\eta\in C^{\infty}_{\rm c}(\mathbb{R}^{n})$ satisfying $\eta(\mathbf{x})\geq 0$ and $\int_{\mathbb{R}^{n}}\eta(\mathbf{x})\,\rm d\mathbf{x}=1$. Let $b\in W^{1,p}_{\rm loc}(\Omega)$ and $u\in L^{q}_{\rm loc}(\Omega)$ for $p,q\in[1,\infty]$ such that $\frac{1}{p}+\frac{1}{q}\leq 1$. Then, for any $i=1,\cdots,n$,
$\partial_{x_{i}}\big{(}(bu)_{\varepsilon}-b_{\varepsilon}u_{\varepsilon}\big{)}\to 0\qquad\mbox{in $L^{1}_{\rm loc}(\Omega)$ as $\varepsilon\to 0^{+}$},$ (A.1)
where we have used the notation $F_{\varepsilon}:=F*\eta_{\varepsilon}$.
Proof.
Let $\mbox{supp}(\eta)\subset B_{M}$. Denote
$A_{\varepsilon}[u,b]:=\partial_{x_{i}}\big{(}(bu)_{\varepsilon}-b_{\varepsilon}u_{\varepsilon}\big{)}.$
We divide the proof into two steps.
1. We first show that, for all open $\Omega^{\prime}\Subset\Omega$ and $\varepsilon\in(0,\,\frac{1}{2M}\mbox{\rm dist}(\Omega^{\prime},\partial\Omega))$,
$\|A_{\varepsilon}[u,b]\|_{L^{1}(\Omega^{\prime})}\leq C(n,\eta,\Omega^{\prime})\|\nabla b\|_{L^{p}(\Omega)}\|u\|_{L^{q}(\Omega)}.$ (A.2)
Notice that
$A_{\varepsilon}[u,b]$ $=\partial_{x_{i}}\big{(}(bu)_{\varepsilon}-bu_{\varepsilon}\big{)}+\partial_{x_{i}}\big{(}(b_{\varepsilon}-b)u_{\varepsilon}\big{)}$
$=\big{(}\partial_{x_{i}}(bu)_{\varepsilon}-b\partial_{x_{i}}u_{\varepsilon}\big{)}-b_{x_{i}}u_{\varepsilon}+u_{\varepsilon}\partial_{x_{i}}(b_{\varepsilon}-b)+(b_{\varepsilon}-b)\partial_{x_{i}}u_{\varepsilon}$
$=:\sum_{m=1}^{4}I_{\varepsilon,m}.$
We now show that, for $\varepsilon\in(0,\,\frac{1}{2M}\mbox{\rm dist}(\Omega^{\prime},\partial\Omega))$,
$\|I_{\varepsilon,m}\|_{L^{1}(\Omega^{\prime})}\leq C(n,\eta,\Omega^{\prime})\|\nabla b\|_{L^{p}(\Omega)}\|u\|_{L^{q}(\Omega)}\qquad\mbox{for $m=1,\cdots,4$}.$ (A.3)
First, we estimate $I_{\varepsilon,1}=\partial_{x_{i}}(bu)_{\varepsilon}-b\partial_{x_{i}}(u)_{\varepsilon}$. For $\mathbf{x}\in\Omega^{\prime}$,
$I_{\varepsilon,1}(\mathbf{x})$ $=\int_{\mathbb{R}^{n}}\partial_{x_{i}}\big{(}\eta_{\varepsilon}(\mathbf{x}-\mathbf{y})\big{)}\,\big{(}b(\mathbf{y})-b(\mathbf{x})\big{)}u(\mathbf{y})\,{\rm d}\mathbf{y}$
$=\int_{B_{M}}\eta_{x_{i}}(\mathbf{y})\frac{b(\mathbf{x}-\varepsilon\mathbf{y})-b(\mathbf{x})}{\varepsilon}u(\mathbf{x}-\varepsilon\mathbf{y})\,{\rm d}\mathbf{y}.$
Then we use the estimate of the difference quotient for a Sobolev function to obtain that, for any $\varepsilon\in(0,\frac{1}{2M}\mbox{\rm dist}(\Omega^{\prime},\partial\Omega))$,
$\Big{\|}\frac{b(\cdot-\varepsilon\mathbf{y})-b(\cdot)}{\varepsilon}\Big{\|}_{L^{p}(\Omega^{\prime})}\leq C(n,\Omega^{\prime})\|\nabla b\|_{L^{p}(\Omega)}|\mathbf{y}|\qquad\,\mbox{for each $\mathbf{y}\in B_{M}$}.$
Noting that $\|u(\cdot-\varepsilon\mathbf{y})\|_{L^{q}(\Omega^{\prime})}\leq\|u\|_{L^{q}(\Omega)}$ for $\mathbf{y}$ and $\varepsilon$ as above, we have
$\|I_{\varepsilon,1}\|_{L^{1}(\Omega^{\prime})}\leq C(n,\Omega^{\prime})\|\nabla b\|_{L^{p}(\Omega)}\|u\|_{L^{q}(\Omega)}\int_{B_{M}}|\eta_{x_{i}}(\mathbf{y})|\,{\rm d}\mathbf{y},$
which is (A.3) for $m=1$.
For $I_{\varepsilon,2}=-b_{x_{i}}u_{\varepsilon}$ and $I_{\varepsilon,3}=u_{\varepsilon}\partial_{x_{i}}(b_{\varepsilon}-b)$, (A.3) follows from the standard estimates.
Finally, we estimate $I_{\varepsilon,4}=(b_{\varepsilon}-b)\partial_{x_{i}}u_{\varepsilon}$. For $\mathbf{x}\in\Omega^{\prime}$,
$I_{\varepsilon,4}(\mathbf{x})$ $=\int_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\eta_{\varepsilon}(\mathbf{x}-\mathbf{y})\partial_{x_{i}}\big{(}\eta_{\varepsilon}(\mathbf{x}-\mathbf{z})\big{)}\,\big{(}b(\mathbf{y})-b(\mathbf{x})\big{)}u(\mathbf{z})\,{\rm d}\mathbf{y}{\rm d}\mathbf{z}$
$=\int_{B_{M}\times B_{M}}\eta(\mathbf{y})\eta_{x_{i}}(\mathbf{z})\frac{b(\mathbf{x}-\varepsilon\mathbf{y})-b(\mathbf{x})}{\varepsilon}u(\mathbf{x}-\varepsilon\mathbf{z})\,{\rm d}\mathbf{y}{\rm d}\mathbf{z}.$
Then we complete the proof of (A.3) for $I_{\varepsilon,4}$ similar as for $I_{\varepsilon,1}$ above, by using the $L^{p}$–estimate the difference quotient for the $b$-term, and the $L^{q}$–estimate of the shift for the $u$-term. Then estimate (A.2) holds.
2. Now we prove the convergence (A.1). We first note that (A.1) holds if $u$ has the higher regularity $u\in W^{1,q}_{\rm loc}(\Omega)$ and $b$ as assumed. Indeed, in this case, $\partial_{x_{i}}(bu)\in L^{1}_{\rm loc}(\Omega)$. Since $\Omega$ is bounded,
$A_{\varepsilon}[u,b]$ $=\big{(}\partial_{x_{i}}(bu)\big{)}*\eta_{\varepsilon}-(\partial_{x_{i}}b)_{\varepsilon}u_{\varepsilon}-b_{\varepsilon}(\partial_{x_{i}}u)_{\varepsilon}$
$\to\partial_{x_{i}}(bu)-\partial_{x_{i}}bu-b\partial_{x_{i}}u=0\qquad\mbox{in $L^{1}_{\rm loc}(\Omega)\,\,$ as $\varepsilon\to 0$}.$
For general $u$ and $b$ as assumed, the same convergence is obtained by approximation, and using (A.2), via a standard argument that we briefly describe now. First, let $p\in[1,\infty)$. To include the case that $q=\infty$, we argue as follows: Choose open $\Omega^{\prime}\Subset\Omega$, then choose open $\Omega^{\prime\prime}$ such that $\Omega^{\prime}\Subset\Omega^{\prime\prime}\Subset\Omega$. Let $(b_{k},u_{k})\in C^{\infty}(\Omega)$ be such that $b_{k}\to b$ in $W^{1,p}_{\rm loc}(\Omega)$ and $u_{k}\to u$ in $L^{1}_{\rm loc}(\Omega)$. Replacing $\{u_{k}\}$ by its subsequence if necessary, we obtain
$\|\nabla b_{k}\|_{L^{\infty}(\Omega^{\prime\prime})}\|u-u_{k}\|_{L^{1}(\Omega^{\prime\prime})}\to 0\qquad\,\,\mbox{as $k\to\infty$.}$ (A.4)
Then, for each $\varepsilon\in(0,\frac{1}{2}\mbox{\rm dist}(\Omega^{\prime},\partial\Omega^{\prime\prime}))$, using the bi-linearity of $A_{\varepsilon}[\cdot,\cdot]$ and estimate (A.2) on sets $\Omega^{\prime}\Subset\Omega^{\prime\prime}$ for $(p,q)$ and $(\infty,1)$ respectively, we have
$\|A_{\varepsilon}[u,b]\|_{L^{1}(\Omega^{\prime})}$
$=\|A_{\varepsilon}[u_{k},b_{k}]+A_{\varepsilon}[u,b-b_{k}]+A_{\varepsilon}[u-u_{k},b_{k}]\|_{L^{1}(\Omega^{\prime})}$
$\leq\|A_{\varepsilon}[u_{k},b_{k}]\|_{L^{1}(\Omega^{\prime})}+C\big{(}\|\nabla b-\nabla b_{k}\|_{L^{p}(\Omega^{\prime\prime})}\|u\|_{L^{q}(\Omega^{\prime\prime})}+\|\nabla b_{k}\|_{L^{\infty}(\Omega^{\prime\prime})}\|u-u_{k}\|_{L^{1}(\Omega^{\prime\prime})}\big{)}.$
Recalling that $\|A_{\varepsilon}[u_{k},b_{k}]\|_{L^{1}(\Omega^{\prime})}\to 0$ as $\varepsilon\to 0$ for each $k$ and using (A.4), we obtain that $\|A_{\varepsilon}[u,b]\|_{L^{1}(\Omega^{\prime})}\to 0$ as $\varepsilon\to 0$. In the remaining case $(p,q)=(\infty,1)$, we argue similarly, interchanging $p$ and $q$ (resp. $b$ and $u$), and, instead of (A.4), we replace $\{b_{k}\}$ by its subsequence if necessary to obtain
$\|\nabla b-\nabla b_{k}\|_{L^{1}(\Omega^{\prime\prime})}\|u_{k}\|_{L^{\infty}(\Omega^{\prime\prime})}\to 0\qquad\mbox{as $k\to\infty$}.$
Then we modify the rest of the argument correspondingly to conclude the proof. ∎
Acknowledgements. The research of Gui-Qiang G. Chen was also supported in part by the UK Engineering and Physical Sciences Research Council Award EP/L015811/1, EP/V008854, and EP/V051121/1. The research of Mikhail Feldman was supported in part by the National Science Foundation under Grants DMS-2054689 and DMS-2219391, and the Steenbock Professorship Award from the University of Wisconsin-Madison. The research of Wei Xiang was supported in part by the Research Grants Council of the HKSAR, China (Project No. CityU 11304820, CityU 11300021, CityU 11311722, and CityU 11305523), and in part by the Research Center for Nonlinear Analysis of the Hong Kong Polytechnic University. For the purpose of open access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
Conflict of Interest: The authors declare that they have no conflict of interest. The authors also declare that this manuscript has not been previously published, and will not be submitted elsewhere before your decision.
Data availability: Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.

### v2

@@
 We establish a general framework for the analysis of the local regularity of such solutions for a class of two-dimensional Riemann problems for the isentropic Euler system, which includes the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the four-shock Riemann problem.
-We prove that it is not possible that both the density and the velocity are in $H^{1}$ in the subsonic domain for the self-similar solutions of these problems in general.
-This indicates that the self-similar solutions of the Riemann problems with shocks for the isentropic Euler system are of much more complicated structure than those for the Euler system for potential flow; in particular, the density and the velocity are not necessarily continuous in the subsonic domain.
+We prove that the velocity is not in $H^{1}$ in the subsonic domain for the self-similar solutions of these problems in general.
+This indicates that the self-similar solutions of the Riemann problems with shocks for the isentropic Euler system are of much more complicated structure than those for the Euler system for potential flow; in particular, the velocity is not necessarily continuous in the subsonic domain.
 The proof is based on a regularization of the isentropic Euler system to derive the transport equation for the vorticity, a renormalization argument extended to the case of domains with boundary, and DiPerna-Lions-type commutator estimates.
 Key words and phrases:
-Low regularity, transonic shock, free boundary, compressible flow, potential flow, self-similar, conservation laws, shock reflection-diffraction, regular shock reflection, Prandtl reflection, Lighthill diffraction, nonlinear, fine properties
+Low regularity, transonic shock, free boundary, compressible flow, potential flow, self-similar, conservation laws, shock reflection-diffraction, regular shock reflection, Prandtl reflection, Lighthill diffraction, nonlinear, fine properties.
+*Corresponding author: Gui-Qiang G. Chen
 2010 Mathematics Subject Classification:
@@
 In particular, the solution has a high regularity in the subsonic domain $\Omega$ (where the solution is not a constant state); see Figs.
-4.24.2 in which density $\rho$ and velocity ${\mathbf{v}}$ are in $C^{\alpha}(\overline{\Omega})\cap C^{\infty}(\Omega)$ for some $\alpha\in(0,1)$.
+33 in which density $\rho$ and velocity $\mathbf{v}$} are in $C^{\alpha}(\overline{\Omega})\cap C^{\infty}(\Omega)$ for some $\alpha\in(0,1)$.
 More precisely, solution $(\rho,{\mathbf{v}})$ is in a weighted Hölder space which implies that $(\rho,{\mathbf{v}})$ is in the Sobolev space $W^{1,p}(\Omega)$ for some $p>2$.
@@
 Since the calculation is formal, it is important to find out whether the low regularity indeed necessarily holds for the regular shock reflection solutions.
-In this paper, under the natural assumptions on the regularity of the self-similar solutions in the subsonic domain $\Omega$ near the shocks, we rigorously prove that it is not possible indeed that both the density and the velocity are in $H^{1}(\Omega)$.
+In this paper, under the natural assumptions on the regularity of the self-similar solutions in the subsonic domain $\Omega$ near the shocks, we rigorously prove that the velocity is not in $H^{1}(\Omega)$ indeed.
 The argument is based on the vortical singularity calculation in [44].
@@
 These self-similar solutions include the above-mentioned shock reflection problems: the regular shock reflection problem, the Prandtl reflection problem, the Lighthill diffraction problem, and the four-shock Riemann problem.
-This indicates that the self-similar solutions of the Riemann problems with shocks for the isentropic Euler system have much more complicated structure than the corresponding solutions for the Euler system for potential flow; in particular, the density and the velocity are not necessarily continuous in $\Omega$, at least their continuity can not be obtained directly by the Sobolev embeddings.
+This indicates that the self-similar solutions of the Riemann problems with shocks for the isentropic Euler system have much more complicated structure than the corresponding solutions for the Euler system for potential flow; in particular, the velocity is not necessarily continuous in $\Omega$, at least their continuity can not be obtained directly by the Sobolev embeddings.
 On the other hand, this argument allows the possibility that $(\rho,{\mathbf{v}})\in W^{1,p}(\Omega)$ for some or even all $p\in[1,2)$, in which case there are no shocks in $\Omega$.
@@
 $({\mathbf{v}}\cdot\nabla){\mathbf{v}}+{\mathbf{v}}+\nabla h(\rho)=0,$ (2.5)
-where $h(\rho)=\frac{\rho^{\gamma-1}}{\gamma-1}$ is the enthalpy. Note that the speed of sound is $c=\rho^{\frac{\gamma-1}{2}}=\sqrt{(\gamma-1)h(\rho)}$.
+where $h(\rho)=\frac{\rho^{\gamma-1}-1}{\gamma-1}$ is the enthalpy. Note that the speed of sound is $c=\rho^{\frac{\gamma-1}{2}}=\sqrt{(\gamma-1)h(\rho)+1}$.
 2.2. Entropy Solutions and the Rankine-Hugoniot Conditions
@@
 Suppose that $S$ is a smooth curve in domain $\Lambda$.
-An entropy solution of (2.3)–(2.4), which is $C^{1}$ near and up to $S$ on both sides of $S$, satisfies the following Rankine-Hugoniot (R-H) conditions on $S$:
-$[\rho{\mathbf{v}}\cdot\boldsymbol{\nu}]=0,\quad\,[(\rho{\mathbf{v}}\cdot\boldsymbol{\nu}){\mathbf{v}}+p\boldsymbol{\nu}]={\bf 0}\qquad\,\,\mbox{on $S$},$ (2.9)
-[…]
-If $\rho^{+}{\mathbf{v}}^{+}\cdot\boldsymbol{\nu}=0$ on $S$, then $\rho^{-}{\mathbf{v}}^{-}\cdot\boldsymbol{\nu}=0$ on $S$ by the first equality in (2.9), and $[p]=0$ from the second equality in (2.9) which implies that $[\rho]=0$.
+An entropy solution of (2.3)–(2.4), which is $C^{1}$ in the open domains near $S$ and $C^{0}$ up to $S$ on both sides, satisfies the following Rankine-Hugoniot (R-H) conditions on $S$:
+$\rho{\mathbf{v}}\cdot\boldsymbol{\nu}]=0,\quad\,[(\rho{\mathbf{v}}\cdot\boldsymbol{\nu}){\mathbf{v}}+p\boldsymbol{\nu}]={\bf 0}\qquad\,\,\mbox{on $S$},$ (2.9)
+where $\boldsymbol{\nu}$ is a unit normal to $S$, and $\cdot\,$] denotes the difference of the concerned quantity across $S$.
+Denote by $(\rho^{\pm},{\mathbf{v}}^{\pm})$ the values of solution $(\rho,{\mathbf{v}})$ on the $\pm$ sides of $S$.
+Assume that $\rho^{\pm}>0$ on $S$ and some of $(\rho,{\mathbf{v}})$ are discontinuous across $S$.
+If $\rho^{+}{\mathbf{v}}^{+}\cdot\boldsymbol{\nu}=0$ on $S$, then $\rho^{-}{\mathbf{v}}^{-}\cdot\boldsymbol{\nu}=0$ on $S$ by the first equality in (2.9), and $[p]=0$ from the second equality in (2.9) which implies that $\rho]=0$.
 This discontinuity is called a vortex sheet.
-In this case, $[{\mathbf{v}}\cdot\boldsymbol{\tau}]\neq 0$ on $S$, unless the solution is continuous across $S$, where $\boldsymbol{\tau}$ is a unit tangent vector to $S$.
+In this case, ${\mathbf{v}}\cdot\boldsymbol{\tau}]\neq 0$ on $S$, unless the solution is continuous across $S$, where $\boldsymbol{\tau}$ is a unit tangent vector to $S$.
 If $\rho^{+}{\mathbf{v}}^{+}\cdot\boldsymbol{\nu}\neq 0$ on $S$, then $\rho^{-}{\mathbf{v}}^{-}\cdot\boldsymbol{\nu}\neq 0$ on $S$ by the first equality in (2.9).
 This discontinuity is called a shock.
-In this case, $\rho{\mathbf{v}}\cdot\boldsymbol{\nu}$ is continuous across $S$ from the first equality in (2.9), and $[{\mathbf{v}}\cdot\boldsymbol{\tau}]=0$ and $[p]=-\rho{\mathbf{v}}\cdot\boldsymbol{\nu}[{\mathbf{v}}\cdot\boldsymbol{\nu}]$ from the second equality in (2.9).
-This implies that $[{\mathbf{v}}\cdot\boldsymbol{\nu}]\neq 0$, unless the solution is continuous across $S$, so that $[\rho]\neq 0$ from the first equality in (2.9).
+In this case, $\rho{\mathbf{v}}\cdot\boldsymbol{\nu}$ is continuous across $S$ from the first equality in (2.9), and ${\mathbf{v}}\cdot\boldsymbol{\tau}]=0$ and $p]=-\rho{\mathbf{v}}\cdot\boldsymbol{\nu}[{\mathbf{v}}\cdot\boldsymbol{\nu}]$ from the second equality in (2.9).
+This implies that ${\mathbf{v}}\cdot\boldsymbol{\nu}]\neq 0$, unless the solution is continuous across $S$, so that $\rho]\neq 0$ from the first equality in (2.9).
 Also, it follows from $\rho^{\pm}>0$ on $S$ that $({\mathbf{v}}^{+}\cdot\boldsymbol{\nu})({\mathbf{v}}^{-}\cdot\boldsymbol{\nu})>0$ on $S$.
@@
 We also give a definition of entropy solutions of system (2.3)–(2.4) in $\Lambda$ with the slip boundary conditions:
-[…]
+${\mathbf{v}}\cdot{\boldsymbol{\nu}}=0\qquad\mbox{on $\partial\Lambda$},$ (2.13)
 where $\boldsymbol{\nu}$} is the outer normal to $\partial\Lambda$.
@@
 $\int_{\Lambda}\big(\rho{\mathbf{v}}\otimes{\mathbf{v}}:D\zeta-3\rho{\mathbf{v}}\cdot\zeta+p\,\mbox{\rm div}\zeta\big)\,{\rm d}{\boldsymbol{\xi}}-\int_{\partial\Lambda}p\,\zeta\cdot{\boldsymbol{\nu}}\,{\rm d}l=0$ (2.14)
-are satisfied for any test function $\zeta\in C^{\infty}_{\rm c}(\mathbb{R}^{2};\,\mathbb{R}^{2})$.
-Note that here the test function is not required to vanish on $\partial\Lambda$.
+are satisfied for any test function $\zeta\in C^{\infty}_{\rm c}(\mathbb{R}^{2};\,\mathbb{R}^{2})$. Note that here the test function is not required to vanish on $\partial\Lambda$.
+From now on, we define the $r$-neighborhood of a curve $\Gamma$ as
+$\mathcal{N}_{r}(\Gamma):=\{{\boldsymbol{\xi}}\,:\,\mbox{\rm dist}\{{\boldsymbol{\xi}},\Gamma\}<r\}.$ (2.15)
 Based on that, we define the notion of entropy solutions of the boundary value problem (2.3)–(2.4) and (2.13) as follows:
@@
 These are achieved by carefully analyzing the vorticity function $\omega:=\partial_{1}v_{2}-\partial_{2}v_{1}$ for the pseudo-velocity ${\mathbf{v}}=(v_{1},v_{2})$.
-Figure 3.1. The Riemann Problem in a General Setting
+Figure 1. The Riemann Problem in a General Setting
 More precisely, as shown in Fig.
-3.1, we consider a Riemann problem in a self-similar wedge domain $\Lambda$ in $\mathbb{R}^{2}$ with the wedge-vertex being the origin, which also includes the case of the whole space.
+1, we consider a Riemann problem in a self-similar wedge domain $\Lambda$ in $\mathbb{R}^{2}$ with the wedge-vertex being the origin, which also includes the case of the whole space.
 That is, in polar coordinates,
-either Λ=2 or $\mbox{either $\,\,\Lambda=\mathbb{R}^{2}\,\,\,$ or $\,\,\Lambda=\big{\{}(r,\theta)\;:\;\theta^{-}<\theta<\theta^{+}\big{\}}$}.$ (3.1)
+$\mbox{either $\,\,\Lambda=\mathbb{R}^{2}\,\,\,$ or $\,\,\Lambda=\big\{(r,\theta)\;:\;\theta^{-}<\theta<\theta^{+}\big\}$}.$ (3.1)
 For a Riemann problem in a domain $\Lambda$ whose boundary contains a wedge-boundary, the initial data are the given constant velocity and density in each sub-sector of $\Lambda$ such that the Rankine-Hugoniot conditions hold on the lines separating them.
-Motivated by the expected configurations of self-similar solutions of the Riemann problems in §4, we consider solutions in self-similar coordinates of the following form, in terms of the density and pseudo-velocity $(\rho,{\mathbf{v}})$ with $\mathcal{N}_{r}(\Gamma):=\{{\boldsymbol{\xi}}\,:\,\mbox{\rm dist}\{{\boldsymbol{\xi}},\Gamma\}<r\}$:
+Motivated by the expected configurations of self-similar solutions of the Riemann problems in §4, we consider solutions in self-similar coordinates of the following form, in terms of the density and pseudo-velocity $(\rho,{\mathbf{v}})$:
 Definition 3.1.
-We say that the vector function $(\rho,{\mathbf{v}})$ on $\Lambda$ is an entropy solution of a Riemann problem of admissible structure if $(\rho,{\mathbf{v}})\in L^{\infty}(\Lambda)$, with $\rho\in BV_{\rm loc}(\Lambda\cap{\mathcal{N}}_{r}(\partial\Lambda))$ for some $r>0$, is an entropy solution of system (2.3)–(2.4) with the slip boundary condition (2.13) in the sense of Definition 2.2, which satisfies the following properties:
+We say that the vector function $(\rho,{\mathbf{v}})$ on $\Lambda$ is an entropy solution of the Riemann problem of admissible structure if $(\rho,{\mathbf{v}})\in L^{\infty}(\Lambda)$, with $\rho\in BV_{\rm loc}(\Lambda\cap{\mathcal{N}}_{r}(\partial\Lambda))$ for some $r>0$, is an entropy solution of system (2.3)–(2.4) with the slip boundary condition (2.13) in the sense of Definition 2.2, which satisfies the following properties:
   1. (i)
-    There exists an open, bounded, connected set $\Omega\subset\Lambda$, an integer $M\geq 1$, and open connected sets $\Lambda_{i}$, $i=1,\cdots,M$, pairwise disjoint, such that
-    $\overline{\Lambda}\setminus\Omega=\cup_{i=1}^{M}\overline{\Lambda_{i}},$
-    and $(\rho,{\mathbf{v}})$ is a constant state in each $\Lambda_{i}$, i.e., $(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{i},{\mathbf{u}}_{i}-{\boldsymbol{\xi}})$ in $\Lambda_{i}$, where $\rho_{i}>0$ is a constant and ${\mathbf{u}}_{i}$ is a constant vector.
+    There exist an open, bounded, connected set $\Omega\subset\Lambda$, an integer $M\geq 1$, and open connected sets $\Lambda_{i}$, $i=1,\cdots,M$, pairwise disjoint, such that ΛΩ=_i=1^MΛ_i, and $(\rho,{\mathbf{v}})$ is a constant state in each $\Lambda_{i}$, i.e., $(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{i},{\mathbf{u}}_{i}-{\boldsymbol{\xi}})$ in $\Lambda_{i}$, where $\rho_{i}>0$ is a constant and ${\mathbf{u}}_{i}$ is a constant vector.
   2. (ii)
-    If, for $i\neq j$, sets $\Lambda_{i}$ and $\Lambda_{j}$ have common boundary within $\Lambda$; that is, if $\partial\Lambda_{i}\cap\partial\Lambda_{j}\cap\Lambda\neq\emptyset$, then the corresponding constant states are not equal to each other: $(\rho_{i},{\mathbf{v}}_{i})\neq(\rho_{j},{\mathbf{v}}_{j})$.
+    If, for $i\neq j$, sets $\Lambda_{i}$ and $\Lambda_{j}$ have common boundary within $\Lambda$, i.e., if $\partial\Lambda_{i}\cap\partial\Lambda_{j}\cap\Lambda\neq\emptyset$, then the corresponding constant states are not equal to each other: $(\rho_{i},{\mathbf{v}}_{i})\neq(\rho_{j},{\mathbf{v}}_{j})$.
   3. (iii)
-    $\partial\Omega$ is Lipschitz. Denote
-    $\Gamma^{\mbox{\rm\footnotesize ext}}=\partial\Omega\cap\partial\Lambda,\qquad\Gamma^{\mbox{\rm\footnotesize int}}=\partial\Omega\cap\Lambda,$
-    so that $\partial\Omega=\Gamma^{\mbox{\rm\footnotesize ext}}\cup\Gamma^{\mbox{\rm\footnotesize int}}$.
+    $\partial\Omega$ is Lipschitz.
+Denote Γ^ext=Ω∩∂Λ,   Γ^int=ΩΛ, so that $\partial\Omega=\Gamma^{\mbox{\rm\footnotesize ext}}\cup\Gamma^{\mbox{\rm\footnotesize int}}$.
 It is possible that $\Gamma^{\mbox{\rm\footnotesize ext}}=\emptyset$ (in particular, this is true when $\partial\Lambda=\emptyset$).
   4. (iv)
-    If $\partial\Lambda\neq\emptyset$, then ${\Gamma^{\mbox{\rm\footnotesize ext}}}=\cup_{i=1}^{N_{1}}\overline{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}}$, where ${N_{1}}\geq 1$ and each $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ is a relatively open segment of straight line.
+    If $\partial\Lambda\neq\emptyset$, then ${\Gamma^{\mbox{\rm\footnotesize ext}}}=\cup_{i=1}^{N_{1}}\overline{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}}$, where ${N_{1}}\geq 1$, each $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ is a relatively open segment of straight line.
 Segments $\{\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\}_{i=1}^{N_{1}}$ are disjoint and, if $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ and $\Gamma^{\mbox{\rm\footnotesize ext}}_{j}$ have a common endpoint, then the interior angle (for $\Omega$) is within $(0,2\pi)$ at that point.
@@
   5. (v)
-    $\overline{\Gamma^{\mbox{\rm\footnotesize int}}}=\cup_{i=1}^{N_{2}}\overline{\Gamma^{\mbox{\rm\footnotesize int}}_{i}}$, where ${N_{2}}\geq 1$ and each $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ is a relatively open segment of curve, and segments $\{\Gamma^{\mbox{\rm\footnotesize int}}_{i}\}_{i=1}^{N_{2}}$ are disjoint.
+    $\overline{\Gamma^{\mbox{\rm\footnotesize int}}}=\cup_{i=1}^{N_{2}}\overline{\Gamma^{\mbox{\rm\footnotesize int}}_{i}}$, where ${N_{2}}\geq 1$, each $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ is a relatively open segment of curve, and segments $\{\Gamma^{\mbox{\rm\footnotesize int}}_{i}\}_{i=1}^{N_{2}}$ are disjoint.
 Each $\Gamma^{\mbox{\rm\footnotesize int}}_{i}$ is $C^{2}$ in its relative interior and $C^{1}$ up to the endpoints.
@@
   8. (viii)
-    There exists a point $\hat{P}$ in the relative interior of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ such that the curvature of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is non-zero at $\hat{P}$, and $({\mathbf{v}}\cdot{\boldsymbol{\tau}})(\hat{P})\neq 0$, where ${\boldsymbol{\tau}}(\hat{P})$ is a unit tangent vector to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at $\hat{P}$.
+[…]
 Then we have the following main theorem.
 Theorem 3.2 (Main Theorem).
-Let $(\rho,{\mathbf{v}})$ be a solution of a Riemann problem of admissible structure in the sense of Definition 3.1. Assume that $(\rho,{\mathbf{v}})$ satisfy
+Let $(\rho,{\mathbf{v}})$ be a solution of the Riemann problem of admissible structure in the sense of Definition 3.1. Assume that $(\rho,{\mathbf{v}})$ satisfy
   1. (i)
@@
     The flow is subsonic on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ from the $\Omega$–side: $|{\mathbf{v}}|<c$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, where $c=\rho^{\frac{\gamma-1}{2}}$ is the speed of sound.
-Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
+Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
 Proof.
-We divide the proof into six steps.
-1. We prove the theorem by establishing that, under the assumption:
+We divide the proof into seven steps. In the first six steps, we prove that both $\rho$ and $\mathbf{v}$} can not be in $H^{1}(\Omega)$:
+$(\rho,{\mathbf{v}})\notin H^{1}(\Omega).$ (3.2)
+This will be proved by showing that, under the assumption that
 $(\rho,{\mathbf{v}})\in H^{1}(\Omega),$ (3.3)
 the calculations of vortical singularity can rigorously be justified, which leads to the contradiction.
-Under assumption (3.2), vorticity $\omega=\nabla\times{\mathbf{v}}$ satisfies $\omega\in L^{2}(\Omega)$ and
+Then, in Step 7, we prove that $\mathbf{v}$} is not in $H^{1}(\Omega)$ indeed: ${\mathbf{v}}\notin H^{1}(\Omega)$.
+1. In Steps 1–6 of this proof, we assume that (3.3) holds. Then vorticity $\omega=\nabla\times{\mathbf{v}}$ satisfies $\omega\in L^{2}(\Omega)$ and
 $X:=\frac{\omega}{\rho}\in L^{2}(\Omega),$ (3.4)
@@
 for any $f\in C^{1}(\mathbb{R})$ with $g$ defined by (3.9).
-From the weak form (2.6) of the conservation law of mass, which holds for all $\phi$ specified in Definition 2.2, and the regularity assumption in (3.2), we obtain that $\rho{\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ holds ${\mathcal{H}}^{1}$a.e.
+From the weak form (2.6) of the conservation law of mass, which holds for all $\phi$ specified in Definition 2.2, and the regularity assumption (3.3), we obtain that $\rho{\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ holds ${\mathcal{H}}^{1}$a.e.
 on $\Gamma^{\mbox{\rm\footnotesize ext}}$ in the sense of traces.
@@
 $\int_{\Omega}\big(\rho f(X){\mathbf{v}}\cdot\nabla\zeta+\rho g(X)\zeta\big){\rm d}{\boldsymbol{\xi}}-\int_{\Gamma^{\mbox{\rm\footnotesize int}}}\rho f(X){\mathbf{v}}\cdot{\boldsymbol{\nu}}\,\zeta\,{\rm d}l=0\qquad\mbox{ for all }\;\zeta\in C^{\infty}(\mathbb{R}^{2})$ (3.12)
-for any $f\in C^{1}(\mathbb{R}^{1})$ with $\|f^{\prime}\|_{C^{0,1}(\mathbb{R}^{1})}<\infty$.
+for any $f\in C^{1}(\mathbb{R})$ with $\|f^{\prime}\|_{C^{0,1}(\mathbb{R})}<\infty$.
 Notice that the boundary integral in (3.12) is taken along a part of $\partial\Omega$.
@@
 Equation (3.12) shows that
-$\rho f(X){\mathbf{v}}\cdot{\boldsymbol{\nu}}=0\qquad\mbox{on $\Gamma^{\mbox{\rm\footnotesize ext}}$ in the weak sense.}$.
+$\rho f(X){\mathbf{v}}\cdot{\boldsymbol{\nu}}=0\qquad\mbox{on $\Gamma^{\mbox{\rm\footnotesize ext}}$ in the weak sense.}$
 Furthermore, the integrand in the boundary integral in (3.12) is well-defined by assumption (i) and Definition 3.1(v).
@@
 Also, there exists $r_{1}>0$ such that, for $i=1,\cdots,{N_{1}}$,
-$\Gamma^{\mbox{\rm\footnotesize ext}}_{i}\cap B_{(L+1)r_{1}}(\mathcal{P}_{0})=\emptyset\qquad\mbox{if $\mathcal{P}_{0}$ is not an endpoint of }\Gamma^{\mbox{\rm\footnotesize ext}}_{i}.$ (3.13)
+[…]
 Now, for each $r\in(0,r_{1}]$, define
@@
 where we recall that $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$ is a relatively open line segment.
-Using (3.2) and assumption (ii), and noting that ${\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ to make the odd extension of ${\mathbf{v}}\cdot{\boldsymbol{\nu}}$ across $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$, we find that the extended $(\rho,{\mathbf{v}})$ satisfies $(\rho,{\mathbf{v}})\in H^{1}((\Omega\cup\overline{V_{i}^{r_{1}}})^{0})\cap L^{\infty}(\Omega\cup\overline{V_{i}^{r_{1}}})$ and $\rho\geq C_{0}^{-1}$ on $\Omega\cup\overline{V_{i}^{r_{1}}}$.
+Using (3.3) and (ii), and noting that ${\mathbf{v}}\cdot{\boldsymbol{\nu}}=0$ to make the odd extension of ${\mathbf{v}}\cdot{\boldsymbol{\nu}}$ across $\Gamma^{\mbox{\rm\footnotesize ext}}_{i}$, we find that the extended $(\rho,{\mathbf{v}})$ satisfies $(\rho,{\mathbf{v}})\in H^{1}((\Omega\cup\overline{V_{i}^{r_{1}}})^{0})\cap L^{\infty}(\Omega\cup\overline{V_{i}^{r_{1}}})$ and $\rho\geq C_{0}^{-1}$ on $\Omega\cup\overline{V_{i}^{r_{1}}}$.
 Combining the extensions for $i=1,\cdots,{N_{1}}$, and using (3.13)–(3.14), we obtain that, for each $r\in(0,r_{1}]$, $(\rho,{\mathbf{v}})$ is extended to the domain:
@@
 Lemma 3.3.
-Under assumptions (i)–(ii) of Theorem 3.2 and (3.2), there exists $r_{1}>0$ depending only on $\Omega$ such that, for any $r\in(0,r_{1}]$, there exists an extension of $(\rho,{\mathbf{v}})$ into $\Omega^{\rm ext}_{r}$, still denoted as $(\rho,{\mathbf{v}})$, such that
+Under assumptions (i)–(ii) of Theorem 3.2 and (3.3), there exists $r_{1}>0$ depending only on $\Omega$ such that, for any $r\in(0,r_{1}]$, there is an extension of $(\rho,{\mathbf{v}})$ into $\Omega^{\rm ext}_{r}$, still denoted as $(\rho,{\mathbf{v}})$, so that
   1. (a)
@@
 Then we define $\eta_{\varepsilon}({\boldsymbol{\xi}})=\frac{1}{\varepsilon^{2}}\eta(\frac{{\boldsymbol{\xi}}}{\varepsilon})$ for $\varepsilon>0$, and denote $F_{\varepsilon}:=F*\eta_{\varepsilon}$ for various functions $F$, specifically
-${\mathbf{v}}_{\varepsilon}={\mathbf{v}}*\eta_{\varepsilon},\quad\omega_{\varepsilon}=\omega*\eta_{\varepsilon},\quad\rho_{\varepsilon}=\rho*\eta_{\varepsilon}\quad(i(\rho))_{\varepsilon}=i(\rho)*\eta_{\varepsilon}\qquad\,\,\mbox{for $\varepsilon\in(0,\frac{\delta}{2})$}.$
+${\mathbf{v}}_{\varepsilon}={\mathbf{v}}*\eta_{\varepsilon},\quad\omega_{\varepsilon}=\omega*\eta_{\varepsilon},\quad\rho_{\varepsilon}=\rho*\eta_{\varepsilon}\quad(h(\rho))_{\varepsilon}=h(\rho)*\eta_{\varepsilon}\qquad\,\,\mbox{for $\varepsilon\in(0,\frac{\delta}{2})$}.$
 Now, since $(\rho,{\mathbf{v}})\in H^{1}(\Omega^{\rm ext}_{r})\cap L^{\infty}(\Omega^{\rm ext}_{r})$ with $\rho\geq C^{-1}$ by Lemma 3.3(a)–(b), and $\Omega^{r}\Subset\Omega^{\rm ext}_{r}$, we obtain that, for each $\varepsilon\in(0,\frac{\delta}{2})$,
-$\rho_{\varepsilon},\,{\mathbf{v}}_{\varepsilon},\,(i(\rho))_{\varepsilon}\in H^{1}(\Omega^{r})\cap L^{\infty}(\Omega^{r}),\quad\|(\rho_{\varepsilon}\,{\mathbf{v}}_{\varepsilon})\|_{L^{\infty}(\Omega^{r})}\leq\|(\rho\,{\mathbf{v}})\|_{L^{\infty}(\Omega^{\rm ext}_{r})},$ (3.23)
+$\rho_{\varepsilon},\,{\mathbf{v}}_{\varepsilon},\,(h(\rho))_{\varepsilon}\in H^{1}(\Omega^{r})\cap L^{\infty}(\Omega^{r}),\quad\|(\rho_{\varepsilon},\,{\mathbf{v}}_{\varepsilon})\|_{L^{\infty}(\Omega^{r})}\leq\|(\rho,\,{\mathbf{v}})\|_{L^{\infty}(\Omega^{\rm ext}_{r})},$ (3.24)
 $(\rho_{\varepsilon},\,{\mathbf{v}}_{\varepsilon})\to(\rho,\,{\mathbf{v}})\,\,\,\,\mbox{in $H^{1}(\Omega^{r})$},\quad\,\omega_{\varepsilon}\to\omega\,\,\,\,\mbox{in $L^{2}(\Omega^{r})\qquad\,\,$ as $\varepsilon\to 0$}.$
@@
 $\mbox{\rm div}(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon})+2\rho_{\varepsilon}+r^{(1)}_{\varepsilon}=0,$ (3.26)
-$({\mathbf{v}}_{\varepsilon}\cdot\nabla){\mathbf{v}}_{\varepsilon}+{\mathbf{v}}_{\varepsilon}+\nabla(i(\rho))_{\varepsilon}+{\bf r}^{(2)}_{\varepsilon}=0,$ (3.26)
+$({\mathbf{v}}_{\varepsilon}\cdot\nabla){\mathbf{v}}_{\varepsilon}+{\mathbf{v}}_{\varepsilon}+\nabla(h(\rho))_{\varepsilon}+\mathbf{r}^{(2)}_{\varepsilon}=0,$ (3.27)
 where
-$r^{(1)}_{\varepsilon}=(\mbox{\rm div}\,(\rho{\mathbf{v}}))_{\varepsilon}-\mbox{\rm div}\,(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon}),\qquad{\bf r}^{(2)}_{\varepsilon}=\big{(}({\mathbf{v}}\cdot\nabla){\mathbf{v}}\big{)}_{\varepsilon}-({\mathbf{v}}_{\varepsilon}\cdot\nabla){\mathbf{v}}_{\varepsilon}.$
+$r^{(1)}_{\varepsilon}=(\mbox{\rm div}\,(\rho{\mathbf{v}}))_{\varepsilon}-\mbox{\rm div}\,(\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon}),\qquad\mathbf{r}^{(2)}_{\varepsilon}=\big(({\mathbf{v}}\cdot\nabla){\mathbf{v}}\big)_{\varepsilon}-({\mathbf{v}}_{\varepsilon}\cdot\nabla){\mathbf{v}}_{\varepsilon}.$
 The functions on the left-hand side of (3.26)–(3.27) are smooth. Taking the curl of (3.27), we obtain
-${\mathbf{v}}_{\varepsilon}\cdot\nabla\omega_{\varepsilon}+(1+\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon})\omega_{\varepsilon}+\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}=0\qquad\,\,\mbox{in $\Omega^{r}$}.$ (3.27)
+${\mathbf{v}}_{\varepsilon}\cdot\nabla\omega_{\varepsilon}+(1+\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon})\omega_{\varepsilon}+\mbox{\rm curl}\,\mathbf{r}^{(2)}_{\varepsilon}=0\qquad\,\,\mbox{in $\Omega^{r}$}.$ (3.28)
 Denote
@@
 $\rho_{\varepsilon}{\mathbf{v}}_{\varepsilon}\cdot\nabla X^{(\varepsilon)}$ $={\mathbf{v}}_{\varepsilon}\cdot\nabla\omega_{\varepsilon}-\nabla\rho_{\varepsilon}\cdot{\mathbf{v}}_{\varepsilon}X^{(\varepsilon)}$
-$=\rho_{\varepsilon}X^{(\varepsilon)}+r^{(1)}_{\varepsilon}-\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}.$
+$=\rho_{\varepsilon}X^{(\varepsilon)}+r^{(1)}_{\varepsilon}-\mbox{\rm curl}\,\mathbf{r}^{(2)}_{\varepsilon}$.
 From this and (3.26), we have
@@
 where
-${\mathcal{R}}_{\varepsilon}=f^{\prime}(X^{(\varepsilon)})\big{(}r^{(1)}_{\varepsilon}-\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}\big{)}-f(X^{(\varepsilon)})r^{(1)}_{\varepsilon}.$ (3.30)
+${\mathcal{R}}_{\varepsilon}=f^{\prime}(X^{(\varepsilon)})\big(r^{(1)}_{\varepsilon}-\mbox{\rm curl}\,\mathbf{r}^{(2)}_{\varepsilon}\big)-f(X^{(\varepsilon)})r^{(1)}_{\varepsilon}$. (3.31)
 Recalling definition (3.9) of $g(\cdot)$, we rewrite (3.30) as
 $\mbox{\rm div}\big(\rho_{\varepsilon}f(X^{(\varepsilon)}){\mathbf{v}}_{\varepsilon}\big)$ $=\rho_{\varepsilon}g(X^{(\varepsilon)})+{\mathcal{R}}_{\varepsilon}\qquad\mbox{in $\Omega^{r}$}.$
-Let $\zeta\in C^{\infty}(\mathbb{R}^{2})$ satisfy (3.21). Multiply the last equation by $\zeta$, integrate over $\Omega$, and integrate by parts via using (3.21) and (3.24) to obtain
+Let $\zeta\in C^{\infty}(\mathbb{R}^{2})$ satisfy (3.22). Multiply the last equation by $\zeta$, integrate over $\Omega$, and integrate by parts with the use of (3.22) and (3.25) to obtain
 $\int_{\Omega^{r}}\big(\rho_{\varepsilon}f(X^{(\varepsilon)}){\mathbf{v}}_{\varepsilon}\cdot\nabla\zeta+\big(\rho_{\varepsilon}g(X^{(\varepsilon)})+{\mathcal{R}}_{\varepsilon}\big)\zeta\big)\,{\rm d}\xi=0,$ (3.32)
@@
 Next, we show that
-${\mathcal{R}}_{\varepsilon}\to 0\qquad\mbox{in $L^{1}(\Omega^{r})\;\;$ as $\varepsilon\to 0$}.$ (3.34)
+[…]
 From (3.31) and (3.33) and the fact that $\Omega^{r}$ is a bounded domain, in order to prove (3.35), it suffices to show that
 $r^{(1)}_{\varepsilon}\to 0\qquad\qquad\mbox{in $L^{2}(\Omega^{r})\;\;$ as $\varepsilon\to 0$},$ (3.36)
-$\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}\to 0\qquad\,\mbox{in$L^{1}(\Omega^{r})\;\;$ as $\varepsilon\to 0$}.$ (3.36)
+$\mbox{\rm curl}\,\mathbf{r}^{(2)}_{\varepsilon}\to 0\qquad\,\mbox{in $L^{1}(\Omega^{r})\;\;$ as $\varepsilon\to 0$}.$ (3.37)
 To show (3.36), we first note that $\Omega^{r}\Subset\Omega^{\rm ext}_{r}$ and $\rho{\mathbf{v}}\in H^{1}(\Omega^{\rm ext}_{r})\cap L^{\infty}(\Omega^{\rm ext}_{r})$ by Lemma 3.3(a), so that
@@
 Also, $(\rho_{\varepsilon}-\rho)(\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon}-\mbox{\rm div}\,{\mathbf{v}})\to 0$ in $L^{2}(\Omega^{r})$, because $\rho-\rho_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$ and $\mbox{\rm div}\,{\mathbf{v}}_{\varepsilon}\to\mbox{\rm div}\,{\mathbf{v}}$ in $L^{2}(\Omega^{r})$.
-Finally, $(\rho_{\varepsilon}-\rho)\mbox{\rm div}\,{\mathbf{v}}\to 0$ in $L^{2}(\Omega^{r})$, because $\rho-\rho_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$ and converge to zero a.e.
-in $\Omega_{r}$, and $\mbox{\rm div}\,{\mathbf{v}}\in L^{2}(\Omega^{r})$, so that
+Finally, $(\rho_{\varepsilon}-\rho)\mbox{\rm div}\,{\mathbf{v}}\to 0$ in $L^{2}(\Omega^{r})$, because $\rho-\rho_{\varepsilon}$ is uniformly bounded in $L^{\infty}(\Omega^{r})$ and converges to zero a.e.
+in $\Omega^{r}$, and $\mbox{\rm div}\,{\mathbf{v}}\in L^{2}(\Omega^{r})$, so that
 $\int_{\Omega_{r}}(\rho_{\varepsilon}-\rho)^{2}(\mbox{\rm div}\,{\mathbf{v}})^{2}\,{\rm d}\xi\to 0\qquad\mbox{as $\varepsilon\to 0$},$
@@
 Now we show (3.37). Note that
-$\mbox{\rm curl}\,{\bf r}^{(2)}_{\varepsilon}=\partial_{\xi_{1}}\big{(}({\mathbf{v}}\cdot\nabla v_{2})_{\varepsilon}-({\mathbf{v}})_{\varepsilon}\cdot\nabla(v_{2})_{\varepsilon}\big{)}-\partial_{\xi_{2}}\big{(}({\mathbf{v}}\cdot\nabla v_{1})_{\varepsilon}-({\mathbf{v}})_{\varepsilon}\cdot\nabla(v_{1})_{\varepsilon}\big{)}.$
-Then (3.36) follows from Lemma A.1 in Appendix A with $p=q=2$, $b=v_{j}$, and $u=\partial_{j}v_{k}$ for the corresponding $j,k=1,2$, and $i=3-k$.
+$\mbox{\rm curl}\,\mathbf{r}^{(2)}_{\varepsilon}=\partial_{\xi_{1}}\big(({\mathbf{v}}\cdot\nabla v_{2})_{\varepsilon}-{\mathbf{v}}_{\varepsilon}\cdot\nabla(v_{2})_{\varepsilon}\big)-\partial_{\xi_{2}}\big(({\mathbf{v}}\cdot\nabla v_{1})_{\varepsilon}-{\mathbf{v}}_{\varepsilon}\cdot\nabla(v_{1})_{\varepsilon}\big).$
+Then (3.37) follows from Lemma A.1 in Appendix A, applied in $\Omega^{\rm ext}_{r}$, with $p=q=2$, $b=v_{j}$, and $u=\partial_{j}v_{k}$ for the corresponding $j,k=1,2$, and i=3ki=3-k.
 Combining the results above, (3.35) is now proved. Then, sending $\varepsilon\to 0$ in (3.32) and using (3.34)–(3.35), we obtain
@@
 $\begin{cases}(\rho v_{1}-\rho_{1}v_{1}^{-})f_{s}^{\prime}-(\rho v_{2}-\rho_{1}v_{2}^{-})=0,\\[2.84526pt] (\rho v_{1}^{2}+p-\rho_{1}(v_{1}^{-})^{2}-p_{1})f^{\prime}_{s}-(\rho v_{1}v_{2}-\rho_{1}v_{1}^{-}v_{2}^{-})=0,\\[2.84526pt] (\rho v_{1}v_{2}-\rho_{1}v_{1}^{-}v_{2}^{-})f_{s}^{\prime}-(\rho v_{2}^{2}+p-\rho_{1}(v_{2}^{-})^{2}-p_{1})=0,\end{cases}$ (3.43)
-where ${\mathbf{v}}^{-}=(v_{1}^{-},v_{2}^{-})$ and ${\mathbf{v}}=(v_{1},v_{2})$.
+where ${\mathbf{v}}^{-}=(v_{1}^{-},v_{2}^{-})=(u_{1}^{-}-\xi_{1},u_{2}^{-}-\xi_{2})$ with some constants $u_{1}^{-}$ and $u_{2}^{-}$, and ${\mathbf{v}}=(v_{1},v_{2})$.
 Taking the tangential derivative $\partial_{\boldsymbol{\tau}}:=\partial_{\xi_{1}}+f_{s}^{\prime}\partial_{\xi_{2}}$ of the Rankine-Hugoniot conditions (3.43) along $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ and using the condition that $f_{s}^{\prime}=0$ at $P$, we obtain that, at point $P$,
@@
 $(a_{1},b_{1},c_{1})=((c^{2}-2v_{1}^{2}-v_{2}^{2})v_{2},\,(c^{2}-v_{2}^{2})v_{1},\,\rho v_{2}^{2}),$
-[…]
+$(a_{2},b_{2},c_{2})=(2(c^{2}-v_{1}^{2})v_{1}v_{2},\,-(c^{2}-v_{2}^{2})(v_{2}^{2}-v_{1}^{2}),\,2\rho v_{1}v_{2}^{2}),$
 $(a_{3},b_{3},c_{3})=(3v_{1}^{2}v_{2}^{2}+v_{1}^{2}c^{2}+v_{2}^{4}-v_{2}^{2}c^{2},\,-2(c^{2}-v_{2}^{2})v_{1}v_{2},\,-2\rho v_{2}^{3}),$
@@
 6. Using Lemma 3.4, we can formally choose $f(s)=s^{2}$ with $g(s)=0$ by (3.9), and use $\zeta\equiv 1$ to obtain that, by (3.12).
-$00=-\int_{\Gamma_{\rm shock}}\big{|}\frac{\omega}{\rho}\big{|}^{2}\rho{\mathbf{v}}\cdot{\boldsymbol{\nu}}\,{\rm d}l>$ 0 .
+$0=-\int_{\Gamma_{\rm shock}}\big|\frac{\omega}{\rho}\big|^{2}\rho{\mathbf{v}}\cdot{\boldsymbol{\nu}}\,{\rm d}l>0.$
 The strict inequality above follows from Definition 3.1(vii) and Lemma 3.5.
@@
 This is a contradiction since $\rho\in[C_{0}^{-1},\,C_{0}]$, ${\mathbf{v}}\cdot{\boldsymbol{\nu}}\leq-C^{-1}$ on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ by Definition 3.1(vii), and $\omega$ (and thus $X$) is continuous and not identically zero on $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ by Lemma 3.5.
-This completes the proof of Theorem 3.2. ∎
+This completes the proof of the fact that, under the assumptions of the theorem, it is not possible that (3.3) holds.
+7. We now complete the proof as claimed. On the contrary, assume that
+${\mathbf{v}}\in H^{1}(\Omega).$ (3.49)
+Then, combining (3.49) with the assumptions in the theorem, we have
+${\mathbf{v}}\in H^{1}(\Omega)\cap L^{\infty}(\Omega),\quad\,\,C^{-1}\leq\rho\leq C,\quad\,\,{\tilde{C}}^{-1}\leq p\leq\tilde{C}$, (3.50)
+where $C,\;\tilde{C}\ \geq 1$, and we recall that $p=\frac{\rho^{\gamma}}{\gamma}$. We show that, with these properties, we can obtain from the weak form of the Euler system (2.6)–(2.7) that
+$\int_{\Omega}\big((\rho{\mathbf{v}}\cdot\nabla{\mathbf{v}}+\rho{\mathbf{v}})\cdot\zeta-p\,\mbox{\rm div}\zeta\big)\,{\rm d}{\boldsymbol{\xi}}=0,$ (3.51)
+for any $\zeta=(\zeta_{1},\zeta_{2})\in C^{1}_{\rm c}(\Omega;\mathbb{R}^{2})$. To show (3.51), we compute for $j=1,2$, using (3.50):
+$\begin{split}\int_{\Omega}\rho\sum_{i=1}^{2}v_{i}v_{j}\partial_{i}\zeta_{j}\,{\rm d}{\boldsymbol{\xi}}&=\int_{\Omega}\rho\sum_{i=1}^{2}v_{i}\partial_{i}(v_{j}\zeta_{j})\,{\rm d}{\boldsymbol{\xi}}-\int_{\Omega}\rho\sum_{i=1}^{2}v_{i}(\partial_{i}v_{j})\zeta_{j}\,{\rm d}{\boldsymbol{\xi}}\\ &=\int_{\Omega}\rho{\mathbf{v}}\cdot\nabla(v_{j}\zeta_{j})\,{\rm d}{\boldsymbol{\xi}}-\int_{\Omega}\rho({\mathbf{v}}\cdot\nabla v_{j})\zeta_{j}\,{\rm d}{\boldsymbol{\xi}}.\end{split}$ (3.52)
+Next, we note that, by the properties in (3.50), we can use the test function $\phi=v_{j}\zeta_{j}$ in (2.6) to obtain
+$\int_{\Lambda}\rho{\mathbf{v}}\cdot\nabla(v_{j}\zeta_{j})\,{\rm d}{\boldsymbol{\xi}}=\int_{\Omega}2\rho v_{j}\zeta_{j}\,{\rm d}{\boldsymbol{\xi}}.$
+Substituting this into the last expression of (3.52) for $j=1,2$, we have
+$\int_{\Omega}\rho{\mathbf{v}}\otimes{\mathbf{v}}:D\zeta\,{\rm d}{\boldsymbol{\xi}}=\int_{\Omega}\big(2\rho v-\rho({\mathbf{v}}\cdot\nabla){\mathbf{v}}\big)\cdot\zeta\,{\rm d}{\boldsymbol{\xi}}.$
+Substituting this into (2.7), we obtain (3.51).
+From (3.51) and the fact that $\rho{\mathbf{v}}\cdot\nabla{\mathbf{v}}+\rho{\mathbf{v}}\in L^{2}(\Omega)$ by (3.50), it follows from (3.51) that the weak derivative of $p$ exists in $\Omega$ and that $\nabla p\in L^{2}(\Omega)$.
+Then, using (3.50), we have
+$p\in H^{1}(\Omega)\cap L^{\infty}(\Omega).$
+From this, since $\rho=(\gamma p)^{1/\gamma}$ and ${\tilde{C}}^{-1}\leq p\leq\tilde{C}$, we obtain
+$\rho\in H^{1}(\Omega)\cap L^{\infty}(\Omega).$
+Combining this with (3.49), we obtain that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$, which contradicts (3.2). Then we conclude that (3.49) is false. ∎
 4. Applications to Transonic Shock problems
@@
 Lemma 4.1.
-Let $(\rho,{\mathbf{v}})$ be a solution of a Riemann problem which satisfies conditions (i)–(vii) of Definition 3.1 and assumption (i) of Theorem 3.2. Assume that
+Let $(\rho,{\mathbf{v}})$ be a solution of the Riemann problem that satisfies conditions (i)–(vii) of Definition 3.1 and assumption (i) of Theorem 3.2. Assume that
   1. (a)
@@
     $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ is not a segment of straight line.
-In particular, denoting by $P_{1}$ and $P_{2}$ the endpoints of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, there exists a point $P^{*}\in\overline{\Gamma_{1}^{int}}\setminus\{P_{1}\}$ such that $\boldsymbol{\tau}(P_{1})\neq\pm\boldsymbol{\tau}(P^{*})$, where $\boldsymbol{\tau}(\cdot)$ is a unit tangent vector to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at a point.
+In particular, denoting by $P_{1}$ and $P_{2}$ the endpoints of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$, there exists a point $P^{*}\in\overline{\Gamma_{1}^{\rm int}}\setminus\{P_{1}\}$ such that $\boldsymbol{\tau}(P_{1})\neq\pm\boldsymbol{\tau}(P^{*})$, where $\boldsymbol{\tau}(\cdot)$ is a unit tangent vector to $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ at a point.
   3. (c)
@@
 Note that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(Q)=0$.
-Moreover, using condition (c), the Rankine-Hugoniot conditions, and the regularity of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ and of $(\rho,{\mathbf{v}})$ given in condition (v) of Definition 3.1 and assumption (i) of Theorem 3.2, we see that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(P_{1})=({\mathbf{v}}\cdot\boldsymbol{\tau}_{S})(P_{1})\neq 0$.
+Moreover, using condition (c), the Rankine-Hugoniot conditions, and the regularity of $\Gamma^{\mbox{\rm\footnotesize int}}_{1}$ and $(\rho,{\mathbf{v}})$ given in condition (v) of Definition 3.1 and assumption (i) of Theorem 3.2, we see that $({\mathbf{v}}_{1}\cdot\boldsymbol{\tau}_{S})(P_{1})=({\mathbf{v}}\cdot\boldsymbol{\tau}_{S})(P_{1})\neq 0$.
 This implies that $Q\neq P_{1}$.
@@
 Then
-$({\mathbf{v}}_{1}\cdot\boldsymbol{\tau})(P)=\frac{-\xi_{1}+(\tilde{v}_{2}-f(\xi_{1}))f^{\prime}(\xi_{1})}{\sqrt{1+(f^{\prime}(\xi_{1}))^{2}}}=\frac{-\xi_{1}+O(\varepsilon)\xi_{1}}{\sqrt{1+(f^{\prime}(\xi_{1}))^{2}}}\neq 0\qquad\mbox{for $\xi_{1}\in(0,\varepsilon)$ if $\varepsilon$ is small.}$.
+[…]
 Thus, ${\mathbf{v}}\cdot\boldsymbol{\tau}={\mathbf{v}}_{1}\cdot\boldsymbol{\tau}\neq 0$ at any $P=(\xi_{1},f(\xi_{1}))$ with $\xi_{1}\in(0,\varepsilon)$.
@@
 The regular reflection problem is a lateral Riemann problem in the region
-[…]
-where +$\mathbb{R}^{2}_{+}=\mathbb{R}^{2}\cap\{x_{1}>0\}$.
+$\Lambda=\mathbb{R}^{2}_{+}\setminus\big\{\mathbf{x}\;:\;x_{1}>0,\,0<x_{2}<x_{1}\tan\theta_{\rm w}\big\},$
+where $\mathbb{R}^{2}_{+}=\mathbb{R}^{2}\cap\{x_{1}>0\}$.
 We seek functions $(\rho,{\mathbf{u}})(\mathbf{x},t)$ satisfying system (2.1) in $\Lambda$ with the boundary condition ${\mathbf{u}}\cdot{\boldsymbol{\nu}}=0$ on $\partial\Lambda$ and the initial data:
@@
 This initial-boundary value problem is invariant under scaling (2.2), so we seek self-similar solutions $(\rho,{\mathbf{v}})=(\rho,{\mathbf{v}})({\boldsymbol{\xi}})$, where the self-similar variables $\boldsymbol{\xi}$} and the pseudo-velocity ${\mathbf{v}}={\mathbf{u}}-{\boldsymbol{\xi}}$ are introduced in §2.1.
-Figure 4.1. Supersonic regular reflection
-Figure 4.2. Subsonic regular reflection
+Figure 2. Supersonic regular reflection
+Figure 3. Subsonic regular reflection
 First, consider the problem with an assumption on the symmetry with respect to the $x_{1}$–axis.
-Then we can consider only the upper half-plane $\{x_{2}>0\}$ and prescribe the slip boundary condition ${\mathbf{u}}\cdot{\boldsymbol{\nu}}=0$ on the symmetry line $\{x_{2}=0\}$, so that there is only one reflection point $P_{0}$ to be considered (see Figs 4.24.2).
+Then we can consider only the upper half-plane $\{x_{2}>0\}$ and prescribe the slip boundary condition ${\mathbf{u}}\cdot{\boldsymbol{\nu}}=0$ on the symmetry line $\{x_{2}=0\}$, so that there is only one reflection point $P_{0}$ to be considered (see Figs 33).
 The regular shock reflection-diffraction configuration is characterized by the fact that the reflection occurs at point $P_{0}$ of the intersection of the incident shock with the wedge boundary.
 Figs.
-4.24.2 show the structure of regular shock reflection-diffraction configurations in self-similar coordinates.
+33 show the structure of regular shock reflection-diffraction configurations in self-similar coordinates.
 A necessary condition for the existence of a regular reflection-diffraction configuration is the existence of the constant state $(2)$ and the reflected shock line such that state $(2)$ satisfies both the slip boundary condition on the wedge and the Rankine-Hugoniot conditions (2.9)–(2.10) on the shock with state $(1)$ across the reflected shock line written at $P_{0}$.
@@
 The type of state $(2)$ at $P_{0}$ for a given wedge angle $\theta_{\rm w}$ determines the type of reflection, supersonic or subsonic, as shown in Fig.
-4.2 and Fig.
-4.2, respectively.
+3 and Fig.
+3, respectively.
 Definition 4.2.
@@
 Next, we define the points and lines in Figs.
-4.24.2.
+33.
 The incident shock $S_{0}$ is line $\{\xi_{1}=\xi_{1}^{0}\}$ with $\xi_{1}^{0}=\frac{\rho_{1}{u}^{(1)}_{1}}{\rho_{1}-\rho_{0}}>0$.
 The center, $O_{2}={\mathbf{u}}^{(2)}=(u_{1}^{(2)},u_{2}^{(2)})$, of the sonic circle $B_{c_{2}}(O_{2})$ of state $(2)$ lies on the wedge boundary between the reflection point $P_{0}$ and the wedge vertex $P_{3}$ for both the supersonic and subsonic cases.
-Then, for the supersonic case, i.e., when $|D\varphi_{2}(P_{0})|=|P_{0}O_{2}|>c_{2}$ so that $P_{0}\notin\overline{B_{c_{2}}(O_{2})}$, we denote by $P_{4}$ the upper point of intersection of $\partial B_{c_{2}}(O_{2})$ with the wedge boundary such that $O_{2}\in P_{3}P_{4}$.
+Then, for the supersonic case, i.e., when $|{\mathbf{v}}_{2}(P_{0})|=|P_{0}O_{2}|>c_{2}$ so that $P_{0}\notin\overline{B_{c_{2}}(O_{2})}$, we denote by $P_{4}$ the upper point of intersection of $\partial B_{c_{2}}(O_{2})$ with the wedge boundary such that $O_{2}\in P_{3}P_{4}$.
 Also, the sonic circle $\partial B_{c_{2}}(O_{2})$ of state $(2)$ intersects line $S_{1}$, and one of the points of intersection, $P_{1}\in\Lambda$, is such that segment $P_{0}P_{1}$ is outside $B_{c_{2}}(O_{2})$.
@@
 Then we denote the line segments $\Gamma_{\rm sym}:=P_{2}P_{3}$ and $\Gamma_{\text{\rm wedge}}:=P_{3}P_{4}$.
-The lines and curves $\Gamma_{\text{\rm shock}}$, $\Gamma_{\text{\rm sonic}}$, $\Gamma_{\rm sym}$, and $\Gamma_{\text{\rm wedge}}$ do not have common points, except their endpoints $P_{1},\cdots,P_{4}$.
+The lines and curves $\Gamma_{\text{\rm shock}}$, $\Gamma_{\text{\rm sonic}}$, $\Gamma_{\rm sym}$, and $\Gamma_{\text{\rm wedge}}$ do not have common points, except for their endpoints $P_{1},\cdots,P_{4}$.
 Thus, $\Gamma_{\text{\rm shock}}\cup\Gamma_{\text{\rm sonic}}\cup\Gamma_{\rm sym}\cup\Gamma_{\text{\rm wedge}}$ is a closed curve without self-intersection.
 Denote by $\Omega$ the bounded domain restricted by this curve.
-For the subsonic/sonic case, i.e., when $|D\varphi_{2}(P_{0})|=|P_{0}O_{2}|\leq c_{2}$ so that $P_{0}\in\overline{B_{c_{2}}(O_{2})}$, the curved reflected-diffracted shock is $\Gamma_{\text{\rm shock}}=P_{0}P_{2}$, which does not have common interior points with the line segments $\Gamma_{\rm sym}=P_{2}P_{3}$ and $\Gamma_{\text{\rm wedge}}=P_{0}P_{3}$.
+For the subsonic/sonic case, i.e., when $|{\mathbf{v}}_{2}(P_{0})|=|P_{0}O_{2}|\leq c_{2}$ so that $P_{0}\in\overline{B_{c_{2}}(O_{2})}$, the curved reflected-diffracted shock is $\Gamma_{\text{\rm shock}}=P_{0}P_{2}$, which does not have common interior points with the line segments $\Gamma_{\rm sym}=P_{2}P_{3}$ and $\Gamma_{\text{\rm wedge}}=P_{0}P_{3}$.
 Then $\Gamma_{\text{\rm shock}}\cup\Gamma_{\rm sym}\cup\Gamma_{\text{\rm wedge}}$ is a closed curve without self-intersection, and $\Omega$ is the bounded domain restricted by this curve.
@@
 The intuition for the definition is the following: The regular shock reflection solution is an entropy solution of the regular shock reflection problem in the sense of Definition 4.2 which has the structure shown in Figs.
-4.24.2, where $(\rho,{\mathbf{v}})$ coincides with those of states $(0)$, $(1)$, and $(2)$ in their respective regions.
+33, where $(\rho,{\mathbf{v}})$ coincides with those of states $(0)$, $(1)$, and $(2)$ in their respective regions.
 As we discussed above, the necessary condition is the existence of state $(2)$, which means that the wedge angle satisfies $\theta_{\rm w}\in(\theta_{\rm w}^{\rm d},\frac{\pi}{2})$.
 Moreover, it is expected that the solution is relatively regular in $\Omega$.
-However, we show below that it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
-On the other hand, from the physical/computational experiments and the theoretical results in the case of potential flow, it is expected that $\Gamma_{\rm shock}$ is a smooth curve and $(\rho,{\mathbf{v}})$ is smooth near and up to $\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}$ in $\Omega$.
+However, we show below that it is not possible that ${\mathbf{v}}\in H^{1}(\Omega)$.
+On the other hand, from the physical/computational experiments and the theoretical results in the case of potential flow, it is expected that $\Gamma_{\rm shock}$ is a smooth curve and $\mathbf{v}$} is smooth near and up to $\Gamma_{\rm shock}\cup\Gamma_{\rm sonic}$ in $\Omega$.
 Moreover, the regularity discussed is expected for the shock reflection-diffraction configuration extended to $\{\xi_{1}<0\}$ by the even reflection about the $\xi_{1}$–axis (since this is the original shock reflection-diffraction configuration).
@@
 If state $(2)$ for $\theta_{\rm w}$ is supersonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$, the solution has the supersonic reflection structure as on Fig.
-4.2.
+3.
 If state $(2)$ for $\theta_{\rm w}$ is subsonic or sonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|\leq c_{2}$, the solution has the subsonic reflection structure as on Fig.
-4.2.
+3.
 More specifically,
@@
   2. (ii)
-    $(\rho,{\mathbf{v}})$ is continuous in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})$ for the supersonic or sonic reflections, and in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock})$ for the subsonic reflection, for some $r>0$.
+    $(\rho,{\mathbf{v}})$ is continuous in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})$ for the supersonic reflection, and in $\overline{\Omega}\cap{\mathcal{N}}_{r}(\Gamma_{\rm shock})$ for the subsonic or sonic reflection, for some $r>0$.
   3. (iii)
     The solution coincides with states $(0)$, $(1)$, and $(2)$ in their respective regions: for the supersonic reflection case,
-    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for $\xi_{1}>\xi_{1}^{0}$ and $\xi_{2}>\xi_{1}\tan\theta_{\rm w}$},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for $\xi_{1}<\xi_{1}^{0}$ and above curve $P_{0}P_{1}P_{2}$},\\[2.84526pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in $P_{0}P_{1}P_{4}$},\end{cases}$
+[…]
     where $\xi_{1}^{0}>0$ is the location of the incident shock $S_{0}$ on the self-similar plane; and for the subsonic or sonic reflection case,
-    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for $\xi_{1}>\xi_{1}^{0}$ and $\xi_{2}>\xi_{1}\tan\theta_{\rm w}$},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for $\xi_{1}<\xi_{1}^{0}$ and above curve $P_{0}P_{2}$},\end{cases}$
+[…]
     and $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{0}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{2},{\mathbf{v}}_{2})(P_{0})$.
@@
   5. (v)
-    The flow is pseudo-subsonic in $\Omega$ on and near $\Gamma_{\rm shock}$, except the except the sonic point $P_{1}$ for the supersonic or sonic reflections, i.e., $|{\mathbf{v}}|<c$ on $\overline{\Gamma}_{\rm shock}\setminus\{P_{1}\}$ for supersonic and sonic reflections, and on $\overline{\Gamma}_{\rm shock}$ for strictly subsonic reflections, where ${\mathbf{v}}$ on $\Gamma_{\rm shock}$ is computed from the $\Omega$–side.
+    The flow is pseudo-subsonic in $\Omega$ on and near $\Gamma_{\rm shock}$, except for the sonic point $P_{1}$ for the supersonic or sonic reflections, i.e., $|{\mathbf{v}}|<c$ on $\overline{\Gamma_{\rm shock}}\setminus\{P_{1}\}$ for supersonic and sonic reflections, and on $\overline{\Gamma_{\rm shock}}$ for strictly subsonic reflections, where $\mathbf{v}$} on $\Gamma_{\rm shock}$ is computed from the $\Omega$–side.
 Remark 4.4.
@@
 Remark 4.5.
-For the supersonic reflections, it follows from conditions (ii)–(iii) of Definition 4.3 and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ are continuous across the sonic arc $P_{1}P_{4}=\overline{\Gamma_{\rm sonic}}$:
+For the supersonic reflections, it follows from conditions (ii)–(iii) of Definition 4.3 and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ is continuous across the sonic arc $P_{1}P_{4}=\overline{\Gamma_{\rm sonic}}$:
 $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{2},{\mathbf{v}}_{2}(P))\qquad\mbox{for any $P\in\overline{\Gamma_{\rm sonic}}$}.$ (4.4)
 Theorem 4.6.
-Let $(\rho,{\mathbf{v}})$ be a regular shock reflection solution in the sense of Definition 4.3. Assume that $(\rho,{\mathbf{v}})$ satisfy the following:
+Let $(\rho,{\mathbf{v}})$ be a regular shock reflection solution in the sense of Definition 4.3. Assume that $(\rho,{\mathbf{v}})$ satisfies the following:
   1. (i)
@@
     $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
-Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
+Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
 The proof of Theorem 4.6 will be given for the more general non-symmetric case below (Theorem 4.10), based on the observation in Remark 4.8.
@@
 As two examples of them, Fig.
-4.4 and Fig.
-4.4 exhibit the structures of two supersonic regular reflections and two subsonic regular reflections, respectively.
+5 and Fig.
+5 exhibit the structures of two supersonic regular reflections and two subsonic regular reflections, respectively.
 In light of Definition 4.3, we define the regular reflection solution for the non-symmetric case.
@@
 Let $\xi_{1}^{0}>0$ be the location of the incident shock $S_{0}$ on the self-similar plane.
-Figure 4.3. Non-symmetric supersonic regular reflection
-Figure 4.4. Non-symmetric subsonic regular reflection
+Figure 4. Non-symmetric supersonic regular reflection
+Figure 5. Non-symmetric subsonic regular reflection
 We call $(\rho,{\mathbf{v}})\in L^{\infty}(\Lambda)$, with $\rho\in BV_{\rm loc}(\Lambda\cap{\mathcal{N}}_{r}(\partial\Lambda))$ for some $r>0$, is an entropy solution of the non-symmetric regular shock reflection problem if $(\rho,{\mathbf{v}})$ is an entropy solution of system (2.3)–(2.4) with the slip boundary condition (2.13) in the sense of Definition 2.2, which satisfies the asymptotic conditions:
@@
 where
-$(\bar{\rho},\bar{{\mathbf{v}}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for $\xi_{1}>\xi_{1}^{0}$},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for $\xi_{1}<\xi_{1}^{0}$}.\end{cases}$
+[…]
 Then we can define the non-symmetric regular shock reflection solution.
@@
     If state $(2)$ at $P_{0}$ and state $(3)$ at $P_{1}$ are both supersonic, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$ and $|{\mathbf{v}}_{2}(P_{1})|>c_{3}$, the solution has the supersonic reflection structure as in Fig.
-4.4 at both points $P_{0}$ and $P_{1}$.
+5 at both points $P_{0}$ and $P_{1}$.
 If state $(2)$ at $P_{0}$ and state $(3)$ at $P_{1}$ are both subsonic or sonic, i.e., $|{\mathbf{v}}_{2}(P_{0})|\leq c_{2}$ and $|{\mathbf{v}}_{2}(P_{1})|\leq c_{3}$, the solution has the subsonic reflection structure as in Fig.
-4.4.
+5.
   2. (ii)
@@
     The solution coincides with states $(0)$, $(1)$, $(2)$, and $(3)$ in their respective regions:
-    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for}\,\,\,\xi_{1}>\xi_{1}^{0},\,\,\,\xi_{2}>\xi_{1}\tan\theta_{\rm w}^{1},\\[2.84526pt] (\rho_{0},{\mathbf{v}}_{0})\qquad\mbox{for}\,\,\,\xi_{1}>\xi_{1}^{0},\,\,\,\xi_{2}<-\xi_{1}\tan\theta_{\rm w}^{2},\\[2.84526pt] (\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{for}\,\,\,\xi_{1}<\xi_{1}^{0},\,\,\,\mbox{and the left to curve }\;P_{0}P_{1},\\[2.84526pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in}\,\,\,\Lambda_{2}\,\,\,\mbox{if state $(2)$ at $P_{0}$ is supersonic},\\[2.84526pt] (\rho_{3},{\mathbf{v}}_{3})\qquad\mbox{in}\,\,\,\Lambda_{3}\,\,\,\mbox{if state $(3)$ at $P_{1}$ is supersonic},\end{cases}$
-    where $\xi_{1}^{0}>0$ is the location of the incident shock $S_{0}$ on the self-similar plane, and
-    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{0}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{2},{\mathbf{v}}_{2})(P_{0})\qquad\mbox{if state $(2)$ at $P_{0}$ is subsonic or sonic},$
-    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{1}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{3},{\mathbf{v}}_{3})(P_{1})\qquad\mbox{if state $(3)$ at $P_{1}$ is subsonic or sonic}.$
+[…]
+    where $\xi_{1}^{0}>0$ is the location of the incident shock $S_{0}$ on the self-similar plane, and limξΩ, ξP0(ρ, v)(ξ)=(ρ2, v2)(P0)  if state $(2)$ at $P_{0}$ is subsonic or sonic, limξΩ, ξP1(ρ, v)(ξ)=(ρ3, v3)(P1)  if state $(3)$ at $P_{1}$ is subsonic or sonic.
   5. (v)
@@
 Remark 4.9.
-For the supersonic reflections, it follows from conditions (ii)–(iv) of Definition 4.7 and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ are continuous across the sonic arcs $\Gamma_{\rm sonic}^{2}\cup\Gamma_{\rm sonic}^{3}$:
+For the supersonic reflections, it follows from conditions (ii)–(iv) of Definition 4.7 and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ is continuous across the sonic arcs $\Gamma_{\rm sonic}^{2}\cup\Gamma_{\rm sonic}^{3}$:
 $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{2},{\mathbf{v}}_{i})(P)\qquad\mbox{for any }\;P\in\overline{\Gamma_{\rm sonic}^{i}}\quad\mbox{if }\;|{\mathbf{v}}_{i}(P_{0})|>c_{i}\;\;\mbox{for }\;i=2,3.$
 Theorem 4.10.
-Let $(\rho,{\mathbf{v}})$ be a non-symmetric regular shock reflection solution in the sense of Definition 4.7. Assume that $(\rho,{\mathbf{v}})$ satisfy the following:
+Let $(\rho,{\mathbf{v}})$ be a non-symmetric regular shock reflection solution in the sense of Definition 4.7. Assume that $(\rho,{\mathbf{v}})$ satisfies the following:
   1. (i)
@@
     $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
-Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
+Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
 Proof.
-It suffices to check that the assumptions of Theorem 3.2 are satisfied. We divide the proof into four steps.
+It suffices to check that the assumptions of Theorem 3.2 are satisfied. We divide the proof into four steps:
 1.
 Since the cases with one supersonic reflection and one subsonic reflection can be treated similarly, we focus only on the non-symmetric supersonic reflection solution (see Fig.
-4.4) and the non-symmetric subsonic reflection solution (see Fig.
-4.4) in the proof below.
+5) and the non-symmetric subsonic reflection solution (see Fig.
+5) in the proof below.
 2. We first show that the non-symmetric regular reflection solutions satisfy conditions (i)–(vii) of Definition 3.1 and assumptions (i)–(iii) of Theorem 3.2.
@@
 Using the boundary conditions that ${\mathbf{v}}={\mathbf{v}}_{2}$ on $\Gamma_{\rm sonic}^{2}$ and ${\mathbf{v}}={\mathbf{v}}_{3}$ on $\Gamma_{\rm sonic}^{3}$ by Definition 4.7(iv), we obtain that
-${\mathbf{v}}\cdot{\boldsymbol{\nu}}=-c_{2}<0\qquad\mbox{on $\Gamma_{\rm sonic}^{2}$ and ${\mathbf{v}}\cdot{\boldsymbol{\nu}}=-c_{3}<0$ on $\Gamma_{\rm sonic}^{3}$}.$
+[…]
 Thus, the requirements of Definition 3.1(vii) hold.
@@
 Condition (a) is satisfied since $\Lambda_{1}$ is the region of state $(1)$ as shown in Figs.
-4.44.4 above, and $\Gamma_{\rm shock}=\partial\Lambda_{1}\cap\partial\Omega$ in both the supersonic and subsonic cases.
+55 above, and $\Gamma_{\rm shock}=\partial\Lambda_{1}\cap\partial\Omega$ in both the supersonic and subsonic cases.
 We now check condition (b) of Lemma 4.1.
@@
 Consider first the case of reflections that are supersonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|>c_{2}$; see Fig.
-4.4.
+5.
 Then we need to show that $({\mathbf{v}}\cdot{\boldsymbol{\tau}}_{\rm shock})(P_{2})\neq 0$.
@@
 In the case of reflections that are sonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|=c_{2}$ (see Fig.
-4.4), the argument is the same as above with only notational change: we use point $P_{0}$ here instead of point $P_{2}$.
+5), the argument is the same as above with only notational change: we use point $P_{0}$ here instead of point $P_{2}$.
 It remains to consider the case of reflections that are subsonic at $P_{0}$, i.e., $|{\mathbf{v}}_{2}(P_{0})|<c_{2}$; see Fig.
-4.4.
+5.
 Assume that $({\mathbf{v}}\cdot\boldsymbol{\tau}_{\rm shock})(P_{0})=0$.
@@
 See Figs.
-4.64.6; also see Bae-Chen-Feldman [2, 3] and Elling-Liu [26].
-Figure 4.5. Prandtl supersonic shock reflection
-Figure 4.6. Prandtl subsonic shock reflection
+77; also see Bae-Chen-Feldman [2, 3] and Elling-Liu [26].
+Figure 6. Prandtl supersonic shock reflection
+Figure 7. Prandtl subsonic shock reflection
 Consider the problem in the self-similar coordinates ${\boldsymbol{\xi}}=(\xi_{1},\xi_{2})$ in the region:
@@
 First, a similar argument as made for the regular shock reflection at the reflection point $P_{0}$ (see Figs.
-4.24.2) yields that, at wedge-vertex $P_{1}$, for a given uniform incoming flow $(\rho_{\infty},{\mathbf{v}}_{\infty})$, there is a detachment angle $\theta_{\rm w}^{\rm d}\in(0,\frac{\pi}{2})$ such that the system of algebraic equations (2.9) and (2.13) for state $(O)$ has two solutions for each wedge-angle $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$ such that the entropy condition is satisfied for the two-shock configuration.
+33) yields that, at wedge-vertex $P_{1}$, for a given uniform incoming flow $(\rho_{\infty},{\mathbf{v}}_{\infty})$, there is a detachment angle $\theta_{\rm w}^{\rm d}\in(0,\frac{\pi}{2})$ such that the system of algebraic equations (2.9) and (2.13) for state $(O)$ has two solutions for each wedge-angle $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$ such that the entropy condition is satisfied for the two-shock configuration.
 The weak state $(O)$ with the smaller density is expected to be physical.
@@
 Similarly, it follows from the Rankine-Hugoniot conditions (2.9) and the slip boundary condition (2.13) that
-The shock line $\Gamma_{\rm shock}^{O}$ between states $(O)$ and ()($\infty$)( ∞ ) is not parallel to $\Gamma_{\rm wedge}$ for all $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$. (4.5)
+The shock line $\Gamma_{\rm shock}^{O}$ between states $(O)$ and $(\infty)$ is not parallel to $\Gamma_{\rm wedge}$ for all $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$. (4.5)
 Depending on the wedge-angle, state $(O)$ can be either supersonic or subsonic at $P_{1}$.
 It determines the type of the reflection, supersonic or subsonic, as shown in Fig.
-4.6 and Fig.
-4.6, respectively.
-Second, by a straightforward computation, we know that there exists a unique constant state $(N)$, which determines the normal reflection of state ()($\infty$)( ∞ ) from the wedge boundary $\Gamma_{\rm wedge}$ so that state $(N)$ satisfies the slip boundary condition (2.13) along $\Gamma_{\rm wedge}$ and the Rankine-Hugoniot conditions (2.9) along a straight line $\Gamma_{\rm shock}^{N}$, which lies in $\Lambda$ and is parallel to $\Gamma_{\rm wedge}$.
+7 and Fig.
+7, respectively.
+Second, by a straightforward computation, we know that there exists a unique constant state $(N)$, which determines the normal reflection of state $(\infty)$ from the wedge boundary $\Gamma_{\rm wedge}$ so that state $(N)$ satisfies the slip boundary condition (2.13) along $\Gamma_{\rm wedge}$ and the Rankine-Hugoniot conditions (2.9) along a straight line $\Gamma_{\rm shock}^{N}$, which lies in $\Lambda$ and is parallel to $\Gamma_{\rm wedge}$.
 Hinted by the solution structures given in [2, 3, 26] for the potential flow, for any given wedge-angle $\theta_{\rm w}\in(0,\theta_{\rm w}^{\rm d})$, an entropy solution $(\rho,{\mathbf{v}})$ of the boundary value problem in the sense of Definition 2.2 is called a regular Prandtl-Meyer reflection solution for the isentropic Euler system (2.3)–(2.4) if $(\rho,{\mathbf{v}})$ satisfies the following further properties:
@@
     If state $(O)$ at $P_{1}$ is supersonic, i.e., $|{\mathbf{v}}_{O}(P_{1})|>c_{O}$, the solution has the supersonic reflection structure as in Fig.
-4.6 at point $P_{1}$.
+7 at point $P_{1}$.
 If state $(O)$ at $P_{1}$ is subsonic or sonic, i.e., $|{\mathbf{v}}_{O}(P_{1})|\leq c_{O}$, the solution has the subsonic reflection structure as in Fig.
-4.6.
+7.
   2. (ii)
@@
   4. (iv)
-    The solution coincides with states ()($\infty$)( ∞ ), $(O)$, and $(N)$ in their respective regions. Specifically,
+    The solution coincides with states $(\infty)$, $(O)$, and $(N)$ in their respective regions. Specifically,
     $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{\infty},{\mathbf{v}}_{\infty})\qquad\,\,&\mbox{in}\,\,\,\Lambda_{\infty},\\[1.42262pt] (\rho_{N},{\mathbf{v}}_{N})\qquad\,\,&\mbox{in}\,\,\,\Lambda_{N},\\[1.42262pt] (\rho_{O},{\mathbf{v}}_{O})\qquad\,\,&\mbox{in}\,\,\,\Lambda_{O}\,\,\,\,\mbox{if state (O) at $P_{1}$ is supersonic},\end{cases}$
-    and
-    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{1}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{O},{\mathbf{v}}_{O})(P_{1})\qquad\mbox{if state (O) at $P_{1}$ is subsonic or sonic}.$
+    and limξ∈Ω, ξ→P1(ρ, v)(ξ)=(ρO, vO)(P1)  if state (O) at $P_{1}$ is subsonic or sonic.
   5. (v)
@@
     The flow is pseudo-subsonic in $\Omega$ on and near $\Gamma_{\rm shock}$, except for the sonic points $P_{4}$ and $P_{5}$ if they exist, or point $P_{1}$ if state $(O)$ is sonic at $P_{1}$.
-We remark that, for the supersonic reflections, it follows from (ii)–(iv) of the definition above and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ are continuous across the sonic arcs $\Gamma_{\rm sonic}^{N}$ and $\Gamma_{\rm sonic}^{O}$.
+We remark that, for the supersonic reflections, it follows from (ii)–(iv) of the definition above and the Rankine-Hugoniot conditions that $(\rho,{\mathbf{v}})$ is continuous across the sonic arcs $\Gamma_{\rm sonic}^{N}$ and $\Gamma_{\rm sonic}^{O}$.
 Then, following the arguments as for the proof of Theorem 4.10, we have the following theorem on the lower regularity of the regular Prandtl-Meyer shock reflection solutions:
@@
     $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
-Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
+Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
 Proof.
@@
 Let us first consider the supersonic Prandtl reflection case; see Fig.
-4.6.
+7.
 Arguing as in the proof of Theorem 4.10, we show that, at point $P_{4}$ (resp.
@@
 In the case of subsonic or sonic Prandtl reflection as in Fig.
-4.6, we argue similarly, except we consider point $P_{1}$ instead of point $P_{5}$ and use the fact that, in the subsonic or sonic case, $\Gamma_{\rm shock}$ is parallel to $\Gamma_{\rm shock}^{O}$ at $P_{1}$ by the last equality in condition (iv) of the definition of admissible solutions.
+7, we argue similarly, except that we consider point $P_{1}$ instead of point $P_{5}$ and use the fact that, in the subsonic or sonic case, $\Gamma_{\rm shock}$ is parallel to $\Gamma_{\rm shock}^{O}$ at $P_{1}$ by the last equality in condition (iv) of the definition of admissible solutions.
 Then we conclude in the same way as in the case of supersonic Prandtl reflection.
@@
 As discussed in [16] for the potential flow and shown in Fig.
-4.7, the Lighthill shock diffraction problem arises as a straight incident shock passes through a wedge stepping down.
-Figure 4.7. Lighthill shock diffraction problem
+8, the Lighthill shock diffraction problem arises as a straight incident shock passes through a wedge stepping down.
+Figure 8. Lighthill shock diffraction problem
 Initially, we consider two piecewise constant Riemann data with the left state $(1)$: $(\rho_{1},u_{1}^{(1)},0)$ for $u_{1}^{(1)}>0$ and the right state $(0)$: $(\rho_{0},0,0)$, separated by a vertical shock, which hits the wedge at the wedge-corner $P_{1}$.
@@
 Thus, as shown in Fig.
-4.7, the incident shock $\Gamma_{\rm shock}^{1}$ interacts with the sonic circle $\Gamma_{\rm sonic}$ of state $(1)$ and becomes a transonic shock $\Gamma_{\rm shock}$, and then $\Gamma_{\rm shock}$ meets wedge $\Gamma_{\rm wedge}^{0}$ perpendicularly.
+8, the incident shock $\Gamma_{\rm shock}^{1}$ interacts with the sonic circle $\Gamma_{\rm sonic}$ of state $(1)$ and becomes a transonic shock $\Gamma_{\rm shock}$, and then $\Gamma_{\rm shock}$ meets wedge $\Gamma_{\rm wedge}^{0}$ perpendicularly.
 Let $\Lambda_{0}$, $\Lambda_{1}$, and $\Omega$ be defined, as shown in Fig.
-4.7.
+8.
 Let $\Lambda=\overline{\Lambda_{0}\cup\Lambda_{1}\cup\Omega}$.
@@
   1. (i)
-    $\rho,{\mathbf{v}}\in C^{1}\left(({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega})\setminus\{P_{3}\}\right)\cap C^{0,1}({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega})$ for some $\sigma>0$;
+    $\rho,{\mathbf{v}}\in C^{1}\left(({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega})\setminus\{P_{3}\}\right)\cap C^{0,1}\left({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}\cup\Gamma_{\rm sonic})\cap\overline{\Omega}\right)$ for some $\sigma>0$;
   2. (ii)
     $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
-Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
+Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
 Since the proof argument is similar to the one for Theorem 4.11, we omit the details.
@@
 As introduced in [12], under the symmetry assumption that $\theta_{12}=\theta_{14}=\theta_{1}$, $\theta_{32}=\theta_{34}=\theta_{2}$, and states in $\Lambda_{2}$ and $\Lambda_{4}$ are the same constants, and under a structure assumption that one forward shock is generated between states $(1)$ and $(j)$ and one backward shock is generated between states $(3)$ and $(j)$, governed by the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12), we expect the Riemann problem develops a Riemann solution with the structure as shown in Fig.
-4.8.
-Figure 4.8. Riemann problem with four-shock interactions
+9.
+Figure 9. Riemann problem with four-shock interactions
 On the symmetry line that is the dashed line shown in Fig.
-4.8, the velocity satisfies (2.13) with $\boldsymbol{\nu}$ being the vertical direction.
+9, the velocity satisfies (2.13) with $\boldsymbol{\nu}$ being the vertical direction.
 At the interaction points $P_{1}$ and $P_{4}$, based on an argument similar to the one for the regular shock reflection problem at the reflection point $P_{0}$, with the symmetry line and state $(2)$ in Fig.
-4.8 corresponding to $\Gamma_{\rm wedge}$ and state $(1)$ in Fig.
-4.2, there exists a detachment angle $\theta^{\rm d}\in(0,\frac{\pi}{2})$ depending on the data such that, for $\theta_{1},\theta_{2}\in(0,\theta_{\rm w}^{\rm d})$, there exist two constant states $(5)$ and $(6)$ in $\Lambda_{5}$ and $\Lambda_{6}$, respectively, satisfying the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12).
+9 corresponding to $\Gamma_{\rm wedge}$ and state $(1)$ in Fig.
+3, there exists a detachment angle $\theta^{\rm d}\in(0,\frac{\pi}{2})$ depending on the data such that, for $\theta_{1},\theta_{2}\in(0,\theta_{\rm w}^{\rm d})$, there exist two constant states $(5)$ and $(6)$ in $\Lambda_{5}$ and $\Lambda_{6}$, respectively, satisfying the Rankine-Hugoniot conditions (2.9) and the entropy condition (2.12).
 One corresponds to the strong shock and the other to the weak shock.
@@
 It is direct to see that
-The tangents to Γshock1 at P2 and P3 are not parallel to each $\mbox{\it The tangents to $\Gamma_{\rm shock}^{1}$ at $P_{2}$ and $P_{3}$ are not parallel to each other},$ (4.6)
+[…]
 where $P_{2}$ is replaced by $P_{1}$ if state $(6)$ is subsonic and $P_{3}$ is replaced by $P_{4}$ if state $(5)$ is subsonic.
 Let $\Lambda_{i}$ for $i=1,\cdots,6$, if they exist, and let $\Omega$ be defined as shown in Fig.
-4.8.
+9.
 Let $\Lambda=\mathbb{R}^{2}$.
@@
     The solution coincides with the corresponding constant states in their respective regions $\Lambda_{i}$, for $i=1,\cdots,6$, if they exist: Specifically,
-    $(\rho,{\mathbf{v}})=\begin{cases}(\rho_{1},{\mathbf{v}}_{1})\qquad\mbox{in}\,\,\,\Lambda_{1},\\[1.42262pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in}\,\,\,\Lambda_{2},\\[1.42262pt] (\rho_{3},{\mathbf{v}}_{3})\qquad\mbox{in}\,\,\,\Lambda_{3},\\[1.42262pt] (\rho_{2},{\mathbf{v}}_{2})\qquad\mbox{in}\,\,\,\Lambda_{4},\\[1.42262pt] (\rho_{5},{\mathbf{v}}_{5})\qquad\mbox{in}\,\,\,\Lambda_{5},\quad\mbox{if state (5) at $P_{4}$ is supersonic}\\[1.42262pt] (\rho_{6},{\mathbf{v}}_{6})\qquad\mbox{in}\,\,\,\Lambda_{6},\quad\mbox{if state (6) at $P_{1}$ is supersonic},\end{cases}$
-    where
-    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{4}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{5},{\mathbf{v}}_{5})(P_{4})\qquad\mbox{if state (5) at $P_{4}$ is subsonic or sonic},$
-    $\lim_{{\boldsymbol{\xi}}\in\Omega,\;{\boldsymbol{\xi}}\to P_{1}}(\rho,{\mathbf{v}})({\boldsymbol{\xi}})=(\rho_{6},{\mathbf{v}}_{6})(P_{1})\qquad\mbox{if state (6) at $P_{1}$ is subsonic or sonic}.$
+[…]
+    where limξ∈Ω, ξ→P4(ρ, v)(ξ)=(ρ5, v5)(P4)  if state (5) at $P_{4}$ is subsonic or sonic, limξ∈Ω, ξ→P1(ρ, v)(ξ)=(ρ6, v6)(P1)  if state (6) at $P_{1}$ is subsonic or sonic.
   4. (iv)
@@
 If state $(5)$ at $P_{4}$ and state $(6)$ at $P_{1}$ are both supersonic, then $M=6$, ${N_{1}}=0$, and ${N_{2}}=4$ with $\Lambda_{i}$ (for $i=1,\cdots,6$) being given as in Fig.
-4.8, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{2}P_{3}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{5}P_{6}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{1}=P_{2}P_{6}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{4}=\Gamma_{\rm sonic}^{2}=P_{3}P_{5}$.
+9, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{2}P_{3}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{5}P_{6}$, $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{1}=P_{2}P_{6}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{4}=\Gamma_{\rm sonic}^{2}=P_{3}P_{5}$.
 If state $(5)$ at $P_{4}$ and state $(6)$ at $P_{1}$ are both subsonic, then $M=4$, ${N_{1}}=0$, and ${N_{2}}=2$ with $\Lambda_{i}$ (for $i=1,2,3,4$) being given as in Fig.
-4.8, $\Lambda_{5}$ and $\Lambda_{6}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{1}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{4}P_{1}$.
+9, $\Lambda_{5}$ and $\Lambda_{6}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{1}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{4}P_{1}$.
 If state $(5)$ at $P_{4}$ is supersonic and state $(6)$ at $P_{1}$ is subsonic, then $M=5$, ${N_{1}}=0$, and ${N_{2}}=2$ with $\Lambda_{i}$ (for $i=1,\cdots,5$) being given as in Fig.
-4.8, $\Lambda_{6}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{1}P_{3}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{5}P_{1}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{2}=P_{3}P_{5}$.
+9, $\Lambda_{6}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{1}P_{3}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{5}P_{1}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{2}=P_{3}P_{5}$.
 It state $(5)$ at $P_{4}$ is subsonic and state $(6)$ at $P_{1}$ is supersonic, then $M=5$, ${N_{1}}=0$ and ${N_{2}}=3$ with $\Lambda_{i}$ (for $i=1,\cdots,6$) being given as in Fig.
-4.8, $\Lambda_{5}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{2}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{4}P_{6}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{1}=P_{2}P_{6}$.
+9, $\Lambda_{5}$ disappearing, $\Gamma^{\mbox{\rm\footnotesize int}}_{1}=\Gamma_{\rm shock}^{1}=P_{2}P_{4}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{2}=\Gamma_{\rm shock}^{2}=P_{4}P_{6}$, and $\Gamma^{\mbox{\rm\footnotesize int}}_{3}=\Gamma_{\rm sonic}^{1}=P_{2}P_{6}$.
 Then, following the proof of Theorem 4.10, we have the following theorem on the lower regularity of the regular Riemann solution, whose proof is omitted since it is similar to the one for Theorem 4.10:
@@
   1. (i)
-    $(\rho,{\mathbf{v}})\in C^{1}\big{(}({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega})\setminus\{P_{2},P_{3},P_{5},P_{6}\}\big{)}\cap C^{0,1}\big{(}{\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega}\big{)}$ for some $\sigma>0$, where $\Gamma_{\rm sonic}^{i}=\emptyset$ when it does not exist for $i=1,2$;
+    For some $\sigma>0$,
+    $(\rho,{\mathbf{v}})\in$ $\,C^{1}\big(({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega})\setminus\{P_{2},P_{3},P_{5},P_{6}\}\big)$
+    $\,\cap C^{0,1}\big({\mathcal{N}}_{\sigma}(\Gamma_{\rm shock}^{1}\cup\Gamma_{\rm shock}^{2}\cup\Gamma_{\rm sonic}^{1}\cup\Gamma_{\rm sonic}^{2})\cap\overline{\Omega}\big),$
+    where $\Gamma_{\rm sonic}^{i}=\emptyset$ when it does not exist for $i=1,2$;
   2. (ii)
     $|{\mathbf{v}}|\leq C_{0}$ and $C_{0}^{-1}\leq\rho\leq C_{0}$ in $\Omega$ for some $C_{0}\geq 1$.
-Then it is not possible that $(\rho,{\mathbf{v}})\in H^{1}(\Omega)$.
+Then ${\mathbf{v}}\notin H^{1}(\Omega)$.
 Appendix A DiPerna-Lions-Type Commutator Estimates
@@
 $=\int_{B_{M}\times B_{M}}\eta(\mathbf{y})\eta_{x_{i}}(\mathbf{z})\frac{b(\mathbf{x}-\varepsilon\mathbf{y})-b(\mathbf{x})}{\varepsilon}u(\mathbf{x}-\varepsilon\mathbf{z})\,{\rm d}\mathbf{y}{\rm d}\mathbf{z}.$
-Then we complete the proof of (A.3) for $I_{\varepsilon,4}$ similar as for $I_{\varepsilon,1}$ above, by using the $L^{p}$–estimate the difference quotient for the $b$-term, and the $L^{q}$–estimate of the shift for the $u$-term.
+Then we complete the proof of (A.3) for $I_{\varepsilon,4}$ similar as for $I_{\varepsilon,1}$ above, by using the $L^{p}$–estimate of the difference quotient for the $b$-term, and the $L^{q}$–estimate of the shift for the $u$-term.
 Then estimate (A.2) holds.
@@
 The research of Gui-Qiang G.
-Chen was also supported in part by the UK Engineering and Physical Sciences Research Council Award EP/L015811/1, EP/V008854, and EP/V051121/1.
+Chen was supported in part by the UK Engineering and Physical Sciences Research Council Awards EP/L015811/1, EP/V008854, and EP/V051121/1.
 The research of Mikhail Feldman was supported in part by the National Science Foundation under Grants DMS-2054689 and DMS-2219391, and the Steenbock Professorship Award from the University of Wisconsin-Madison.
 The research of Wei Xiang was supported in part by the Research Grants Council of the HKSAR, China (Project No.
-CityU 11304820, CityU 11300021, CityU 11311722, and CityU 11305523), and in part by the Research Center for Nonlinear Analysis of the Hong Kong Polytechnic University.
+CityU 11304820, CityU 11300021, CityU 11311722, and CityU 11305523), and in part by the Hong Kong Polytechnic University internal grant (No.
+P0045335).
 For the purpose of open access, the authors have applied a CC BY public copyright license to any Author Accepted Manuscript (AAM) version arising from this submission.
@@
 Data availability: Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.
+Funding Declaration: The research was supported in part by the UK Engineering and Physical Sciences Research Council Awards EP/L015811/1, EP/V008854, and EP/V051121/1; the US National Science Foundation under Grants DMS-2054689 and DMS-2219391, and the Steenbock Professorship Award from the University of Wisconsin-Madison; and the Research Grants Council of the HKSAR (China) Project Nos.
+CityU 11304820, CityU 11300021, CityU 11311722, and CityU 11305523, and the Hong Kong Polytechnic University internal grant (No.
+P0045335).